### v1

# Data Compression for Fast Online Stochastic Optimization

Abstract
We propose an online data compression approach for efficiently solving distributionally robust optimization (DRO) problems with streaming data while maintaining out-of-sample performance guarantees. Our method dynamically constructs ambiguity sets using online clustering, allowing the clustered configuration to evolve over time for an accurate representation of the underlying distribution. We establish theoretical conditions for clustering algorithms to ensure robustness, and show that the performance gap between our online solution and the nominal DRO solution is controlled by the Wasserstein distance between the true and compressed distributions, which is approximated using empirical measures. We provide a regret analysis, proving that the upper bound on this performance gap converges sublinearly to a fixed clustering-dependent distance, even when nominal DRO has access, in hindsight, to the subsequent realization of the uncertainty. Numerical experiments in mixed-integer portfolio optimization demonstrate significant computational savings, with minimal loss in solution quality.
1 Introduction
Many decision-making problems in engineering, operations research, and computer science involve solving optimization problems affected by uncertain parameters. A common approach to address this class of problems is stochastic optimization (SO), which models uncertain parameters as random variables with known probability distributions. Typically, SO minimizes an expected cost based on this information [48]. While effective, this method relies on the assumption of complete knowledge of the underlying distributions—an assumption that is often unrealistic. If violated, SO produces suboptimal solutions that can lead to catastrophic consequences in the case of safety-critical systems. When faced with distributional ambiguity, i.e., uncertainty in the probability distribution, DRO provides a disciplined framework to robustify the decision-making process against distributional misspecification. Instead of assuming a known distribution, DRO considers an ambiguity set of possible distributions and optimizes for the worst-case expected cost under all distributions within this set. This approach is particularly appealing as it offers out-of-sample performance guarantees with a finite number of data samples. By accounting for ambiguity, DRO often outperforms non-robust methods such as sample average approximation (SAA) with respect to worst-case performance [22].
Since its appearance, significant effort has been devoted to constructing ambiguity sets primarily in static settings, where a single set is used throughout; see [16, 46, 40, 22], and the references therein. Even in multi-stage DRO, where uncertain data is revealed sequentially, a single ambiguity set is typically constructed, with time progression modeled through adjustable wait-and-see variables [18, 8, 6]. Fundamentally, this approach determines decisions based on a priori information. However, in many real-time applications—such as healthcare resource allocation, air traffic control, and energy trading—new data becomes available sequentially over time. As a result, the optimizer’s confidence in the nominal distribution can improve through the continuous assimilation of new data. Despite its potential, DRO with streaming data remains largely unexplored. In the few cases where it has been studied, updates to the DRO problem rely on gradient-based methods that either restrict the distribution to a finite set of scenarios [3], or require continuous decision variables and objective functions with local strong concavity [37]. In this work, we focus on the popular Wasserstein-based DRO [40, 36, 22], where the ambiguity set is defined as a Wasserstein ball around the empirical distribution of a dataset. As new data-points arrive, the problem dimension grows, increasing computational complexity [36, 56]. While a larger dataset improves confidence in the empirical measure as an approximation of the true distribution, the curse of dimensionality [21] ensures that the reduction in the ambiguity set’s radius is too slow to offset this complexity, even for moderately sized uncertain parameters. This leads to a computational bottleneck, and is especially limiting for real-time applications.
1.1 Our contributions
We propose an online data compression approach for efficiently solving Wasserstein DRO problems with streaming data, while maintaining out-of-sample performance guarantees. Our key contributions are:
  • Adaptive ambiguity sets via online clustering: We apply online clustering to construct adaptive ambiguity sets, formulated as Wasserstein balls of any order around a clustered empirical distribution. We establish theoretical conditions on clustering algorithms to ensure the robustness of our framework, which is compatible with any online clustering method that meets these conditions. We formalize the concept of optimal clustering, and provide a few fast and memory-efficient approximation algorithms suitable for our purposes.
  • Clustering effect analysis: We provide certificates, i.e., upper bounds on the out-of-sample performance of our algorithm, and prove that the performance gap between our compressed solution and the non-compressed DRO solution is controlled by the Wasserstein distance between the true distribution and its clustered approximation. This distance is approximated using empirical measures. We show that the objective converges to a value dependent on the number of clusters and the quality of the clustering algorithm. This quantifies the impact of data compression, and highlights the trade-off between computational effort and optimality.
  • Online performance analysis: We provide a regret analysis comparing the performance of our algorithm to the one of non-compressed DROin hindsight, i.e., with access to the subsequent realization of the uncertainty. We show that the regret converges sublinearly to a function of the Wasserstein distance between the true and compressed distributions, and with probability 1, the difference induced by the time-discrepancy converges to 0.
  • Computational gains in numerical experiments: We demonstrate the efficiency of our approach in sparse portfolio optimization. Our results show significant computational savings, even compared to SAA, with minimal loss in solution quality. We also demonstrate the possibility for the online framework to be memory-efficient, allowing the optimizer to discard data-points once seen, with minimal impact on solution times and quality.
1.2 Related work
Distributionally robust optimization
DRO has been extensively explored in recent years, with successful applications (among others) in machine learning [47], finance [10], and medicine [52]. Typical ambiguity sets that appeared in the literature include support, moment, or distance-based sets of distributions or mixtures thereof. We focus our attention on discrepancy-based ambiguity sets, defined as a ball in the space of probability distributions around a nominal or most-likely distribution, which is constructed from data. In this setting, the distance, commonly expressed in terms of, e.g., the $\phi-$divergence [5], the total variation norm [53], the kernel mean embedding [25], contamination techniques [35], or optimal transport based-distances including the celebrated Wasserstein distance [40, 32], signifies the “trust” in the data at hand. More recently, extensions, such as the trade-off ambiguity set [51] and the globalized ambiguity set [38], have been introduced to mitigate some of the conservatism of the classical DRO models. Due to the favorable properties of Wasserstein distance in terms of expressivity and statistical properties, we will focus on Wasserstein DRO.
Online learning and DRO
Online learning is a well-established framework providing algorithms for solving repetitive problems over time. Recently, it has been applied to robust optimization problems [28, 17], as well as to DRO formulations. For example, [43] considers an online DRO model with $\phi-$divergence ambiguity sets and proposes an alternating mirror descent algorithm to solve it, while [45] proposes a duality-free online stochastic method for regularized DRO problem with KL-divergence regularization. We focus here on a different problem, where the same DRO problem needs to be solved repeatedly over time with a growing knowledge of the uncertainty distribution. In this sense, our work is more closely related to [3], where a projected gradient descent method is used to adapt the ambiguity set to the samples collected progressively over time. However, [3] only considers discrete distributions. As a result, this work naturally disregards the fundamental computational challenge of incorporating streaming data in continuous settings, which is instead our main focus. Recently, [37] also considers the online DRO problem with data assimilation, but restricts their framework to continuous variables and strongly concave functions with accessible gradients. Our approach is therefore more general, and importantly, allows for mixed-integer formulations.
Scenario reduction and DRO
It is well-known that the size and computational complexity of data-driven optimization problems generally increase with the number of samples, resulting in a fundamental trade-off between statistics and computation. To overcome the computational bottleneck, scenario reduction techniques are often applied. They aim to reduce the number of scenarios while retaining a good enough representation of the underlying uncertainty and thus an accurate solution to the data-driven optimization problem. A popular approach consists of generating new scenarios and assigning probabilities to minimize the distance to the original distribution [19]. More recently, decision-focused scenario reduction techniques have been proposed, where the loss function itself is used in the construction of metrics to aggregate scenarios [7, 58]. A variety of techniques to aggregate scenarios have been suggested in the literature, based on clustering [27], moment matching [39], objective approximation [57], and nested distances [30]. We consider scenario reduction for DRO problems closely related to recent developments in two-stage robust optimization, such as [7, 58]. Our previous work [56] studies the impact of clustering in static and finite-sample DRO problems, using it as a tool to bridge robust and distributionally robust optimization. Closely related is also [4], which embeds scenario reduction into a DRO framework and provides suboptimality bounds—albeit limited to monotonically homogeneous uncertain objectives with strictly positive uncertainty. However, both [56] and [4] assume that data are available a priori, while our focus here is on online scenario reduction for DRO problems with streaming data.
Clustering with streaming data
Due to the ongoing data revolution, data stream clustering has recently attracted attention for emerging applications that involve large amounts of streaming data. They can be broadly classified into partition-based algorithms that partition data into clusters using distance-based similarity metrics, density-based algorithms that define clusters as dense partitions separated by sparse areas that dynamically change over time to adapt to data evolving distributions (such as DenStream [14]), hierarchical-based algorithms that maintain a tree-like structure by grouping similar clusters at different levels [42]. Among the first class, the most popular ones are undoubtedly incremental $k$-mean [1] and CluStream [2], which we adapt to fit our framework. We highlight that a major strength of our approach is its ability to incorporate state-of-the-art online clustering algorithms, with various memory and run-time complexities, with the choice left to the user’s discretion. In this work, we adapt two algorithms for our purposes.
1.3 Layout of the paper
In Section 2, we state the problem and introduce our online algorithm, and in Section 3, analyze the clustering effect by providing finite sample and asymptotic performance guarantees. In Section 4, we give dynamic regret bounds, and in Section 6, we describe various online clustering algorithms. In Section 7, we give guidelines for choosing the number of clusters and the ambiguity set radii experimentally, and in Section 8 we demonstrate our results on a portfolio optimization example.
2 The online stochastic problem
Consider a stochastic optimization problem of the form
$H_{\star}=\min_{x\in\mathcal{X}}\mathbf{E}_{u\sim\mathbf{P}}[f(u,x)],$ (1)
where $x\in\mathcal{X}\subseteq\mathbf{R}^{n}$ is the decision variable, $\mathcal{X}$ a compact set, $u\in S\subseteq\mathbf{R}^{d}$ the vector of uncertain parameters that is governed by some probability distribution $\mathbf{P}$, and $H_{\star}$ the optimal objective value. We assume that the support $S$ of ${\mathbf{P}}$ lives within the domain of $f$ for the variable $u$, which we will refer to as $\mathop{\bf dom}_{u}{f}$, i.e., $S\subseteq\mathop{\bf dom}_{u}{f}$. The function $f:\mathcal{X}\times S\rightarrow(-\infty,\infty]$ is assumed to be of the form
$f(u,x)=\max_{j\leq J}f_{j}(u,x),$
with each $f_{j}$ being proper, concave, and upper-semicontinuous in $u$ for all $x$. Additionally, we require the functions $f_{j}$ to be either Lipschitz or smooth for all $x\in\mathcal{X}$, according to the following definitions. Note that we assume the existence of the global constants by the compactness of $\mathcal{X}$.
Definition 2.1 (Lipschitzness).
A function $f(u,x)$ is $M$-Lipschitz on its domain with respect to the $\ell_{2}$-norm, if for all $x\in\mathcal{X}$,
$\left|f(v,x)-f(u,x)\right|\leq M\|u-v\|_{2},\quad\forall u,v\in\mathop{\bf dom}_{u}f.$
Definition 2.2 (Smoothness).
A differentiable function $f(u,x)$ is $L$-smooth on its domain, with respect to the $\ell_{2}$-norm, if for all $x\in\mathcal{X}$,
$\left\|\nabla f(v,x)-\nabla f(u,x)\right\|_{2}\leq L\|u-v\|_{2},\quad\forall u,v\in\mathop{\bf dom}_{u}f.$
Analogously to [56], we also assume $\mathop{\bf dom}_{u}{f}$ is independent of $x$, and satisfies the following assumption:
Assumption 2.1.
The domain $\mathop{\bf dom}_{u}{f}$ is ${\mbox{\bf R}}^{d}$. Otherwise, $f$ is either element-wise monotonically increasing in $u$ and only has a (potentially) lower-bounded domain, or element-wise monotonically decreasing in $u$ and only has a (potentially) upper-bounded domain.
Streaming data
We assume ${\mathbf{P}}$ to be unknown, and that the value $H_{\star}$ cannot be calculated as is. However, we have access to a streaming dataset of i.i.d. realization of $u$, which we use to construct an uncertainty framework. We begin with an initial dataset with $n_{0}\geq 1$ data-points, i.e., $\mathcal{D}_{0}=\{\hat{u}^{i}\}_{i=1}^{n_{0}}$. Then, at the end of each time step $t\geq 1$, the dataset is updated to be $\mathcal{D}_{t}=\{\hat{u}^{i}\}_{i=1}^{n_{0}+t}$, with $|\mathcal{D}_{t}|=n_{t}=n_{0}+t$. That is, we assume w.l.o.g. that one realization gets disclosed at the end of each round. Furthermore, we define the empirical distribution
$\hat{{\mathbf{P}}}_{t}=\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\delta_{\hat{u}^{i}},$
where $\delta_{u}$ is the Dirac distribution concentrating unit mass on $u$.
Distributionally robust optimization over time
[…]
$\mathcal{P}_{t}=\mathbf{B}_{\varepsilon_{t}}^{p}(\hat{{\mathbf{P}}}_{t})=\left\{\mathbf{Q}\in\mathcal{W}({S})\mid W_{p}(\hat{{\mathbf{P}}}_{t},\mathbf{Q})\leq\varepsilon_{t}\right\},$ (2)
where the Wasserstein type-$p$ distance between two probability distributions $\mathbf{P},\mathbf{Q}\in\mathcal{W}(S)$ is defined by
$W_{p}(\mathbf{P},\mathbf{Q})=\inf_{\pi\in\Pi(\mathbf{P},\mathbf{Q})}\left(\int_{S\times S}\left\|u_{1}-u_{2}\right\|^{p}\mathrm{d}\pi\left(u_{1},u_{2}\right)\right)^{1/p},$ (3)
and $\Pi(\mathbf{P},\mathbf{Q})$ is the set of all probability distributions over $S\times S$ with marginals $\mathbf{P}$ and $\mathbf{Q}$. Intuitively, the ambiguity set $\mathcal{P}_{t}$ contains all probability distribution in $\mathcal{W}(S)$ that can obtained by transporting probability mass from $\hat{{\mathbf{P}}}_{t}$ when the transportation budget is at most $\varepsilon_{t}$. With this uncertainty modeling framework in place, we compute the DRO solution $x_{t}\in\mathcal{X}$ at time step $t$ by solving the optimization problem
$H_{t}=\min_{x\in\mathcal{X}}\max_{\mathbf{Q}\in\mathcal{P}_{t-1}}\mathbf{E}_{u\sim\mathbf{Q}}[f(u,x)],$ (4)
where we seek to minimize the worst-case (w.r.t. any distribution in the ambiguity set) expected value of $f$. Note that we use the ambiguity set $\mathcal{P}_{t-1}$ constructed at time $t-1$: in the online framework, we assume to solve the problem before the new data-point is revealed.
As $t$ increases, our confidence in $\hat{{\mathbf{P}}}_{t}$ grows, allowing us to safely reduce the radius $\varepsilon_{t}$. This radius can be interpreted as a measure of trust in the data, and the cost of robustness. This intuition is supported by the results in [20], which show that under mild assumptions, choosing $\varepsilon_{t}\sim\text{o}(t^{-\frac{1}{d}})$ ensures that the true distribution $\mathbf{P}\in\mathcal{P}_{t}$ with high probability. From [40], this choice of radius leads to the following finite sample performance guarantee
${\mathbf{P}}^{t-1}\left(H_{\star}\leq{\mbox{\bf E}}_{{\mathbf{P}}}[f(u,x_{t})]\leq H_{t}\right)\geq 1-\beta_{t-1},$ (5)
where for all $t\geq 1$, $\beta_{t-1}>0$ is a specified time-varying probability of constraint violation, and ${\mathbf{P}}^{t-1}$ is the product distribution of the dataset $\mathcal{D}_{t-1}$. The in-sample objective value $H_{t}$ is then a certificate for the out-of-sample performance ${\mbox{\bf E}}_{{\mathbf{P}}}[f(u,x_{t})]$. Furthermore, [40] prove that an appropriately decreasing choice of $\varepsilon_{t}$ and $\beta_{t}$ as $t\to\infty$ leads to asymptotic guarantees on the convergence of $H_{t}$ to the true value $H_{\star}$. These results imply, by incorporating streaming data into Problem (4), we can adapt the ambiguity set and gradually reduce the cost of robustness.
Our online algorithm
While incorporating streaming data improves the solution quality, it also increases the computational burden, scaling with the number of data-points [40]. We thus propose a data-compression framework for fast online DRO. Our goal is to solve an adjusted DRO problem over time, where
$H^{K}_{t}=\min_{x\in\mathcal{X}}\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t-1}}\mathbf{E}_{u\sim\mathbf{Q}}[f(u,x)],\quad t=1,2,\dots$ (6)
and $\mathcal{P}^{K}_{t-1}$ is an adaptive ambiguity set, with adaptive radius $\varepsilon^{K}_{t-1}\geq 0$, constructed from the data $\mathcal{D}_{t-1}$ observed so far. From here onward, we refer to the solution $x^{K}_{t}$ as the online solution, and the solution $x_{t}$ from (4) as the nominal DRO solution.
Definition 2.3 (Adaptive ambiguity set).
The adaptive ambiguity set at the end of time period $t$, with radius $\varepsilon^{K}_{t}\geq 0$ and reference distribution $\hat{\mathbf{P}}^{K}_{t}$, is given by
$\mathcal{P}^{K}_{t}=\mathbf{B}_{\varepsilon^{K}_{t}}^{p}(\hat{\mathbf{P}}^{K}_{t})=\left\{\mathbf{Q}\in\mathcal{W}({S})\mid W_{p}(\hat{\mathbf{P}}^{K}_{t},\mathbf{Q})\leq\varepsilon^{K}_{t}\right\}.$ (7)
Unlike the standard DRO ambiguity set, the adaptive ambiguity set is compressed in that we allow the reference distribution to be supported on only $K\leq n_{t}$ points, to greatly reduce the computational burden. With access to the dataset $\mathcal{D}_{t}$, the reference distribution is formulated by clustering the data-points into sets $C^{k}_{t}$, $k=1,\dots,K_{t},$ where $K_{t}\leq K$. The weight $\theta^{k}_{t}$ of a cluster $C^{k}_{t}$ is defined as the proportion of total points in the cluster, and the cluster centroid $\bar{u}^{k}_{t}$ is the mean of these points. These values characterize the clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$i.e.,
$\theta_{t}^{k}=\frac{n^{k}_{t}}{n_{t}},\quad\bar{u}^{k}_{t}=\frac{1}{n^{t}_{k}}\sum_{\hat{u}\in C^{k}_{t}}\hat{u},\quad\hat{\mathbf{P}}^{K}_{t}=\sum_{k=1}^{K_{t}}\theta_{t}^{k}\delta_{\bar{u}_{t}^{k}},$
where $n^{k}_{t}=|C^{k}_{t}|$. In Section 3.2, we provide additional details on our clustering requirements, and in Section 6, provide efficient online clustering methods.
With an adaptive ambiguity set $\mathcal{P}^{K}_{t-1}$, we solve the online compressed DRO problem (6) through a direct reformulation approach [56, Section 2.4], given as
$\begin{array}[]{lll}\!\!H^{K}_{t}\!=\!\!\!&\underset{x\in\mathcal{X},\lambda_{t}\geq 0,s_{t},z_{t},y_{t}}{\mbox{minimize}}&\sum_{k=1}^{K_{t}}\theta^{k}_{t-1}s^{k}_{t}\\ &\mbox{subject to}&[-f_{j}]^{*}(z^{jk}_{t}-y^{jk}_{t},x)+\sigma_{{S}}(y^{jk}_{t})-(z{{}^{jk}_{t}})^{T}\bar{u}^{k}_{t}+\phi(q)\lambda_{t}\|z^{jk}_{t}/\lambda_{t}\|_{*}^{q}\\ &&\hskip 80.00012pt+\lambda_{t}\varepsilon_{t-1}^{p}\leq s^{k}_{t},\quad k=1,\dots,K_{t},\,j=1,\dots,J\\ \end{array}\!\!\!\!\!\!\!\!\!\!\!\!\!$ (8)
[…]
In summary, the online problem follows a sequential process for $t\geq 1$:
  1. 1.
    We compute decision $x^{K}_{t}$ by solving (8) with ambiguity set $\mathcal{P}^{K}_{t-1}$;
  2. 2.
    We observe a new realization of the uncertainty $u$, denoted as $\hat{u}$;
  3. 3.
    We update the ambiguity set to $\mathcal{P}^{K}_{t}$ with the new data point.
To assess the performance of our online solution scheme, in Section 3 we provide finite-sample and asymptotic performance guarantees of a similar form as (5), where we analyze the effects of clustering. In Section 4, we further consider the dynamic regret, which measures the performance gap between our online solution and the DRO solution in hindsight at time step $t$, both evaluated with respect to the non-compressed and updated ambiguity set $\mathcal{P}_{t}$. The notion of regret is in line with literature on online learning [3, 44, 9], where a time-discrepancy (which indicates an information-discrepancy) is introduced between the online solution and an oracle. Formally, we quantify this gap as
$R(T,K)=\frac{1}{T}\sum_{t=1}^{T}\left(\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]-\min_{x\in\mathcal{X}}\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x)]\right),$ (9)
where we compare the average performance of our online solution $x^{K}_{t}$ against the average performance of the oracle: the DRO solution in hindsight.
3 Clustering effect analysis
In this section, we state our main results: the effect of clustering on the finite samples and asymptotic performance guarantees for the online algorithm. These are confidence bounds, similar to (5), on the in-sample and out-of-sample performance of the online solution $x^{K}_{t}$. To provide the preliminaries, we break this section into several subsections. In Section 3.1, we state key measure concentration results. Next, in Section 3.2, we provide requirements on the online clustered distribution $\hat{{\mathbf{P}}}^{K}_{t}$; later, in Section 6, we provide specific algorithms to formulate this distribution. In Section 3.3, we then analyze this clustered distribution, including its distance from the empirical distribution $\hat{{\mathbf{P}}}_{t}$, its convergence to a limiting distribution, and its convergence rate. Finally, making use of the above information, in Section 3.4 we give the main results.
3.1 Measure concentration and nominal results
In order to obtain performance guarantees, we require the distribution ${\mathbf{P}}$ to satisfy certain assumptions. We impose the following light-tailed assumption, which is satisfied by the common class of sub-Gaussian distributions, and by bounded distributions.
Assumption 3.1 (Light-tailed distribution).
There exists an exponent $r>1$, such that $R={\mbox{\bf E}}_{\mathbf{P}}[\exp(\|u\|^{r})]=\int_{S}\exp(\|u\|^{r}){\mathbf{P}}(du)<\infty.$
This leads to the following measure concentration results.
Theorem 3.1 (Measure concentration, light-tailed [20]).
If Assumption 3.1 holds and $p=1$, for the distribution ${\mathbf{P}}$ and empirical distribution $\hat{{\mathbf{P}}}^{N}$ supported on $N$ points, we have
$\mathbf{P}^{N}(W_{1}(\mathbf{P},\hat{\mathbf{P}}^{N})\geq\varepsilon)\leq\begin{cases}c_{1}\exp(-c_{2}N\varepsilon^{\max\{d,2\}})\quad\textit{if }\varepsilon\leq 1\\ c_{1}\exp(-c_{2}N\varepsilon^{r})\quad\textit{if }\varepsilon>1,\end{cases}$
for all $\varepsilon>0$, $d\neq 2$, and $N\geq 1$. When $p\geq 2$ and $r\in(0,p)$, we have
$\mathbf{P}^{N}(W_{p}(\mathbf{P},\hat{\mathbf{P}}^{N})\geq\varepsilon)\leq\begin{cases}c_{1}\exp(-c_{2}N\varepsilon^{\max\{d,2p\}})+\exp(-c_{2}(N\varepsilon^{p})^{\frac{r-s}{p}})\\ \hfill\textit{if }\varepsilon\leq 1,\forall s\in(0,r)\\ c_{1}\exp(-c_{2}(N\varepsilon^{p})^{r/p})\quad\textit{if }\varepsilon>1,\end{cases}$
for all $\varepsilon>0$, $d\neq 4$, and $N\geq 1$. For all cases, $c_{1},c_{2}$ are positive constants depending only on $R,r,s,$ and $d$.
For the special case $p=\infty$, we point to [50, Theorem 1.1] for an analogous result, which assumes a stronger condition than Assumption 3.1. Using these relations, the radius for a Wasserstein DRO problem with $N=n_{t-1}$ points can be chosen by setting the right-hand-side to $\beta_{t-1}$, then solving for $\varepsilon_{t-1}$. See [40, Theorem 3.4] for the case $p=1$. The procedure for $p\geq 2$ follow similarly. In Section 5, we further calculate explicit radii for the specific case of bounded support $S$. In all cases, we note that the radius is inversely related to the number of data-points, i.e., the larger the number $n_{t}$, the smaller $\varepsilon_{t}$ can be. With the radius chosen in this way, we obtain the DRO performance guarantees (5) using [40, Theorem 3.5].
From the above characterization, it is evident that larger datasets are beneficial for reducing the distributional ambiguity around the nominal distribution. Indeed, for $n_{t}\rightarrow\infty$, we can choose a confidence sequence $\beta_{t}\in(0,1)$ such that $\sum_{t=1}^{\infty}\beta_{t}<\infty$ and $\lim_{t\rightarrow\infty}\varepsilon_{t}=0$ to conclude that $\mathbf{P}^{\infty}\{\lim_{t\rightarrow\infty}W_{p}(\mathbf{P},\hat{\mathbf{P}}_{t})\}=1$, which then implies $\lim_{t\rightarrow\infty}H_{t}\downarrow H_{\star}$ [40, Theorem 3.7]. However, the increase in the number of data-points makes the resolution of the DRO problem in (4) computationally challenging, as a constraint needs to be added for each data-point constituting the empirical distribution $\hat{\mathbf{P}}_{t}$ (see the problem formulation (8)). Clearly, for $t\rightarrow\infty$, we would end up with an infinite-dimensional problem that cannot be solved.
3.2 Clustering procedure requirements
To address the above dilemma, the key of our approach is the adaptive ambiguity set $\mathcal{P}^{K}_{t}$, which is constructed using a compressed representation of the available data, $\hat{{\mathbf{P}}}^{K}_{t}$. In this section, we formalize the requirements on generating $\hat{{\mathbf{P}}}^{K}_{t}$, and note that the decision-maker has the flexibility to use any method to generate $\hat{{\mathbf{P}}}^{K}_{t}$, provided it satisfies the given assumptions. These requirements ensure the validity of the asymptotic performance guarantees established later in this section, and can be satisfied by any partitioning of the support set $S$.
Regardless of the method, the overall goal is to find the clustering that minimizes the discrepancy between the clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$ and the true distribution ${\mathbf{P}}$, which is approximated by the empirical distribution $\hat{{\mathbf{P}}}_{t}$. In this section, we also outline the theoretical procedure to find the optimal clustering, and note that the goal of the online procedure is to find the best approximation of the optimal clustering, while also maintaining a low computational complexity.
3.2.1 Requirements
Let $K$ denote the maximum number of data points (that is, of constraints) that we can handle based on the available computational budget. We thus cluster data points together, and allow up to $K$ distinct clusters. For all methods and time periods $t$, we then maintain a set of $K_{t}\leq K$ clusters $\{C_{t}^{k}\}_{k=1}^{K_{t}}$ and a corresponding set of supports $\mathcal{S}_{t}=\{S_{t}^{k}\}_{k=1}^{K_{t}},$ where each one is defined as follows.
Definition 3.1 (Cluster support).
For each cluster $k$ and time $t$, the cluster support $S_{t}^{k}\subset S$ is defined as a region in $S$ such that any data-point $\hat{u}\in\mathcal{D}_{t}$ that falls within this region is contained in cluster $C_{t}^{k}$. We also call these clusters the clustering induced by the cluster support.
The cluster supports must satisfy the following assumption.
Assumption 3.2.
The following holds for any time step $t$,
  • The elements in the set $\mathcal{S}_{t}=\{S_{t}^{k}\}_{k=1}^{K_{t}}$ have pairwise disjoint interiors, i.e.,
    $\mathrm{int}(S^{k}_{t})\cap\mathrm{int}(S^{k^{\prime}}_{t})=\emptyset,~{}\forall S^{k},S^{k^{\prime}}\in\mathcal{S}_{t},k\neq k^{\prime}.$
  • The cluster supports cover $S$, i.e., $S=\cup_{k\leq K_{t}}S_{t}^{k}$.
  • The cluster supports have nonzero measure, i.e., ${\mathbf{P}}(S^{k}_{t})>0$ for $k=1,\dots,K_{t}$ and for all $t$.
Furthermore, there is some finite time $\tau<\infty$ such that $\mathcal{S}_{t}=\mathcal{S}_{\tau}$ for all $t\geq\tau$.
These assumptions ensure that all data-points fall in a single cluster support, and could be satisfied by any partitioning of the original support set $S$. Note that until some finite time $\tau$, the data-points are allowed to switch cluster assignments, which increases the flexibility of the clustering algorithm. By Assumption 3.2, the cluster weights $\sum_{k=1}^{K_{t}}\theta^{k}_{t}=1$, and $\hat{\mathbf{P}}^{K}_{t}$ represents an approximation of the non-clustered empirical distribution $\hat{\mathbf{P}}_{t}$, supported on $K_{t}\leq n_{t}$ atoms.
3.2.2 Optimal clustering
If the distribution ${\mathbf{P}}$ is known, then for any given $K$, the optimal set of supports can be determined using the $k$-centers clustering approach described in [24, Chapter 1]. Specifically, we solve the $k$-centers problem to obtain an optimal set of centers $A^{\star}\subset S$, then construct the Voronoi diagram of $A^{\star}$. Each cluster support is defined as $S^{k}=\left\{u\in S\,\middle|\,\|u-a^{k}\|=\min_{a\in A^{\star}}\|u-a\|\right\},$ where each point $u$ is assigned to the region corresponding to its closest center $a^{k}$.
Definition 3.2 ($k$-centers problem).
The $k$-centers problem, with respect to a distribution ${\mathbf{P}}$ and orders $p\geq 1$ (left), $p=\infty$ (right) is given as
$\inf_{A\subset S,|A|\leq K}{\mbox{\bf E}}_{\mathbf{P}}\left[\min_{a\in A}\|u-a\|^{p}\right],~{}\quad\inf_{A\subset S,|A|\leq K}\sup_{u\in S}\min_{a\in A}\|u-a\|,$ (10)
where $A^{\star}$ is the set that achieves this infimum.
We also define the minimium quantization error of ${\mathbf{P}}$, and note a subsequent proposition.
Definition 3.3 (Minimum quantization error [24, Chapter 1]).
The minimum quantization error of a distribution ${\mathbf{P}}$ is the minimum Wasserstein-$p$ distance between ${\mathbf{P}}$ and any discretization of ${\mathbf{P}}$ to at most $K$ atoms, given as
$d^{K}_{\star,p}({\mathbf{P}})=\left({\mbox{\bf E}}_{\mathbf{P}}\left[\min_{a\in A^{\star}}\|u-a\|^{p}\right]\right)^{1/p},\quad d^{K}_{\star,\infty}=\sup_{u\in S}\min_{a\in A^{\star}}\|u-a\|,$
where $A^{\star}$ is the set of optimal centers.
Proposition 3.1.
[…]
Proof.
This holds since $\hat{{\mathbf{P}}}^{K}_{t}$ is a particular discretization of ${\mathbf{P}}$ to $K$ atoms. ∎
Without knowledge of ${\mathbf{P}}$ but having access to a dataset $\mathcal{D}_{t}$, the optimal clustering with respect to the empirical distribution $\hat{\mathbf{P}}_{t}$ can also be found through the above process, where ${\mathbf{P}}$ is replaced by $\hat{\mathbf{P}}_{t}$. However, the $k$-centers problem is NP-hard [29], and is therefore inefficient to solve. Furthermore, with $\hat{\mathbf{P}}_{t}$ as the reference distribution, when $t\rightarrow\infty$, the computational complexity of the $k$-centers problem increases, conflicting with our goal of decreasing computational effort. Thus, the goal of the online clustering algorithm is to find a good approximation of the $k$-centers problem, while also minimizing computational complexity. In Section 6, we give a few algorithms that satisfy the given requirements.
3.3 Clustered distribution analysis
Since the performance bounds will be affected by the sequence of clustered empirical distributions $\hat{{\mathbf{P}}}^{K}_{t}$, we first derive some results pertaining these distributions. In particular, we formalize its distances, under different metrics, to the empirical distribution $\hat{{\mathbf{P}}}_{t}$, as well as its convergence to a limiting distribution ${\mathbf{P}}^{K}_{\star}$. We define the following values.
Definition 3.4.
For any clustered configuration of the dataset $\mathcal{D}_{t}$, with clusters $C^{k}_{t}$, cluster centroids $\bar{u}^{k}_{t}$, and clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$, we define the Wasserstein cost $d^{K}_{t,p}$ and clustering value $D^{K}_{t,p}$ as
$d^{K}_{t,p}=W_{p}(\hat{{\mathbf{P}}}_{t},\hat{{\mathbf{P}}}^{K}_{t}),\quad D^{K}_{t,p}=\left(\frac{1}{n_{t}}\sum_{k=1}^{K_{t}}\sum_{\hat{u}\in C^{k}_{t}}\left\|\hat{u}-\bar{u}^{k}_{t}\right\|_{2}^{p}\right)^{1/p}.$
At each time step, we also define a function of dual variables $z^{jk}_{t}$ of (8),
$\Phi^{K}_{t}=\mathbf{1}\{J\geq 2\}\frac{1}{n_{t}}\sum_{k=1}^{K_{t}}\sum_{\hat{u}\in C^{k}_{t}}\max_{j\leq J}(-z_{t}^{{jk}})^{T}(\hat{u}-\bar{u}^{k}_{t}),$
where $\mathbf{1}\{\cdot\}$ is the 00-$1$ indicator function. Note that $\Phi^{K}_{t}=0$ when $J=1$, i.e., when $f$ is a concave function.
By construction, the above values all quantify the distance between the clustered and the non-clustered empirical distributions, under different metrics, and are nonnegative. In addition, we establish a hierarchy between these distances and provide a few bounds, making use of the following definition.
Definition 3.5.
The diameter of a set $S$ is defined
$\mathrm{diam}_{q}(S)=\max\{\|u-v\|_{q}\mid u,v\in S\}.$
Proposition 3.2.
[…]
$D^{K}_{t,p}\leq 2\rho,\quad\Phi^{K}_{t}\leq\mathbf{1}\{J\geq 2\}\max_{k\leq{K_{t}}}\max_{j\leq J}2\rho\|z^{jk}_{t}\|.$
Proof.
By definition, the Wasserstein distance is computed with the optimal coupling (joint distribution) between the two reference distributions, while $D^{K}_{t,p}$ is calculated from one particular coupling. Therefore, $d^{K}_{t,p}\leq D^{K}_{t,p}$. The lower bound on $d^{K}_{t,p}$ follows from Proposition 3.1, and the equalities follow from the fact that $\hat{{\mathbf{P}}}^{K}_{t}=\hat{{\mathbf{P}}}_{t}$ when $K\geq n_{t}$. The final inequalities follow from the boundedness of the support. ∎
Furthermore, we can derive asymptotic bounds, for $t\to\infty$. This relies on Assumption 3.2, that for $t\geq\tau$, where $\tau<\infty$ is some finite time, the cluster supports $S_{t}^{k}$ are fixed. We make use of the following lemma.
Lemma 3.1 (Clustering convergence).
Under Assumption 3.2 on the cluster supports, the corresponding sequence of clustered empirical distributions $\hat{{\mathbf{P}}}^{K}_{t}$ converges almost surely with respect to ${\mathbf{P}}^{\infty}$ to a distribution ${\mathbf{P}}^{K}_{\star}$. Furthermore, the distributions achieve almost sure convergence with respect to the Wasserstein metric, that is,
${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}W_{p}({\mathbf{P}}^{K}_{\star},\hat{{\mathbf{P}}}^{K}_{t})=0\right\}=1.$
Proof.
Without loss of generality, let $K_{\tau}=K$. By Assumption 3.2, the cluster supports are fixed. By the Strong Law of Large Numbers (SLLN) [34], for each cluster $k\leq K$,
${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\mathbf{1}\{\hat{u}^{i}\in S^{k}_{\tau}\}=\theta_{\star}^{k}\right\}=1.$
This shows that the cluster weights converge almost surely to ${\mathbf{P}}(S^{k}_{\tau})=\theta_{\star}^{k}$, the true weights on each cluster support.
Since, for all clusters $k$, ${\mathbf{P}}(S^{k}_{\tau})>0$, we note that $n^{k}_{t}\to\infty$ when $t\to\infty$. Furthermore, since ${\mathbf{P}}$ has finite $p$-th moments, the conditional distributions of ${\mathbf{P}}$ on these supports $S^{k}_{\tau}$ must also have finite $p$-th moments. We denote these conditional distributions ${\mathbf{P}}^{k}$, and their means $\bar{u}^{k}_{\star}$. We then observe, by the SLLN,
${\mathbf{P}}^{\infty}\left\{\lim_{n_{t}^{k}\to\infty}\frac{1}{n_{t}^{k}}\sum_{\hat{u}\in C^{k}_{\tau}}\hat{u}=\bar{u}^{k}_{\star}\right\}=1,$
i.e., the cluster centroids converge almost surely to the true means of the conditional distributions ${\mathbf{P}}^{k}$.
By Assumption 3.2, the cluster supports are pairwise disjoint, and cover the entire support $S$. It follows that $\sum_{k=1}^{K}{\mathbf{P}}(S^{k}_{\tau})=1$. This implies
${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}\hat{{\mathbf{P}}}^{K}_{t}=\lim_{t\to\infty}\sum_{k=1}^{K}\theta^{k}_{t}\delta_{\bar{u}_{t}^{k}}=\sum_{k=1}^{K}\theta^{k}_{\star}\delta_{\bar{u}_{\star}^{k}}={\mathbf{P}}^{K}_{\star}\right\}=1,$
where ${\mathbf{P}}^{K}_{\star}$ is a discretized distribution of ${\mathbf{P}}$ onto $K$ atoms. Since ${\mathbf{P}}^{K}_{\star}$ has finite $p$-th moments by construction, this implies almost sure convergence with respect to the Wasserstein-$p$ metric [55, Theorem 6.9]. ∎
We can then give an asymptotic bound on $d^{K}_{t,p}$ for $t\to\infty$.
Theorem 3.2 (Asymptotic clustering bounds).
Under the Lemma 3.1, we have
$\mathbf{P}^{\infty}\left\{d^{K}_{\star,p}({\mathbf{P}})\leq\lim_{t\rightarrow\infty}d^{K}_{t,p}\leq W_{p}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\right\}=1.$
Furthermore, as $K\to\infty$, $\lim_{K\rightarrow\infty}W_{p}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})=0$ for all $p$.
Proof.
By the triangle inequality, we note that for all $t$,
$d^{K}_{t,p}=W_{p}(\hat{{\mathbf{P}}}_{t},\hat{{\mathbf{P}}}^{K}_{t})\leq W_{p}(\hat{{\mathbf{P}}}_{t},{\mathbf{P}})+W_{p}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})+W_{p}({\mathbf{P}}^{K}_{\star},\hat{{\mathbf{P}}}^{K}_{t}),$
that is, the distance between the two empirical distributions is bounded by the sum of the distances between the nominal empirical distribution and ${\mathbf{P}}$, between ${\mathbf{P}}$ and the converged clustered distribution, and between converged clustered distribution and the clustered empirical distribution. Taking the limit as $t\to\infty$, we have
$\lim_{t\to\infty}d^{K}_{t,p}\leq\lim_{t\to\infty}W_{p}(\hat{{\mathbf{P}}}_{t},{\mathbf{P}})+W_{p}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})+\lim_{t\to\infty}W_{p}({\mathbf{P}}^{K}_{\star},\hat{{\mathbf{P}}}^{K}_{t}).$
Under the given assumptions, Theorem 3.1 implies
$\mathbf{P}^{\infty}\left\{\lim_{t\rightarrow\infty}W_{p}(\mathbf{P},\hat{{\mathbf{P}}}_{t})=0\right\}=1.$ (11)
Together with Lemma 3.1, we obtain $\mathbf{P}^{\infty}\left\{\lim_{t\rightarrow\infty}d^{K}_{t,p}\leq W_{p}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\right\}=1.$ For the lower bound, recall that by Proposition 3.1, $d^{K}_{\star,p}(\hat{{\mathbf{P}}}_{t})\leq d^{K}_{t,p}.$ The result then holds by the asymptotic consistency of finite quantizations, given in [24, Corollary 4.24], which holds due to (11). Specifically, we note that $\mathbf{P}^{\infty}\left\{\lim_{t\rightarrow\infty}d^{K}_{\star,p}(\hat{{\mathbf{P}}}_{t})=d^{K}_{\star,p}({\mathbf{P}})\right\}=1.$ Lastly, the asymptotic result for $K\to\infty$ holds from [24, Lemma 6.1]. ∎
The asymptotic limits of $D^{K}_{t,p}$ and $\Phi^{K}_{t}$ follow similarly.
Theorem 3.3 (Asymptotic clustering values).
Under Lemma 3.1, we have that
${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}D^{K}_{t,p}=D^{K}_{\star,p}=\left(\mathbf{E}_{{\mathbf{P}}}\left[\sum_{k=1}^{K}\mathbf{1}\{u\in S^{k}_{\tau}\}\|u-\bar{u}^{k}_{\star}\|^{p}_{2}\right]\right)^{1/p}\right\}.$
Furthermore, with accumulation points $\hat{z}^{jk}_{\star}=\lim_{t\to\infty}z_{t}^{jk}$, we have
${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}\Phi^{K}_{t}=\Phi^{K}_{\star}=\mathbf{1}\{J\geq 2\}\mathbf{E}_{{\mathbf{P}}}\left[\sum_{k=1}^{K}\mathbf{1}\{u\in S^{k}_{\tau}\}\max_{j\leq J}(z^{jk}_{\star})^{T}(u-\bar{u}^{k}_{\star})\right]\right\}.$
Proof.
This follows from the fixed cluster supports and the almost-sure convergence of both empirical distributions. ∎
From Lemma 3.1, we established that the clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$ converges to the distribution ${\mathbf{P}}^{K}_{\star}$. We can also prove that its convergence rate follows the convergence rate of the non-clustered empirical distribution.
Theorem 3.4 (Clustering convergence rate).
Suppose Assumption 3.2 holds, and we have a sequence $\varepsilon_{t}>0$, computed using Theorem 3.1, such that $\lim_{t\rightarrow\infty}\varepsilon_{t}=0$ and the corresponding confidence sequence $\beta_{t}\in(0,1)$ satisfies $\sum_{t=1}^{\infty}\beta_{t}<\infty$. Then, for $t\geq\tau$, with rate $O(\varepsilon_{t})$, the sequence of centroids $\bar{u}^{k}_{t}$ converges to the true centroids $\bar{u}^{k}_{\star}$, and the distance $W_{p}(\hat{{\mathbf{P}}}^{K}_{t},{\mathbf{P}}^{K}_{\star})$ converges to 0.
Proof.
By assumption, all cluster supports have nonzero measure, i.e., ${\mathbf{P}}(S^{k}_{\tau})>0$. Since ${\mathbf{P}}$ is light-tailed, we also know that for some $r>1$, there is some finite $R<\infty$ defined by Assumption 3.1. It follows that the conditional distributions of ${\mathbf{P}}$ on the supports $S^{k}_{\tau}$, denoted ${\mathbf{P}}^{k}$, are also light-tailed, with the same $r$ and a constant $R^{k}\leq O(R)$, i.e., $R^{k}=\int_{S_{\tau}^{k}}\exp(\|u\|^{r}){\mathbf{P}}(u|u\in S^{k}_{\tau})(du)<\infty.$ By Theorem 3.1, we have
${\mathbf{P}}^{n_{t}}(W_{p}({\mathbf{P}}^{k},\hat{{\mathbf{P}}}^{t,k})\leq O(\varepsilon_{t}))\geq 1-\beta_{t},$
where $\hat{{\mathbf{P}}}^{t,k}$ is the conditional distribution of $\hat{{\mathbf{P}}}_{t}$ on support $S^{k}_{\tau}$. By Kantorovich-Rubinstein duality and the ordering of Wasserstein distances,
$\left|{\mbox{\bf E}}_{{\mathbf{P}}^{k}}[u_{i}]-{\mbox{\bf E}}_{\hat{{\mathbf{P}}}^{k,t}}[u_{i}]\right|\leq W_{1}({\mathbf{P}}^{k},\hat{{\mathbf{P}}}^{k,t})\leq W_{p}({\mathbf{P}}^{k},\hat{{\mathbf{P}}}^{k,t}),\quad i=1,\dots,d.$
This implies that the centroids $\bar{u}^{k}_{t}$ converge to the true means $\bar{u}^{k}_{\star}$ with rate $O(\varepsilon_{t})$, with respect to the infinity norm, i.e., $\|\bar{u}^{k}_{t}-\bar{u}^{k}_{\star}\|_{\infty}\leq O(\varepsilon_{t}).$ Then, by the ordering of norms and the definition of the Wasserstein distance on discrete distributions,
$W_{p}(\hat{{\mathbf{P}}}^{K}_{t},{\mathbf{P}}^{K}_{\star})\leq\left(\sum_{k=1}^{K}|\theta^{k}_{t}-\theta^{k}_{\star}|\|\bar{u}^{k}_{t}-\bar{u}^{k}_{\star}\|^{p}\right)^{1/p}\leq O(\varepsilon_{t}),$
since the weights $\theta$ are bounded. ∎
3.4 Performance guarantees
We can now derive performance guarantees for the online problem. We begin by summarizing the assumptions.
Assumption 3.3 (Ambiguity set construction).
Let the following hold.
  1. 1.
    For all $t$, $\varepsilon_{t}$ is computed using Theorem 3.1, with a sequence $\beta_{t}$ such that $\sum_{t=0}^{\infty}\beta_{t}<\infty$ and $\lim_{t\to\infty}\varepsilon_{t}=0$.
  2. 2.
    The distribution $\hat{{\mathbf{P}}}^{K}_{t}$ is constructed using $\mathcal{S}^{k}$ that satisfy Assumption 3.2.
  3. 3.
    The adaptive ambiguity set $\mathcal{P}^{K}_{t}=\mathbf{B}_{\varepsilon^{K}_{t}}^{p}(\hat{\mathbf{P}}^{K}_{t})$, with $\varepsilon^{K}_{t}=\varepsilon_{t}$.
In addition, either the Lipschitz or smoothness condition must hold.
Assumption 3.4 (Lipschitzness and smoothness).
Let at least one hold.
  1. 1.
    For all $x\in\mathcal{X}$, the constituent functions $f_{j}$ of $f$ satisfy the Lipschitz condition given in Definition 2.1, with constants $M_{j}$.
  2. 2.
    For all $x\in\mathcal{X}$, the constituent functions $f_{j}$ of $f$ satisfy the smoothness condition given in Definition 2.2, with constants $L_{j}$.
We present the main results after the following lemma. Proofs are in Appendix A.1.
Lemma 3.2 (Effect of the support $S$).
Let $\Delta_{t}=\tilde{H}_{t}-H_{t}$, where $\tilde{H}_{t}$ is the value of the nominal DRO problem where the support $S={\mbox{\bf R}}^{d}$, i.e., the ambiguity set is defined $\tilde{\mathbf{B}}_{\varepsilon_{t-1}}^{p}(\hat{{\mathbf{P}}}_{t-1})=\left\{\mathbf{Q}\in\mathcal{W}({{\mbox{\bf R}}^{d}})\mid W_{p}(\hat{{\mathbf{P}}}_{t-1},\mathbf{Q})\leq\varepsilon_{t-1}\right\},$ and the solution is fixed at $x_{t}$. If the sequence of $\varepsilon_{t}$ decreases following clause 1 of Assumption 3.3, we have ${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}\Delta_{t}=0\right\}=1$.
This lemma formalizes the discrepancy between nominal DRO objectives with and without explicit support constraints. The discrepancy appears in the following bound, but we note that it converges to 0 as $\varepsilon_{t}\rightarrow 0$, and is often negligible.
Theorem 3.5 (Finite sample guarantee).
Suppose assumptions 3.3 and 3.4 hold. Then, the optimal solution $x^{K}_{t}$ and the optimal value $H^{K}_{t}$ of the compressed DRO problem (6) at time $t$ satisfies
$-\underline{\psi}_{t}^{K}\leq H_{t}^{K}-H_{t}\leq\bar{\psi}_{t}^{K},$ (12)
where $H_{t}$ is the nominal DRO value, and the bounds are defined as
$\underline{\psi}_{t}^{K}$ $=\min\left\{\Phi^{K}_{t-1},\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+d^{K}_{t-1,1})\right\},$
$\bar{\psi}_{t}^{K}$ $=\min\left\{\Delta_{t}+{\max_{j\leq J}(L_{j}/2)}(D^{K}_{t-1,2})^{2},\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+d^{K}_{t-1,1})\right\},$
with $\Phi^{K}_{t}$ given in Definition 3.4 and $\Delta_{t}$ given in Lemma 3.2. Furthermore, we obtain the finite-sample probabilistic guarantee
${\mathbf{P}}^{n_{t-1}}\left(H_{\star}\leq\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]\leq H^{K}_{t}+\underline{\psi}_{t}^{K}\right)\geq 1-\beta_{t-1},$ (13)
where the certificate $H^{K}_{t}+\underline{\psi}_{t}^{K}$ satisfies
${\mathbf{P}}^{n_{t-1}}\left(H_{\star}\leq H_{t}\leq H^{K}_{t}+\underline{\psi}_{t}^{K}\leq H_{t}+\underline{\psi}_{t}^{K}+\bar{\psi}_{t}^{K}\right)\geq 1-\beta_{t-1}.$ (14)
If either the Lipschitz condition or smoothness condition from Assumption 3.4 does not hold, the corresponding term(s) in the bounds are set to $\infty$.
Theorem 3.5 shows that, when the adaptive ambiguity set $\mathcal{P}^{K}_{t-1}$ is constructed with the same radius as nominal DRO, we can relate their optimal values in terms of the distances given in Definition 3.4. We can thus obtain finite-sample guaranteess, with a certificate within a bounded distance from the nominal DRO certificate.
In particular, equation (13) states that with probability $1-\beta_{t-1}$, the out-of-sample performance $\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]$ upper bounds the true optimal value $H_{\star}$, and this value is in turn upper bounded by the in-sample objective $H^{K}_{t}$ and a function $\underline{\psi}_{t}^{K}$ of the clustering discrepancy. This value $H^{K}_{t}+\underline{\psi}_{t}^{K}$ is our certificate.
Using the bounds established in (12), equation (14) then relates this certificate to the one obtained by nominal DRO, providing upper and lower bounds. Specifically, the certificate is at least an upper bound on $H_{t}$, but the increase in suboptimality (as we are solving a minimization problem, a larger certificate is less optimal)is at most $\underline{\psi}_{t}^{K}+\bar{\psi}_{t}^{K}$, which is once again controlled by the clustering discrepancy.
Depending on the smoothness and Lipschitz conditions on the function $f$, we provide multiple bounds for $\underline{\psi}_{t}^{K}$ and $\bar{\psi}_{t}^{K}$. For each application, we choose the best out of the available bounds. Note that in special cases, the terms $\Phi_{t-1}^{K}$ and $\Delta_{t}+{\max_{j\leq J}(L_{j}/2)}(D^{K}_{t-1,2})^{2}$ may reduce to 0. Specifically, for DRO problems with support $S={\mbox{\bf R}}^{d}$ (therefore $\Delta_{t}=0$), both values are 0 when the objective function $f$ is affine; we thus recover nominal DRO guarantees regardless of $K$. If the objective function is instead maximum-of-affine, the term involving $\max_{j\leq J}L_{j}$ is reduced to zero. On the other hand, if the objective $f$ is not affine, but we have $J=1$, then $\Phi^{K}_{t-1}=0$, as noted in Definition 3.4.
When the curvature of $f$ is not covered by the special cases above, the clustering discrepancies need to be explicitly computed. When the smoothness condition is satisfied, the term $\Delta_{t}+{\max_{j\leq J}(L_{j}/2)}(D^{K}_{t-1,2})^{2}$ can be easily computed using the clustered and non-clustered empirical distributions, and by approximating $\Delta_{t}\simeq 0$. As shown in Lemma 3.2, for ambiguity sets with small radii, this is often the case. When the Lipchitz condition is satisfied, we can compute the bound involving $d^{K}_{t-1,1}$. This is the Wasserstein-1 distance between two empirical distributions, and can be calculated using a linear program. Lastly, the value $\Phi^{K}_{t}$ can be computed regardless of smoothness and Lipschitz conditions, using the dual variables of the solved optimization problem. Therefore, all values in these bounds are readily assessable. In the case of bounded support, we can also use Proposition 3.2 to obtain upper bounds in terms of $\rho$.
Asymptotically, we note that these bounds converge to the following limits.
Theorem 3.6 (Asymptotic guarantee).
Suppose assumptions 3.23.3, and 3.4 hold. Asymptotically, the sequence of optimal solutions $x^{K}_{t}$ and optimal values $H^{K}_{t}$ of the compressed DRO problem (6) satisfies
$\mathbf{P}^{\infty}\left\{H_{\star}\leq\limsup_{t\rightarrow\infty}\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]\ \leq\limsup_{t\rightarrow\infty}H^{K}_{t}+\underline{\psi}_{\star}^{K}\leq H_{\star}+\underline{\psi}_{\star}^{K}+\bar{\psi}_{\star}^{K}\right\}=1,$
where
$\underline{\psi}_{\star}^{K}$ $=\min\left\{\limsup_{t\rightarrow\infty}{\Phi^{K}_{t}},\max_{j\leq J}M_{j}W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\right\},$
$\bar{\psi}_{\star}^{K}$ $=\min\left\{{\max_{j\leq J}(L_{j}/2)}(D^{K}_{\star,2})^{2},\max_{j\leq J}M_{j}W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\right\}.$
Proof.
This follows from the summability of $\beta_{t}$ and the Borel–Cantelli Lemma [31, Theorem 2.18], as well as Lemma 3.2. The convergence of the bounds follow from the results in Section 3.3. ∎
In order to obtain a closed-form limit for ${\Phi^{K}_{t}}$, we require the convergence of the online algorithm. Below, we prove that the online optimal value converges to a clustering-dependent value, at the same rate of convergence as nominal DRO.
Theorem 3.7.
Suppose Assumption 3.2 and the first clause of Assumption 3.4 holds, and let $\varepsilon^{K}_{t-1}\sim O(\varepsilon_{t-1})$, where $\varepsilon_{t-1}$ is computed as in the first clause of Assumption 3.3. With this radius, the optimal value $H^{K}_{t}$ of the compressed DRO problem (6) at time $t\geq\tau$ satisfies
${\mathbf{P}}^{n_{t-1}}\left(H^{K}_{\star}\leq{\mbox{\bf E}}_{{\mathbf{P}}^{K}_{\star}}[f(u,x^{K}_{t})]\leq H^{K}_{t}\right)\geq 1-\beta_{t-1},$
and asymptotically, ${\mathbf{P}}^{\infty}$-almost surely we have $H^{K}_{t}\downarrow H^{K}_{\star}$ as $t\to\infty$, where $H^{K}_{\star}$ is the optimal value of the stochastic optimization problem (1) with distribution ${\mathbf{P}}^{K}_{\star}$. In addition, if $f(u,x)$ is lower-semicontinuous in $x$ for all $u\in S$, then any accumulation point $\hat{x}^{K}_{\star}=\lim_{t\to\infty}x_{t}^{K}$ is almost surely, with respect to ${\mathbf{P}}^{\infty}$, an optimal solution to the stochastic optimization problem with distribution ${\mathbf{P}}^{K}_{\star}$.
Proof.
This follows from Theorem 3.4 and an adaptation of [40, Theorem 3.6], by treating ${\mathbf{P}}^{K}_{\star}$ as the true distribution. ∎
With this result, for special cases, we can obtain a limit for $\Phi^{K}_{t}$.
Theorem 3.8.
Suppose the assumptions of Theorem 3.7 hold. When the dual variables $z^{jk}$ of the reformulated problem (8) can be expressed as a continuous function of only the primal variables $x\in\mathcal{X}$, we have $\lim_{t\to\infty}\Phi^{K}_{t}=\Phi^{K}_{\star},$ where the latter is defined in Theorem 3.3.
Proof.
This follows from Theorem 3.7, where we note that the accumulation points $\hat{z}^{jk}_{\star}$ are defined by the accumulation point $\hat{x}^{K}_{\star}$. ∎
For maximum-of-affine $f$ without additional support constraints, the above is always satisfied. Well-known conjugation results show that the variables $z^{jk}=P^{jk}x$, for constant matrices $P^{jk}$ [40, Corollary 5.1].
Under optimal clustering, we obtain the following tighter bounds.
Theorem 3.9 (Optimal clustering bounds).
Suppose the assumptions of Theorem 3.6 hold. If the clustering algorithm is optimal with respect to $p=2$ and the $\ell_{2}$-norm, i.e., for all $t$, we achieve the minimum quantization error of $\hat{{\mathbf{P}}}_{t}$ on $K$ clusters, such that $D^{K}_{t,2}=d^{K}_{t,2}=d^{K}_{\star,2}(\hat{{\mathbf{P}}}_{t}),$ then almost surely with respect to ${\mathbf{P}}^{\infty}$,
$\bar{\psi}_{\star}^{K}=\min\left\{{\max_{j\leq J}(L_{j}/2)}(d^{K}_{\star,2}({\mathbf{P}}))^{2},\max_{j\leq J}M_{j}W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\right\}.$
On the other hand, suppose the assumptions of Theorem 3.8 hold, and $\|z^{jk}_{t}\|\leq M_{j}$ for $j=1,\dots,J$, $k=1,\dots,K$, and $t\geq 1$. If the clustering-induced coupling is optimal with respect to $p=1$, i.e., $D^{K}_{t,1}=d^{K}_{t,1},$ then almost surely with respect to ${\mathbf{P}}^{\infty}$,
$\underline{\psi}_{\star}^{K}$ $={\Phi^{K}_{\star}}\leq\max_{j\leq J}M_{j}W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star}).$
Proof.
The first result follows from Proposition 3.2 and Theorem 3.2, where the pertinent inequalities are made tight by the assumption of optimal clustering. For the second result, note that $\Phi^{K}_{t}$ is computed with respect to the clustering-induced coupling, i.e., each $\hat{u}$ is associated with the centroid $\bar{u}^{k}_{t}$ of the cluster it belongs to. If this is equivalent to the optimal coupling, then by the Cauchy-Schwarz inequality,
$(-z^{jk}_{t})^{T}(\hat{u}-\bar{u}^{k}_{t})\leq\|z^{jk}_{t}\|\|\hat{u}-\bar{u}^{k}_{t}\|\leq M_{j}\|\hat{u}-\bar{u}^{k}_{t}\|=M_{j}\min\left\{\|\hat{u}-\bar{u}\|\mid\bar{u}\in\{\bar{u}^{k}_{t}\}_{k=1}^{K}\right\},$
which holds for all $t$, $j$, and $\hat{u}$. This implies
$\Phi^{K}_{t}\leq\max_{j\leq J}M_{j}\frac{1}{n_{t}}\sum_{k=1}^{K}\sum_{\hat{u}\in C^{k}_{t}}\|\hat{u}-\bar{u}^{k}_{t}\|=\max_{j\leq J}M_{j}d^{K}_{t,1}.$
The result follows by taking the limit $t\to\infty$. ∎
Note that, the requirement on $z^{jk}_{t}$ is once again satisfied by maximum-of-affine $f$ without extra support constraints. Furthermore, while we only prove that $\Phi^{K}_{t}\leq\max_{j\leq J}M_{j}d^{K}_{t,1}$ under the conditions above, the large gain from the Cauchy-Schwarz inequality generally implies that this relationship, even without optimal coupling.
In all cases, our bounds scale with the clustering quality; when the clustering algorithm is optimized, the bound is tighter. Asymptotically, the performance-gap between the online solution and the nominal DRO solution is controlled by different distance metrics between the true distribution and its discretization to $K$ atoms, and, in the case of Lipschitz $f$, is upper bounded by the Wasserstein-1 distance. Theorem 3.7 shows a direct correlation between the optimality of the clustering procedure and the optimality of the online solution; the algorithm will converge to some value dependent on the supports of the final clustering, so it follows that a well-chosen clustering procedure will lead to a better performing solution.
For all these bounds, we can verify using either Proposition 3.2 or Theorem 3.2 that as the number of clusters $K\to\infty$, we recover the DRO asymptotic performance guarantee: as ${t\to\infty}$, we have ${\mathbf{P}}^{\infty}$-almost surely $H^{K}_{t}\rightarrow H_{\star}$. This illustrates the importance of finding a $K<\infty$ which balances the tradeoff between performance and computational effort. In fact, this leads to an interesting result: if we do not have a strict budget $K$, but would like to decrease computational effort, we can obtain the following more controllable bounds.
Theorem 3.10 (Fixed radii clustering).
Suppose assumptions 3.3 and 3.4 hold. If we fix a radius $\eta$ such that all data-points are either clustered into an existing ball with at most radius $\eta$, or initialized as a new cluster, then ${\mathbf{P}}^{\infty}$-almost surely,
$H_{\star}-2\max_{j\leq J}M_{j}\eta\leq\limsup_{t\rightarrow\infty}H^{K_{t}}_{t}\leq H_{\star}+2\min\left\{{\max_{j\leq J}L_{j}}\eta^{2},\max_{j\leq J}M_{j}\eta\right\}.$
Proof.
Under this clustering regime, the distance $W_{p}(\hat{{\mathbf{P}}}_{t},\hat{{\mathbf{P}}}^{K_{t}}_{t})\leq 2\eta$, where $K_{t}$ is allowed to increase with $t$. The result then follows from Theorem 3.5. ∎
4 Online performance analysis
So far, we have obtained guarantees with respect to the out-of-sample performance, and to the nominal DRO performance with the same information, i.e., both having observed $\mathcal{D}_{t-1}$. However, in online settings, we are interested in robustifying our solution with respect to the subsequent realization of the uncertainty, i.e., we want to compare with an oracle that has observed $\mathcal{D}_{t}$. This notion of regret is common in online learning [3, 44, 9], where the online solution is compared against the best solution in hindsight, and measures how well an algorithm can adapt to anticipated data.
In this section, we thus compute the dynamic regret bound, where we compare our online solution with information up to time $t-1$, against the nominal DRO solution with information up to time $t$. In particular, recall that we solve for the online solution at time $t$ before the new data-point is observed, with the ambiguity set $\mathcal{P}^{K}_{t-1}$. The dynamic regret, defined in equation (9), then calculates the difference between the worst-case performance of our online solution $x^{K}_{t}$, hedged against the updated non-compressed ambiguity set $\mathcal{P}_{t}$, and the optimal DRO value over the same set. Note that the performance of the online solution over the set $\mathcal{P}_{t}$, which is built around the full empirical distribution at time $t$, will differ from its performance over the set $\mathcal{P}^{K}_{t-1}$, which it was optimized against. We present the following results; all proofs are delayed to Appendix A.2.
Lemma 4.1.
Suppose assumptions 3.3 and 3.4 hold. The dynamic regret satisfies
$R(T,K)\leq\frac{1}{T}\sum_{t=1}^{T}\left(\underline{\psi}_{t+1}^{K}+\bar{\psi}_{t}^{K}\right)$ (15)
$\quad+\frac{1}{T}\left(\max_{j\leq J}M_{j}\left(\varepsilon_{0}+\varepsilon_{T}+W_{1}(\hat{{\mathbf{P}}}^{0},\hat{{\mathbf{P}}}_{t})+\sum_{t=1}^{T}(2\varepsilon_{t-1}+W_{1}(\hat{{\mathbf{P}}}^{K}_{t},\hat{{\mathbf{P}}}^{K}_{t-1}))\right)\right),$
with $\underline{\psi}_{t}^{K}$ and $\bar{\psi}^{K}_{t}$ as defined in Theorem 3.1.
Note that there are two sources of discrepancy: one from clustering, and one from the difference in time step. The difference in time step results in an offset in the index of $\underline{\psi}_{t+1}^{K}$, as well as the addition of the second set of terms. However, these differences once again boil down to the discrepancy between the distributions involved, for which we have convergence results. We thus obtain the following bounds.
Theorem 4.1 (Dynamic regret bound).
Suppose Assumption 3.3 and the first clause of Assumption 3.4 holds. For finite $T\gg\tau$, with probability $1-\zeta$, the dynamic regret is bounded as
$R(T,K)\leq O\left(\left(\log(\beta_{T}^{-1})/T\right)^{1/d}\right)+O\left(W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\right),$
where $\beta_{T}$ decays at least sublinearly in $T$, and $\zeta$ scales with $\beta_{T}$.
Corollary 4.1.1.
If we choose $\beta_{t}=O(\exp(-\sqrt{n_{t}}))$ in Assumption 3.3, the dynamic regret is bounded as
$R(T,K)\leq O\left(T^{-\frac{1}{2d}}\right)+O\left(W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\right),$
with some probability $1-\zeta\geq 1-\sum_{t=0}^{T}\beta_{t}$.
In Theorem 4.1, we assume the Lipschitz condition in order to obtain a clear convergence rate; the regret converges sublinearly to a value at most $O(W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star}))$, a factor of the Wasserstein-1 distance between the true distribution and the converged compressed distribution. This is in line with our expectations. Recall that over time, we gain confidence in the empirical measures as approximations of the true distributions. Therefore, both the ambiguity set radii we select and the theoretical distances between the empirical and true distributions converge sublinearly to 0. What remains is then a function of the discrepancy between the converged (true) distributions.
Theorem 4.1 thus consolidates theorems 3.53.6 and 3.7, and shows that the difference arising from the one-time-step information discrepancy converges to 0. Overall, only the choice of $K$ and the clustering algorithm directly affect the asymptotic performance of the online algorithm: they fully control the trade-off between efficiency and optimality. Below, we note this asymptotic behavior, which hold due to ${\mathbf{P}}^{\infty}$-almost sure convergence.
Corollary 4.1.2.
Suppose assumptions 3.3 and 3.4 hold. With probability 1, asymptotically we observe $\lim_{T\to\infty}R(T,K)\leq O(\underline{\psi}_{\star}^{K}+\bar{\psi}_{\star}^{K}),$ and additionally as $K\to\infty$$\lim_{K\to\infty}\lim_{T\to\infty}R(T,K)=0.$
5 Special case: bounded support
In Section 3, we gave performance guarantees for light-tailed supports $S$, and set the ambiguity set radius $\varepsilon^{K}_{t}$ with the nominal DRO radius $\varepsilon_{t}$, calculated using Theorem 3.1. In this section, for the case of bounded support, we give an explicit formulation of the radius in terms of the diameter of the support $S$. We also derive performance guarantees using a different theoretical approach from that of Section 3: using measure concentration results, we instead increase the ambiguity set radius by a finite value. In this case, the online solution is guaranteed to be an upper bound on the DRO solution, regardless of the curvature of $f$.
5.1 Wasserstein radius with bounded support
We begin with calculating the ambiguity set radius for the case of bounded support. Suppose the following assumption holds true.
Assumption 5.1 (Compact support).
The true distribution $\mathbf{P}$ has compact support $S\subset\mathbf{R}^{d}$.
Then, the result follows from propositions 24 and 25 of the online version of [13], which makes use of the concentration bound below.
Proposition 5.1 (Concentration bound).
Let Assumption 5.1 hold, and let $p<d/2$. Then, for the true distribution ${\mathbf{P}}$ and an empirical distribution $\hat{{\mathbf{P}}}^{N}$ supported on $N$ points,
$\mathbf{P}^{N}\left(W_{p}\left(\hat{\mathbf{P}}^{N},\mathbf{P}\right)\geq\mathbf{E}\left[W_{p}\left(\hat{\mathbf{P}}^{N},\mathbf{P}\right)\right]+t\right)\leq e^{-Nt^{2p}/\left(2\tilde{\rho}^{2p}\right)}\quad\forall t\geq 0,$
where $\tilde{\rho}=\mathrm{diam}_{2}(S)$.
This follows from [11, Proposition A.2] and exploits the fact that bounded distributions are sub-Gaussians. Then, by bounding the expected Wasserstein distance using a function of $d,p,\tilde{\rho}$, and $N$, we obtain the following explicit characterization of the nominal ambiguity radius. For details, see the proofs in the online version of the referenced paper.
Theorem 5.1 (Explicit nominal radius [13][Propositions 23, 24]).
Assume that the probability measure $\mathbf{P}$ is supported on $S\subset\mathbf{R}^{d}$ with $\rho=\frac{1}{2}\operatorname{diam}_{\infty}(S)<\infty$ and that $p<d/2$. Then, for a given confidence $\beta\in(0,1)$, the nominal ambiguity radius is given by
$\varepsilon_{N}(\beta,\rho)=2\rho\left(CN^{-\frac{1}{d}}+\sqrt{d}\left(2\ln\beta^{-1}\right)^{\frac{1}{2p}}N^{-\frac{1}{2p}}\right),$
where
$C=\sqrt{d}2^{(d-2)/(2p)}\left(\frac{1}{1-2^{p-d/2}}+\frac{1}{1-2^{-p}}\right)^{1/p}.$
Furthermore, the radius can be written with the compact form
$\varepsilon_{N}(\beta,\rho)=2\rho\left(\frac{\ln\left(C^{\star}\beta^{-1}\right)}{c^{\star}}\right)^{\frac{1}{d}}N^{-\frac{1}{d}},$
with $C^{\star}=C^{d}/(2\sqrt{d}^{d})$ and $c^{\star}=1/(2^{d}\sqrt{d}^{d})$.
For any time step $t$ and $\beta_{t}$, the radius $\varepsilon_{t}$ calculated from Theorem 5.1 can be used to obtain the performance guarantees in Section 3, since distributions with bounded support are trivially light-tailed.
On the other hand, we can also use this radius to obtain performance guarantees specifically for bounded distributions. In the following, we show that by suitably enlarging the radius $\varepsilon_{t}$, we can also obtain performance guarantees for the online compressed DRO problem.
5.2 Performance guarantees with bounded support
In this approach, we assume the clustering algorithm to satisfy the requirements in Section 3.2, and additionally restrict the diameter of each cluster to a value $2\eta_{K}$. In Section 6.1, we give such an algorithm, which is based on covering the bounded support with $K$ balls of finite radius $\eta_{K}$. By setting the ambiguity set radius $\varepsilon^{K}_{t}$ as the nominal radius $\varepsilon_{t}$ plus a function of $\eta_{K}$, we ensure that the true distribution lies within the adaptive ambiguity set with high probability. This then allows us to derive guarantees on the out-of-sample performance of the online solution $x^{K}_{t}$. To show this, we first quantify the discrepancy between the distributions of interest. Note that the proofs of this section are delayed to Appendix A.3.
Lemma 5.1 (Distance between distributions).
Let $\hat{{\mathbf{P}}}^{K}_{t}$ be the clustered empirical distribution at time $t$, with cluster diameters at most $2\eta_{K}$. For all $t$, we have $d_{W_{p}}(K,t)\leq 2\eta_{K}.$ With probability $1-\beta_{t}$, we have $W_{p}(\mathbf{P},\hat{\mathbf{P}}^{K}_{t})\leq\varepsilon_{t}+2\eta_{K},$ where $\beta_{t}$ and $\varepsilon_{t}$ are computed as in Theorem 5.1.
Lemma 5.1 implies that by selecting the radius $\varepsilon^{K}_{t}=\varepsilon_{t}+2\eta_{K}$, where $\varepsilon_{t}$ is the nominal radius for Wasserstein DRO, we can obtain the same finite sample performance guarantee for the online algorithm as Wasserstein DRO. We formalize this result in the following theorem.
Theorem 5.2 (Finite sample performance guarantee).
Under the assumptions of Theorem 5.1, let the adaptive ambiguity set $\mathcal{P}^{K}_{t-1}$ be defined with $\varepsilon^{K}_{t-1}=\varepsilon_{t-1}+2\eta_{K}$. Then, the solution $x^{K}_{t}$ and optimal value $H^{K}_{t}$ of the compressed DRO problem (6) implies the finite sample performance guarantee
${\mathbf{P}}^{n_{t-1}}\left(H_{\star}\leq{\mbox{\bf E}}_{{\mathbf{P}}}[f(u,x^{K}_{t})]\leq H^{K}_{t}\right)\geq 1-\beta_{t-1}.$
Proof.
This follows from Lemma 5.1 and the same logic as nominal DRO results [40, Theorem 3.5]. ∎
In addition, we can derive the following asymptotic result by applying a similar approach as [40, Lemma 3.7].
Lemma 5.2 (Convergence of distributions).
Assume for all $t$, $\varepsilon_{t}$ is computed as in Theorem 5.1, with a sequence $\beta_{t}$ such that $\sum_{t=0}^{\infty}\beta_{t}<\infty$ and $\lim_{t\to\infty}\varepsilon_{t}=0$. Then, any sequence $\mathbf{Q}_{t}\in\mathcal{P}^{K}_{t}$, where $\mathcal{P}^{K}_{t}$ has radius $\varepsilon^{K}_{t}=\varepsilon_{t}+2\eta_{K}$, satisfies
$\mathbf{P}^{\infty}\left\{\lim_{t\rightarrow\infty}W_{p}(\mathbf{P},\mathbf{Q}_{t})\leq 4\eta_{K}\right\}=1.$
This lemma dictates that asymptotically, the true distribution lies in a ball around the clustered distribution with radius $4\eta_{K}$; the increased radius is the price to pay due to the limited computational budget available to solve (4), and is directly related to the quality of our approximation $\hat{{\mathbf{P}}}^{K}_{t}$. Clearly, if $K$ is high, then $\eta_{K}$ will be low and our asymptotic estimate will be more accurate; viceversa for low $K$. Following [40, Theorem 3.6], we now derive the following result.
Theorem 5.3 (Asymptotic performance guarantee).
Assume for all $t$, $\varepsilon_{t}$ is computed as in Theorem 5.1, with a sequence $\beta_{t}$ such that $\sum_{t=0}^{\infty}\beta_{t}<\infty$ and $\lim_{t\to\infty}\varepsilon_{t}=0$. Also assume the adaptive ambiguity set $\mathcal{P}^{K}_{t}$ is defined with $\varepsilon^{K}_{t}=\varepsilon_{t}+2\eta_{K}$, and that for all $x\in\mathcal{X}$, the constituent functions $f_{j}$ of $f$ satisfy the $M$-Lipschitz condition given in Definition 2.1 with constants $M_{j}$. Then, the sequence of solutions and optimal values $x^{K}_{t}$ and $H^{K}_{t}$ of the compressed DRO problem (6) satisfies
$\mathbf{P}^{\infty}\left\{H_{\star}\leq\limsup_{t\rightarrow\infty}{\mbox{\bf E}}_{\mathbf{P}}[f(u,x^{K}_{t})]\leq\limsup_{t\rightarrow\infty}H^{K}_{t}\leq H_{\star}+4\max_{j\leq J}M_{j}\eta_{K}\right\}=1.$
Theorem 5.3 shows that, asymptotically, our online in-sample value is an upper bound on the out-of-sample and true values, but is also no more than $4\max_{j\leq J}M_{j}\eta_{K}$ suboptimal compared to the true solution. In other words, there exists some value $\tau$ such that for all $t\geq\tau$, the online solution $H^{K}_{t}\leq H_{\star}+4\max_{j\leq J}M_{j}\eta_{K}$. We note that the upper bound on $H^{K}_{t}$ may be less tight than the ones derived in Section 3.4, but $H^{K}_{t}$ is now an upper bound on $H_{\star}$, regardless of the curvature of $f$, without the need for the term $\Phi^{K}_{t}$. Therefore, for maximum-of-concave $f$, this bound may perform better than the ones in Section 3.4, depending on the specific robust optimization problem.
Overall, to tune the tradeoff between performance and computational effort, we would again be inclined to choose a number $K$ that provides a reasonable $\eta_{K}$ while not being too computationally expensive.
6 Online algorithms for data-compression
In Section 3.2, we gave requirements for the clustering algorithm. In this section, we provide three online clustering algorithms satisfying the given assumptions. These algorithms are expected to maintain a reasonably accurate approximation of the optimal discretization while minimizing computational effort. Below, we summarize their time and memory complexities.
Table 1: Time and memory complexities of our clustering algorithms. Recall that $n_{t}$ is the total number of data-points at time $t$, $K$ is the number of clusters, and $d$ is the dimension of the data. We also let $I$ be the number of iterations required for the clustering algorithm’s convergence. The values $G$ and $Q$ are defined in the respective algorithms.
Algorithm Time (initialization) Time (online) Memory
SupCover $O(GK)$ $O(Kd)$ $O(Kd)$
Reclustering $O(In_{0}Kd)$ $O(In_{t}Kd)$ $O((K+n_{t})d)$
OnlineClustering $O(In_{0}Qd)$ $O(IQKd)$ $O((K+Q)d)$
The most time and memory efficient algorithm, as we will describe below, is the SupCover algorithm, but it is the least flexible in terms of cluster assignments, and is only for bounded supports. The Reclustering approach gives the best approximation, but has the highest time and memory complexities. The OnlineClustering approach is suboptimal compared to reclustering, but has lower runtime and memory complexities; it doesn’t need to store the observed dataset.
6.1 A bounded coverage of the support (SupCover)
We begin with an intuitive approach for a bounded support set $S$: partitioning $S$. When $S$ is bounded, we can always cover $S$ with $K$ balls $B_{\eta_{K}}(a^{k})$ with centers $A=\{a^{k}\}_{k=1}^{K}\subseteq S$ and fixed finite radius $\eta_{K}>0$. Specifically, the centers and the radius are selected as follows.
Assumption 6.1.
The parameters $\{a^{k}\}_{k=1}^{K}$ and $\eta_{K}$ are chosen such that, $\forall u\in S$
  1. 1.
    $\min_{k\leq K}\|u-a^{k}\|\leq\eta_{K}$;
  2. 2.
    $u\subseteq\cup_{k\in K}B_{\eta_{K}}(a^{k})$;
  3. 3.
    $\eta_{K}=\min\{\eta\mid\min_{k\leq K}\|u-a^{k}\|\leq\eta,~{}u\in S\}$.
In fact, such a partitioning is equivalent to an $\eta_{K}$-net of the support $S$, or a $k$-centers problem with order 1. Solving this problem is NP-hard, but we can use an efficient greedy-algorithm approximation [23]. In this greedy algorithm, we begin with a random point, and iteratively find the furthest point from the current point, and set these $K$ total points as the cluster centers. With a given point $a^{k}$, the process of finding the next furthest point depends on the support $S$, and can be solved using the following optimization problem
$\mathop{\rm argmax}_{u\in S}\|u-a^{k}\|,$
whose complexity depends on the shape of $S$. We let the time complexity of this operation be $G$, and note that it needs to be solved $K$ times. The radius $\eta_{K}$ can be set to the maximum of the distances found.
Once the centers are found, we can construct a simple procedure for updating $\hat{{\mathbf{P}}}^{K}_{t}$. The entire process is summarized in Algorithm 1.
Algorithm 1 SupCover
1:given $\mathcal{D}_{0},K,S$
2:choose centers $\{a^{k}\}_{k=1}^{K}$ and radius $\eta_{K}$ satisfying Assumption 6.1
3:assign data-points to clusters $\{C^{k}_{0}\}_{k=1}^{K}$, and compute $\{n^{k}_{0}\}_{k=1}^{K}$, $\{\theta^{k}_{0}\}_{k=1}^{K}$, $\{\bar{u}_{0}^{k}\}_{k=1}^{K}$
4:for $t=0,1,\dots$ do
5:     $\hat{\mathbf{P}}^{K}_{t}\leftarrow\sum_{k=1}^{K}\theta_{t}^{k}\delta_{\bar{u}_{t}^{k}}$
6:     $C_{t}^{k^{\prime}}\leftarrow C_{t}^{k^{\prime}}\cup\{\hat{u}\}$triangleright assign $\hat{u}$ to cluster $k^{\prime}$ following  (16)
7:     $\theta^{k}_{t}\leftarrow$ (17) $\triangleright$ update weights
8:     $\bar{u}^{k^{\prime}}_{t}\leftarrow(n^{k^{\prime}}_{t-1}\bar{u}^{k^{\prime}}_{t-1}+\hat{u})/n^{k^{\prime}}_{t}$triangleright update centroid
Specifically, the cluster support set $\mathcal{S}_{t}$ can be seen as the Voronoi diagram of $A$, which covers the entire support $S$, and remains constant for all $t$. In addition, each Voronoi region $S^{k}$ has a diameter bounded by $\eta_{K}$. The clustering induced by this support is then
$\hat{u}\in C_{t}^{k^{\prime}}\quad\text{if}\;\hat{u}\in S^{k^{\prime}},\quad\text{where}\quad k^{\prime}=\mathop{\rm argmin}_{k\leq K}\|\hat{u}-a^{k}\|,$ (16)
and $\|\hat{u}-\zeta_{k^{\prime}}\|\leq\eta_{K}$. The set of weights $\{\theta^{k}_{t}\}_{k=1}^{K}$ follow the first-order dynamics
$\theta^{k}_{t}=\begin{cases}\frac{n^{k}_{t-1}+1}{n_{t}}&\text{if}\>\hat{u}\in C^{k}_{t}\\ \frac{n^{k}_{t-1}}{n_{t}}&\text{otherwise}\end{cases}\quad k=1,\dots,K,$ (17)
where $\hat{u}$ is the data-point at time $t$. Additionally at each time $t$, the centroid of the cluster that gained a data-point is updated. Notably, we do not need to store $\mathcal{D}_{t}$. By keeping track of the centers, centroids and weights of each cluster, together with a count of the total number of data-points seen, we can perform the above updates with only the newest data-point. After performing the update, this data-point can also be discarded. The algorithm is thus memory efficient; it only requires $O(Kd)$ space.
However, while the updating procedure of this algorithm is simple, its performance is highly dependent on the initial selection of centers $\{a^{k}\}_{k=1}^{K}$. If the covering found is a suboptimal representation of the data, i.e., a single cluster contains multiple distinct modes of the data, then the performance is expected to be suboptimal as well.
6.2 Approximate $k$-centers clustering with warm starts (Reclustering)
To track the given data more closely, we can recompute the approximate $k$-centers clustering at each time step, until the stopping threshold $\tau$. At each time $t$, the cluster support for each $k$ is the Voronoi region around the center $a_{t}^{k}\in A_{t}$. After the threshold $\tau$, we cluster points using the fixed centers $A_{\tau}$, similarly as above. At times $t\leq\tau$, the subsequent clustering can be warm-started with the previous set of centers. We assume without loss of generality that the number of clusters remain $K$, and summarize the algorithm in Algorithm 2.
Algorithm 2 Reclustering
1:given $\mathcal{D}_{0},K,\tau$
2:$A_{0}\leftarrow$ find $K$ centers using $k$-centers on $\mathcal{D}_{0}$
3:assign data-points to clusters $\{C^{k}_{0}\}_{k=1}^{K}$, and compute $\{n^{k}_{0}\}_{k=1}^{K}$, $\{\theta^{k}_{0}\}_{k=1}^{K}$, $\{\bar{u}_{0}^{k}\}_{k=1}^{K}$
4:for $t=0,1,\dots$ do
5:     $\hat{\mathbf{P}}^{K}_{t}\leftarrow\sum_{k=1}^{K}\theta_{t}^{k}\delta_{\bar{u}_{t}^{k}}$
6:     observe data-point $\hat{u}$
7:     if $t\geq\tau$ then
8:         $C_{t}^{k^{\prime}}\leftarrow C_{t}^{k^{\prime}}\cup\{\hat{u}\}$triangleright assign $\hat{u}$ to cluster $k^{\prime}$ following  (16)
9:         $\bar{u}^{k^{\prime}}_{t}\leftarrow(n^{k^{\prime}}_{t-1}\bar{u}^{k^{\prime}}_{t-1}+\hat{u})/n^{k^{\prime}}_{t}$triangleright update centroid
10:         $\{\theta^{k}_{t}\}_{k=1}^{K}\leftarrow$ (17) $\triangleright$ update weights
11:     else  repeat lines 2 and 3, warm starting at $A_{t-1}$
In the complexity analysis, we assume the $k$-centers approximation to be $k$-means [26], with an iterative algorithm converging in $I$ steps. We note that this method is not memory conservative, as we need to store all data-points to recompute the cluster information. However, with warm-starts, it can be quite computationally efficient; while the theoretical upper bound on the time-complexity is unchanged compared to not warm-starting, in practice, warm-starting greatly reduces the required number of iterations for convergence.
In the complexity analysis, we assume the $k$-centers approximation to be $k$-means [26], with an iterative algorithm converging in $I$ steps. This method is not memory conservative, as we need to store all data-points to recompute the cluster information. However, with warm-starts, it can be quite computationally efficient; in practice, warm-starting greatly reduces the required number of iterations.
6.3 Online clustering (OnlineClustering)
For an algorithm that is also memory conservative, we propose an online updating procedure that takes inspiration from a well-known online clustering approach, CluStream [2]. We summarize the algorithm here, and give a more detailed version in Appendix A.4. Similarly to CluStream, we maintain two sets of clusters: a set of $Q\geq K$ microclusters, and a set of $K$ macroclusters. Data-points are assigned to microclusters on arrival, or initialized as a new cluster based on its distance to current clusters. When new microclusters are initialized, the closest microclusters are merged to maintain the budget $Q$. The $K$ macroclusters are obtained by clustering the microclusters, and can be warm-started.
Differently from CluStream, we allow for separate cluster centers and centroids; the centers do not shift with the addition of new data-points, while the centroids are updated. For times $t\leq\tau$, all clustering assignments can be made according to either the centers or the centroids. After $t>\tau$, cluster assignments must be made according to the centers of the $K$ macroclusters. We require the macrocluster update to also be performed online, so that they have up-to-date information. For all clusters, we formalize the notion of the cluster supports, such that they are consistent with Definition 3.1 and Assumption 3.2. This requires adjustments to the Voronoi regions around the centers/centroids, based on the already-clustered data-points. For continuous distributions ${\mathbf{P}}$, since the data-points have measure 0, this does not need to be explicitly maintained. Therefore, this algorithm can discard all data-points once seen, and keep only the aggregated information of size $O((K+Q)d)$. For discrete ${\mathbf{P}}$, however, the adjustment to the support is no longer negligible. Nevertheless, discrete distributions generally have a finite set of atoms, so memory is not a bottleneck.
For the complexity analysis, we assume the $k$-centers approximation on the $Q$ microclusters to be $k$-means [26]. Compared to the reclustering approach, we are solving a smaller problem, which reduces the theoretical time complexity; the set of $Q$ microclusters allows this method to interpolate between SupCover and Reclustering. Practically, however, the main advantage of this approach is the lower memory requirement.
7 Choosing hyper-parameter values
Above, we detailed various theoretical bounds for the online algorithm, with a given $K$ and with $\varepsilon^{K}_{t}$ chosen according to measure concentration results. In practice, however, these values are hyper-parameters that can be tuned. In this section, we give some guidelines for their practical selection.
Choosing $K$
As noted in the previous sections, the value of $K$ controls the discrepancy between the online and nominal DRO performance. A low value of $K$ is computationally efficient, while higher values allow for better approximations of the underlying distribution, thus increasing solution quality. In practice, we expect a value of $K$ that doesn’t far exceed the number of modes of the underlying distribution. If the initial dataset $\mathcal{D}_{0}$ is large enough, this value of $K$ can be chosen a-priori, using the elbow method on the clustering value $(D^{K}_{0,2})^{2}$ [49]. If the initial dataset is small, we can initialize with a small value of $K$, and allow it to increase or decrease, based on its performance. Our asymptotic analysis requires only that after a finite time $\tau$, the cluster supports are fixed; prior to this time, the clusters are flexible, and could be adjusted to better approximate the true distribution.
Choosing $\varepsilon^{K}_{t}$
In Wasserstein DRO, for a tighter practical certificate on the out-of-sample performance, the radius $\varepsilon$ is often chosen through cross-validation. We employ a similar technique for the compressed online algorithm, where we tune $\varepsilon^{K}_{t}$ to minimize the out-of-sample performance. We assume to have a validation dataset of size $N^{\rm val}$, in addition to our growing training data. For every time step $t$, upon constructing the clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$, we solve the optimization problem (8) with a finite set of radii $\varepsilon^{K}_{t}$. For each in-sample solution, we use the validation dataset to estimate the out-of-sample performance via SAA, and record whether or not it is upper bounded by the certificate $H^{K}_{t}+\underline{\psi}_{t}^{K}$, defined in Theorem 3.5. We select the sequence of $\varepsilon^{K}_{t}$ that gives the lowest out-of-sample performance, and report the corresponding sequence of data-driven solutions, certificates, and empirical confidence.
8 Numerical examples
We now illustrate the computational performance and robustness of the proposed method on a numerical example. All the code to reproduce our experiments is available, in Python, at
https://github.com/stellatogrp/online_mro.
We run the experiments on the Princeton Institute for Computational Science and Engineering (PICSciE) facility with 35 parallel 2.4 GHz Skylake cores. We solve all optimization problems with the MOSEK [41] optimizer with default settings.
Baselines and metrics
We compare four different approaches, described below.
  • Online clustering (our memory-efficient method). We use our online clustering algorithm, as described in Algorithm 3, to update cluster assignments. This method discards the data-points once seen, and is thus memory-efficient. Upon constructing the empirical clustered distribution, we solve (8) for the online solution.
  • $k$-means reclustering (our method). Same as above, except using $k$-means clustering at each time step, warm-starting at the previous centers. The procedure is described in Algorithm 2.
  • Wasserstein DRO. At time $t$, we solve (8) with the ambiguity set $\mathcal{P}_{t-1}$.
  • Sample average approximation SAA. At time $t$, we solve the stochastic optimization problem with respect to the empirical measure $\hat{{\mathbf{P}}}_{t-1}$ of the training data. We denote this in-sample value as $H^{\rm SAA}_{t}$.
For all approaches, we use the same training (in-sample) and testing (out-of-sample) datasets. The testing dataset is of size 200, and we consider up to $T=2000$ time steps. For the DRO methods, we select the radii $\varepsilon_{t}$ using the procedure given in Section 7. Specifically, we choose the best $\varepsilon_{t}$ determined through 30 repetitions of the experiment. We compare the following metrics, with values averaged over all repetitions. We additionally show the 25-75th percentiles using shaded regions.
  • In-sample certificate. For all approaches, we compare the in-sample certificates obtained with respect to the training data. For nominal Wasserstein DRO, the certificate is the in-sample objective $H_{t}$. For our online methods, the certificate is the in-sample objective plus the clustering discrepancy, $H^{K}_{t}+\underline{\psi}_{t}^{K}$.
  • Empirical confidence. The empirical confidence is computed as the probability that the in-sample certificate upper bounds the out-of-sample expected value, which is the stochastic optimization objective computed using the empirical measure of the testing dataset, at the given solutions. This value is denoted $1-\hat{\beta}_{t}$.
  • Computation times. We compare the per-iteration computation times, which include both clustering and solving times for our data-compressed methods.
  • Regret. For our methods, we compute the dynamic regret $R(T,K)$ (9), compared with the theoretical upper bound given in Lemma 4.1.
8.1 Sparse portfolio optimization
Similarly as in [56, 36], we consider a market that forbids short-selling and has $d$ assets. The uncertain parameters are the daily returns of these assets, given by the random vector $r=(r_{1},\dots,r_{d})\in{\mbox{\bf R}}^{d}$. The percentage weights (of the total capital) invested in the assets are given by the decision vector ${x=(x_{1},\dots,x_{n})\in{\mbox{\bf R}}^{n}}$. We restrict our selection to at most $\gamma$ assets, given by the 00-th norm cardinality constraint below. The distribution ${\mathbf{P}}$ is unknown, but we have access to a streaming dataset $\mathcal{D}_{t}$, updated each day with a new returns vector. Our objective is to minimize the CVaR with respect to variable $x$,
minimize $\begin{array}[]{ll}\text{minimize }&\mathop{\bf CVaR}(-r^{T}x,\alpha)\\ \text{subject to}&\mathbf{1}^{T}x=1,\quad x\geq 0,\quad\|x\|_{0}\leq\gamma,\end{array}$
[…]
Problem setup
We take stock data from the past 5 years of S&P500 (1/1/2020 to 1/1/2025) daily returns, and generate synthetic data from their fitted general Pareto distributions, with correlations preserved using a Gaussian copula. These distributions are chosen to model the data more accurately than a Gaussian fit. We let $\alpha=20\%$, $d=50$ stocks, and restrict our portfolio to at most $\gamma=8$ stocks. We initialize with a dataset of size $n_{0}=5$, and show results for $K=15$ and $K=25$. This range is obtained using the elbow method on the initial dataset, and allowing for adjustments.
Figure 1: Sparse portfolio, $K=25$. In-sample certificates, empirical confidence, and per-iteration computation times for the different methods, at $\varepsilon^{K}_{t}=0.0025{(t+n_{0})^{-1/40}}$.
Figure 2: Sparse portfolio. Left: $K=15$, right: $K=25$. Out-of-sample expected values for $\varepsilon_{t}=0.0025{(t+n_{0})^{-1/40}}$.
Results
For all DRO settings, the radius we obtain through cross-validation is $\varepsilon_{t}=0.0025{(t+n_{0})^{-1/40}}$. In Figure 1, we compare the certificates, empirical confidence, and per-iteration computation times at the chosen $\varepsilon_{t}$ sequence, and in Figure 2, compare the out-of-sample performance. We observe that our online data-compressed approaches introduce slight sub-optimality in out-of-sample performance, but provide high-confidence certificates, and offer significant speed-ups compared to nominal DRO. As expected, $K=25$ clusters out-performs $K=15$ clusters, but $K=15$ already gives a good approximation of the nominal DRO performance. The SAA approach improves in performance as the number data points increases, but does not offer a certificate of optimality: the empirical confidence is near 0 for all $t$. Furthermore, it also grows in complexity with sample size and is less efficient than our online approaches. In the long-term, the computation times for our data-compressed approaches can be multiple orders of magnitude faster than that of both SAA and nominal DRO. We find that the $k$-means warm-starting algorithm (reclustering approach) is particularly efficient, achieving lower computation times and better solution quality than the online clustering method, which entails an extra trade-off between optimality and memory efficiency. Nonetheless, the differences are minimal.
In Figure 3, we show the various certificates given in equation (14), and note that the relationships follow the hierarchy presented. The data-compressed in-sample objective value for this maximum-of-affine problem is lower than the in-sample objective of nominal DRO, but adding the clustering discrepancy makes it a valid certificate.
Figure 3: Sparse portfolio, $K=25$. The radius $\varepsilon_{t}=0.0025{(t+n_{0})^{-1/40}}$ for all methods. Left: certificates from (14). Right: dynamic regret (9), compared to the theoretical upper bounds.
In Figure 3, we also show the dynamic regret $R(T,K)$ and the regret bound as calculated in Lemma 4.1. As $T\to\infty$, the upper bound is expected to converge to a fixed function of the clustering discrepancy; indeed, we observe that it coincides with $\Phi^{K}_{t}$. This is in line with the theory; since we are solving a maximum-of-affine problem, the term $\bar{\psi}^{K}_{t}$ is reduced to 0. The limiting distance becomes $\Phi^{K}_{t}$, which, as we show in Figure 4 below, is a tighter bound than the Wasserstein-1 distance. This value should converge to its theoretical limit $\Phi_{\star}^{K}$.
Figure 4: Sparse portfolio, $K=25$, $T=10000$, at $\varepsilon_{t}=0.0025{(t+n_{0})^{-1/40}}$. Left: in-sample certificates. Right: clustering distances given in Definition 3.4.
In Figure 4, we show the results of our online approaches for a longer horizon, up to $T=10000$, with $\tau=8500$. We observe that the $p=1$ and $p=2$ clustering distances, $d^{K}_{t,1}$ and $(D^{K}_{t,2})^{2}$, as well as $\Phi^{K}_{t}$, are converging, and the out-of-sample performances are also converging in mean. As remarked above, we observe that $d^{K}_{t,1}$ upper bounds $\Phi^{K}_{t}$ without even multiplying by the maximum Lipschitz constant (which is $51/\alpha=$ 5). This illustrates the relative optimality of $\Phi^{K}_{t}$, as opposed to the Wasserstein-1 distance, as proven in Theorem 3.9. In fact, although we have not assumed (or achieved) optimal clustering-induced coupling, the result still held.
9 Conclusions
We have introduced an online data compression framework for solving Wasserstein DRO problems with streaming data. Our method constructs adaptive ambiguity sets using online clustering, allowing the uncertainty model to evolve as new data arrives, while maintaining out-of-sample performance guarantees. We analyzed the impact of data compression on solution quality, providing finite-sample and asymptotic performance bounds, as well as a sublinear regret analysis with respect to the full-information DRO solution. The framework is compatible with a broad class of clustering algorithms and supports efficient, memory-aware implementations. Numerical experiments in sparse portfolio optimization demonstrate significant computational savings with minimal loss in solution quality.
Acknowledgments
I. Wang and B. Stellato are supported by the NSF CAREER Award ECCS-2239771. I. Wang is also supported by the Princeton Wallace Memorial Fellowship. B. Stellato is also supported by the ONR YIP Award N000142512147. The authors are pleased to acknowledge that the work reported on in this paper was substantially performed using the Princeton Research Computing resources at Princeton University which is a consortium of groups led by the Princeton Institute for Computational Science and Engineering (PICSciE) and Research Computing.
Appendix A Appendices
A.1 Proof of the performance guarantees in Section 3.4
Proof of Lemma 3.2.
For all $t$, let $\tilde{\mathbf{Q}}_{t}\in\tilde{\mathbf{B}}_{\varepsilon_{t}}^{p}(\hat{{\mathbf{P}}}_{t})$ and $\mathbf{Q}_{t}\in\mathcal{P}_{t}$. By [40, Lemma 3.7], we observe ${\mathbf{P}}^{\infty}\{\lim_{t\to\infty}\tilde{\mathbf{Q}}_{t}\rightarrow{\mathbf{P}}\}=1$ and ${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}\mathbf{Q}_{t}\rightarrow{\mathbf{P}}\right\}=1$. The result follows. ∎
Proof of Theorem 3.5.
Recall that under the radius $\varepsilon_{t-1}$, the solution $x_{t}$ and optimal value $H_{t}=\max_{\mathbf{Q}\in\mathcal{P}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x_{t})]$ of the DRO problem (4) imply the finite sample performance guarantee (5). We then observe
$\max_{\mathbf{Q}\in\mathcal{P}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x_{t})]\leq\max_{\mathbf{Q}\in\mathcal{P}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]$ $\leq\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]+\Phi^{K}_{t-1},$
with $\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]=H^{K}_{t}$ in the final clause. The first inequality follows from the optimality of $x_{t}$, the second follows from [56, Theorem 5], and $\Phi^{K}_{t}$ is defined in Definition 3.4. From [56, Theorem 4], we note that $\Phi^{K}_{t-1}=0$ for concave functions $f$i.e., when $J=1$. We also observe, using Kantorovich-Rubinstein duality [33],
$\max_{\mathbf{Q}\in\mathcal{P}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]$ $\leq\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]+\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+W_{1}(\hat{{\mathbf{P}}}_{t-1},\hat{{\mathbf{P}}}^{K}_{t-1}))$
$\leq H^{K}_{t}+\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+d^{K}_{t-1,1}).$
Combining the above relations, we obtain
${\mathbf{P}}^{n_{t-1}}\left(H_{\star}\leq H_{t}\leq\hat{H}^{K}_{t}\leq H^{K}_{t}+\min\left\{\Phi^{K}_{t-1},\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+d^{K}_{t-1,1})\right\}\right)\geq 1-\beta_{t-1}.$
To derive (13), we also note that, with probability $1-\beta_{t-1}$,
$H_{\star}=\min_{x\in\mathcal{X}}\mathbf{E}_{\mathbf{P}}[f(u,x)]\leq\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]\leq\max_{\mathbf{Q}\in\mathcal{P}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]\leq H^{K}_{t}+\Phi^{K}_{t-1}.$
Now for the upper bound, from the optimality of $H^{K}_{t}$ and [56, Theorem 5],
$H^{K}_{t}\leq\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x_{t})]$ $\leq H_{t}+(\tilde{H}_{t}-H_{t})+{\max_{j\leq J}(L_{j}/2)}(D^{K}_{t-1,2})^{2}$
$=H_{t}+\Delta_{t}+{\max_{j\leq J}(L_{j}/2)}(D^{K}_{t-1,2})^{2},$
where $\Delta_{t}$ is given in Lemma 3.2. By applying Kantorovich-Rubinstein duality,
$H^{K}_{t}\leq\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x_{t})]\leq H_{t}+\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+d^{K}_{t-1,1}).$
Combining both inequalities, we obtain
$H^{K}_{t}\leq H_{t}+\min\left\{\Delta_{t}+{\max_{j\leq J}(L_{j}/2)}(D^{K}_{t-1,2})^{2},\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+d^{K}_{t-1,1})\right\}.$
Combining the upper and lower bounds gives the desired result.Rearranging terms gives equation (12). ∎
A.2 Proof of the regret bound in Section 4
Proof of Lemma 4.1.
For brevity, let us denote the following shorthands: $\bar{f}_{t}(x)=\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x)]$, $\bar{f}^{K}_{t}(x)=\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x)],$ and the corresponding solutions $x_{t}=\mathop{\rm argmin}_{x\in\mathcal{X}}\bar{f}_{t}(x)$, $x^{K}_{t}=\mathop{\rm argmin}_{x\in\mathcal{X}}\bar{f}^{K}_{t}(x).$ We wish to bound $R(T)=(1/T)\sum_{t=1}^{T}\left(\bar{f}_{t}(x^{K}_{t})-\bar{f}_{t}(x_{t})\right),$ which can be rewritten as
$R(T,K)$ $=\frac{1}{T}\sum_{t=1}^{T}\left(\bar{f}_{t}(x^{K}_{t})-\bar{f}_{t-1}(x_{t-1})\right)+\frac{1}{T}\sum_{t=1}^{T}\left(\bar{f}_{t-1}(x_{t})-\bar{f}_{t}(x_{t})\right).$
Note that the second summation telescopes to $(1/T)\left(\bar{f}_{0}(x_{0})-\bar{f}_{T}(x_{T})\right).$ We then begin by bounding the terms in the first summation. We observe that
$\bar{f}_{t}(x^{K}_{t})\leq\bar{f}^{K}_{t+1}(x^{K}_{t})+\min\left\{\Phi^{K}_{t},\max_{j\leq J}M_{j}(2\varepsilon_{t}+d^{K}_{t,1})\right\}=\bar{f}^{K}_{t+1}(x^{K}_{t})+\underline{\psi}_{t+1}^{K},$
using [56, Theorem 5] and Kantorovich-Rubinstein duality. Using the same theorems, we observe $\bar{f}_{t-1}(x_{t-1})\geq\bar{f}_{t}^{K}(x_{t-1})-\bar{\psi}_{t}^{K},$ and by the optimality of $x^{K}_{t}$ for the particular problem, $\bar{f}_{t}^{K}(x_{t-1})\geq\bar{f}_{t}^{K}(x^{K}_{t})$. Furthermore, by Kantorovich-Rubinstein duality
$\bar{f}^{K}_{t+1}(x^{K}_{t})-\bar{f}_{t}^{K}(x^{K}_{t})\leq\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+W_{1}(\hat{{\mathbf{P}}}^{K}_{t},\hat{{\mathbf{P}}}^{K}_{t-1})).$
Combining these relations, we obtain
$\frac{1}{T}$ $\sum_{t=1}^{T}(\bar{f}_{t}(x^{K}_{t})-\bar{f}_{t-1}(x_{t-1}))\leq\frac{1}{T}\sum_{t=1}^{T}(\underline{\psi}_{t+1}^{K}+\bar{\psi}_{t}^{K}+\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+W_{1}(\hat{{\mathbf{P}}}^{K}_{t},\hat{{\mathbf{P}}}^{K}_{t-1})).$ (18)
For the final term, we show again by Kantorovich-Rubinstein duality
$\bar{f}_{0}(x_{0})-\bar{f}_{T}(x_{T})\leq\bar{f}_{0}(x_{T})-\bar{f}_{T}(x_{T})\leq\max_{j\leq J}M_{j}(\varepsilon_{0}+\varepsilon_{T}+W_{1}(\hat{{\mathbf{P}}}^{0},\hat{{\mathbf{P}}}_{t})).$
We use this to prove Theorem 4.1 and its corollaries.
Proof of Theorem 4.1 and Corollary 4.1.1.
We examine the terms in $R(T,K)$ individually, for a finite value $T\gg\tau$. As we assume the Lipschitz condition, we use the corresponding terms in $\underline{\psi}^{K}_{t-1}$ and $\bar{\psi}^{K}_{t}$ as the upper bounds. By assumption, $\varepsilon_{t}\sim(\log(\beta_{t}^{-1})/t)^{1/d}$, so their averages are $O((\log(\beta_{T}^{-1})/T)^{1/d}).$ Note that by the summability of $\beta_{t}$, the convergence rate of $\beta_{t}$ is at least sublinear. Next, recall by the triangle inequality,
$d^{K}_{t,1}\leq W_{1}(\hat{{\mathbf{P}}}_{t},{\mathbf{P}})+W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})+W_{1}({\mathbf{P}}^{K}_{\star},{\mathbf{P}}^{K}_{t}).$
By Theorem 3.1, we can construct a summable sequence $\beta_{t}$ such that ${\mathbf{P}}(W_{1}(\hat{{\mathbf{P}}}_{t},{\mathbf{P}})\leq\varepsilon_{t})\geq 1-\beta_{t}$, where $\varepsilon_{t}$ is the same as above. By Theorem 3.4, a similar result holds for $W_{1}({\mathbf{P}}^{K}_{\star},\hat{{\mathbf{P}}}^{K}_{t})$, with a sequence of radii $\tilde{\varepsilon}_{t}\leq O(\varepsilon_{t})$ for $t\geq\tau$. Since $T\gg\tau$, the finite values $\tilde{\varepsilon}_{t}/T$ for $t\leq\tau$ are dominated by $O(\varepsilon_{T})$. Therefore, the terms involving $d^{K}_{t,1}$ and $d^{K}_{t-1,1}$ are together $O(W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})+\left(\log(\beta_{T}^{-1})/T\right)^{1/d})$, with probability at least $\Pi_{t=0}^{T-1}(1-\beta_{t})^{2}$. This probability can be constructed to converge to a nonzero value due to the summability of $\beta_{t}$. Next, we note again by the triangle inequality,
$W_{1}(\hat{{\mathbf{P}}}^{0},\hat{{\mathbf{P}}}_{t})\leq W_{1}(\hat{{\mathbf{P}}}^{0},{\mathbf{P}})+W_{1}({\mathbf{P}},\hat{{\mathbf{P}}}_{t}),$
where the first term is constant in $T$ and the second term, by measure concentration results, is upper bounded by $\varepsilon_{T}$ with some probability $1-\beta_{T}$. Therefore, with probability $1-\beta_{T}$, $W_{1}(\hat{{\mathbf{P}}}^{0},\hat{{\mathbf{P}}}_{t})/T$ is $O(T^{-1})$. Lastly, by the triangle inequality,
$W_{1}(\hat{{\mathbf{P}}}^{K}_{t},\hat{{\mathbf{P}}}^{K}_{t-1})\leq W_{1}(\hat{{\mathbf{P}}}^{K}_{t},{\mathbf{P}}^{K}_{\star})+W_{1}({\mathbf{P}}^{K}_{\star},\hat{{\mathbf{P}}}^{K}_{t-1})\leq 2\tilde{\varepsilon}_{t-1},$
with probability $(1-\beta_{t-1})^{2}$. The average over $T$ is again $O(\varepsilon_{T})$, with probability $\Pi_{t=0}^{T-1}(1-\beta_{t})^{2}$. Combining all terms, we have $R(T,K)\leq O(\left(\log(\beta_{T}^{-1})/T\right)^{1/d})+O(W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star}))$, with a consolidated probability $1-\zeta\geq\Pi_{t=0}^{T}(1-\beta_{t})^{4}$. Again, we remark that the probability $1-\zeta$ can be constructed to be high, with well-chosen sequences $\beta_{t}$ and $\varepsilon_{t}$. Note that their convergence rates are inversely related; the higher the desired probability $1-\zeta$, the slower the convergence with respect to $T$. Using the Bonferroni Inequality [12], the probability can be estimated as $1-\zeta\geq 1-\sum_{t=0}^{T}4\beta_{t}$.
Corollary 4.1.1 follows by setting $\beta_{t}=O(\exp(-\sqrt{n_{t}})$ and simplifying. ∎
Proof of Corollary 4.1.2.
The asymptotic result for $T\rightarrow\infty$ follows from Lemma 4.1 and the results from Section 3. As $t\to\infty$ and $T\to\infty$, by the almost sure convergence of $\underline{\psi}_{t}^{K}+\bar{\psi}_{t}^{K}$ to $\underline{\psi}_{\star}^{K}+\bar{\psi}_{\star}^{K}$, $W_{1}(\hat{{\mathbf{P}}}^{0},\hat{{\mathbf{P}}}_{t})/T$ to $W_{1}(\hat{{\mathbf{P}}}^{0},{{\mathbf{P}}})/T$ to 00, and $W_{1}(\hat{{\mathbf{P}}}^{K}_{t},\hat{{\mathbf{P}}}^{K}_{t-1})$ to 00, by the Cesàro Mean Theorem [15], $R(T,K)$ converges to some value that is $O(\underline{\psi}_{\star}^{K}+\bar{\psi}_{\star}^{K})$ with probability 1. The result for $K\rightarrow\infty$ follows from Theorem 3.2. ∎
A.3 Proofs of Section 5 results
Proof of Lemma 5.1.
Let the center of the $k$-th cluster be $a^{k}$. By definition of the Wasserstein metric,
$d_{W_{p}}(K,t)^{p}=W_{p}(\hat{{\mathbf{P}}}_{t},\hat{\mathbf{P}}^{K}_{t})^{p}$ $=\inf_{\pi\in\Pi}\left\{\int_{{S}^{2}}\|u-u^{\prime}\|^{p}\pi(du,du^{\prime})\right\}$
$\leq\sum_{k=1}^{K}\frac{n^{t}_{k}}{n_{t}}\int_{S^{2}}\|u-\bar{u}^{k}_{t}\|^{p}\hat{{\mathbf{P}}}_{t}(u|u^{\prime}=\bar{u}^{k}_{t})(du)$
$\leq\sum_{k=1}^{K}\frac{n_{k}^{t}}{n_{t}}\frac{1}{n_{k}^{t}}\sum_{\hat{u}\in C^{k}_{t}}\|\hat{u}-\bar{u}^{k}_{t}\|^{p}$
$=\frac{1}{n_{t}}\sum_{k=1}^{K}\sum_{\hat{u}\in C^{k}_{t}}\|\hat{u}-\bar{u}^{k}_{t}\|^{p}$
$\leq\frac{1}{n_{t}}\sum_{k=1}^{K}\sum_{\hat{u}\in C^{k}_{t}}(\|\hat{u}-a^{k}\|+\|a^{k}-\bar{u}^{k}_{t}\|)^{p}$
$\leq(2\eta_{K})^{p},$
where the final inequality follows from Assumption 6.1. Next, by exploiting the triangular inequality, with probability $1-\beta$, we get
$W_{p}(\mathbf{P},\hat{\mathbf{P}}^{K}_{t})\leq W_{p}(\mathbf{P},\hat{{\mathbf{P}}}_{t})+W_{p}(\hat{\mathbf{P}}_{t},\hat{\mathbf{P}}^{K}_{t})\leq\varepsilon_{t}+2\eta_{K},$ (19)
where $\varepsilon_{t}$ is computed as in Theorem 5.1. ∎
Proof of Lemma 5.2.
As $\mathbf{Q}_{t}\in\mathbf{B}_{\varepsilon^{K}_{t}}(\hat{\mathbf{P}}_{t}^{K})$, from the triangular inequality,
$W_{p}(\mathbf{P},\mathbf{Q}_{t})\leq W_{p}(\mathbf{P},\hat{\mathbf{P}}^{K}_{t})+W_{p}(\hat{\mathbf{P}}^{K}_{t},\mathbf{Q}_{t})\leq W_{p}(\mathbf{P},\hat{\mathbf{P}}^{K}_{t})+\varepsilon^{K}_{t}$
From (19) we have $W_{p}(\mathbf{P},\hat{\mathbf{P}}^{K}_{t})\leq\varepsilon_{t}+2\eta_{K}=\varepsilon^{K}_{t}$. Thus, we have $\mathbf{P}(W_{p}(\mathbf{P},\mathbf{Q}_{t})\leq 2\varepsilon^{K}_{t})\geq 1-\beta_{t}$. Since by definition $\sum_{t=0}^{\infty}\beta_{t}<\infty$, then by the Borel-Cantelli Lemma we get
$\mathbf{P}^{\infty}\{W_{p}(\mathbf{P},\mathbf{Q}_{t})\leq 2\varepsilon^{K}_{t}\text{~{}~{}for all large enough $t$}\}=1.$
Since by definition $\lim_{t\rightarrow\infty}2\varepsilon^{K}_{t}=4\eta_{K}$, we conclude $\lim_{t\rightarrow\infty}W_{p}(\mathbf{P},\mathbf{Q}_{t})\leq 4\eta_{K}$ almost surely. ∎
Proof of Theorem 5.3.
By the finite sample performance guarantee in Theorem 5.2 and the summability of $\beta_{t}$, the Borel–Cantelli Lemma implies that
$\mathbf{P}^{\infty}\left\{H_{\star}\leq\limsup_{t\rightarrow\infty}{\mbox{\bf E}}_{\mathbf{P}}[f(u,x^{K}_{t})]\leq\limsup_{t\rightarrow\infty}H^{K}_{t}\right\}=1.$
Now, choose any $\gamma>0$, and fix a $\gamma$-optimal solution $x_{\gamma}\in\mathcal{X}$ of (1) such that ${\mbox{\bf E}}_{{\mathbf{P}}}[f(x_{\gamma},u)]\leq H_{\star}+\gamma$. For each $t$, we choose a $\gamma$-optimal distribution $\mathbf{Q}_{t}\in\mathcal{P}^{t-1}$ such that $\sup_{\mathbf{Q}\in\mathcal{P}^{t-1}}{\mbox{\bf E}}_{\mathbf{Q}}[f(x_{\gamma},u)]\leq{\mbox{\bf E}}_{\mathbf{Q}_{t}}[f(x_{\gamma},u)]+\gamma.$ Then, we observe that
$\limsup_{t\rightarrow\infty}H^{K}_{t}$ $\leq\limsup_{t\rightarrow\infty}\sup_{\mathbf{Q}\in\mathcal{P}^{t-1}}{\mbox{\bf E}}_{\mathbf{Q}}[f(x_{\gamma},u)]$
$\leq\limsup_{t\rightarrow\infty}{\mbox{\bf E}}_{\mathbf{Q}_{t}}[f(x_{\gamma},u)]+\gamma$
$\leq\limsup_{t\rightarrow\infty}{\mbox{\bf E}}_{\mathbf{{\mathbf{P}}}}[f(x_{\gamma},u)]+\max_{j\leq J}M_{j}W_{1}(\mathbf{Q}_{t},{\mathbf{P}})+\gamma$
$\leq H_{\star}+4\max_{j\leq J}M_{j}\eta_{K}+2\gamma,$
where the first inequality follows from the optimality of $x^{K}_{t}$ and $H^{K}_{t}$, and the third inequality follows from Kantorovich-Rubinstein duality [33] and the Lipschitz condition of $f$. The final inequality follows from Lemma 5.2. Since we chose $\gamma>0$ arbitrarily, we can conclude that $\limsup_{t\rightarrow\infty}H^{K}_{t}\leq H_{\star}+4\max_{j\leq J}M_{j}\eta_{K}$. ∎
A.4 Online clustering algorithm
We give details for the online clustering algorithm, introduced in Section 6. We keep track of two sets of clusters: a set of $Q\geq K$ microclusters, and a set of $K$ macroclusters. For simplicity, below we assume the number of microclusters to be $Q$ and the number of macroclusters to be $K$; if the number of data-points is smaller, i.e., $Q_{t}\leq Q$ and $K_{t}\leq K$, the arrival of new data-points result in new clusters.
Initialization
We initialize the problem by solving approximately the $k$-centers problem (10) with respect to $\hat{{\mathbf{P}}}^{0}$, allowing up to $Q$ clusters. This set of centers is denoted $A^{Q}_{0}$. We assign all points to the closest center, and define the support of the $q$-th cluster $S^{q}_{0}$ as the Voronoi region around its center $a_{0}^{q}\in A^{Q}_{0}$. Note that the Voronoi regions need not be explicitly stored; they are implicitly defined by their centers. For each cluster, we note its center, centroid, root-mean-squared-error (RMSE), and weight. These are denoted the microclusters. We refer to the RMSE of each cluster as $\eta^{q}_{t}$, and for clusters with only a single data-point, heuristically initialize it as twice the minimum RMSE of all clusters.
To create $K$ macroclusters, we solve approximately the $k$-centers problem with respect to $A^{Q}_{0}$, and obtain a set of centers $A^{K}_{0}$. Each microcluster is assigned to the macrocluster with the closest center; each macrocluster then contains all the points of its constituent microclusters, and has their combined weight. The centroid of the macrocluster is therefore the weighted average of the centroids of the constituent microclusters. The support of the $k$-th macrocluster is the Voronoi region around the $k$-th center $a_{0}^{k}\in A^{K}_{0}$, plus the data-points assigned to it, and minus the data-points assigned to other clusters, i.e.,
$S^{k}_{t}=(V(a_{t}^{k})\cup\{C^{k}_{t}\})/\{C^{k^{\prime}}_{t}\}_{k^{\prime}\neq k}.$ (20)
In this manner, the set of supports $\mathcal{S}_{t}$ satisfy Assumption 3.2. The necessity of this adjustment of the support follows from the two-layer nature of the clustering algorithm. It is possible for a data-point to be clustered into the $q$-th microcluster, which is assigned to the $k$-th macrocluster, but distance-wise is actually closer to the center of $k^{\prime}$-th macrocluster for some $k^{\prime}\neq k$.
Online updating procedure
For $t\geq 1$, when we observe a new data-point $\hat{u}$, we calculate its distance to the microcluster whose support it falls on, i.e., if it falls on support $S^{q}_{t-1},$
$d_{t}=\|\hat{u}-a^{q}_{t-1}\|.$
If this value is below $2\eta^{q}_{t-1}$, where $\eta^{q}_{t-1}$ is the RMSE, we assign it to this cluster. The multiplier of 2 is selected according to CluStream sensitivity analysis [2, Section 6.4]. We then update all microclusters weights, as well as the centroid and RMSE of the assigned microcluster. The macroclusters are also adjusted accordingly, based on their constituent microclusters.
On the other hand, if $d_{t}>2\eta^{q}_{t-1}$, we create a new microcluster with this data-point as its center, and assign it the weight $1/n_{t}$. The weights of the other microclusters are decreased accordingly, following (17). The RMSE of this cluster is initialized heuristically to be twice the minimum RMSE of all other existing clusters. If the total number of microclusters exceeds $Q$ with this addition, we merge the two closest microclusters based on the distances between their centers. The center of the merged microcluster will be the weighted average of the centers of the two constituent microclusters, with the new centroid, RMSE, and weight calculated accordingly. The supports of the microclusters will be reassigned as in (20). Then, we again perform approximate $k$-centers on the $Q$ microclusters, warm-staring at the previous centers, and generate $K$ macroclusters. The parameters of the macroclusters are assigned the same manner as in initialization.
At some time step $\tau<\infty$, we terminate the support updating procedure, and only cluster points based on the latest support $\hat{S}^{K}_{\tau}$.
The online DRO algorithm with online clustering is summarized in Algorithm 3. If a parameter is not explicitly updated, we assume it inherits the value from the previous time step.
Remark
In the algorithm and description, the cluster assignments are assumed to use the cluster centers. The centroid can also be used, however, up to time $t\leq\tau$.
Overall, by keeping $Q\geq K$ microclusters, we allow for a finer clustering algorithm than keeping only $K$ clusters at all times. In this manner, the microclusters are allowed to switch macrocluster assignments when the $k$-centers update is performed, thereby minimizing the information loss induced by the $K$-cluster budget.
Algorithm 3 OnlineClustering
1:given $\mathcal{D}_{0},K,Q,S,\tau,\{\varepsilon^{K}_{t}\}_{t=0,1,\dots}$
2:$A^{Q}_{0}\leftarrow$ find $Q$ centers using approximate $k$-centers on $\mathcal{D}_{0}$
3:assign all data-points to microclusters $C^{q}_{0}$, and compute $n^{q}_{0}$, $\theta^{q}_{0}$, $\bar{u}_{0}^{q}$, and $\eta^{q}_{0}$
4:$\hat{S}^{Q}_{0}\leftarrow$ (20)
5:$A^{K}_{0}\leftarrow$ find $K$ centers using approximate $k$-centers on $A^{Q}_{0}$
6:assign corresponding microclusters to macroclusters $C^{k}_{0}$, and compute $n^{k}_{0}$, $\theta^{k}_{0}$, and $\bar{u}_{0}^{k}$
7:$\hat{S}^{K}_{0}\leftarrow$ (20)
8:for $t=1,2,\dots$ do
9:     $\hat{\mathbf{P}}^{K}_{t}\leftarrow\sum_{k=1}^{K}\theta_{t}^{k}\delta_{\bar{u}_{t}^{k}}$
10:     observe data-point $\hat{u}$, and set $d_{t}\leftarrow\|\hat{u}-a^{q^{\prime}}_{t-1}\|$, where $q^{\prime}$ is chosen such that $\hat{u}\in S^{q^{\prime}}_{t-1}$
11:     if $t\geq\tau$ then
12:         $C_{t}^{k^{\prime}}\leftarrow C_{t}^{k^{\prime}}\cup\{\hat{u}\}$triangleright assign $\hat{u}$ following  (16)
13:         $\theta^{k}_{t}\leftarrow$ (17) $\triangleright$ update weights
14:         $\bar{u}^{k^{\prime}}_{t}\leftarrow(n^{k^{\prime}}_{t-1}\bar{u}^{k^{\prime}}_{t-1}+\hat{u})/n^{k^{\prime}}_{t}$triangleright update centroid
15:     else if $d_{t}\leq 2\eta^{q^{\prime}}_{t}$ then
16:         $C_{t}^{q^{\prime}}\leftarrow C_{t}^{q^{\prime}}\cup\{\hat{u}\}$
17:         $\theta^{q}_{t}\leftarrow$ (17) $\triangleright$ update microcluster weights
18:         $\bar{u}^{q^{\prime}}_{t}\leftarrow(n^{q^{\prime}}_{t-1}\bar{u}^{q^{\prime}}_{t-1}+\hat{u})/n^{q^{\prime}}_{t}$$\triangleright$ update centroid
19:         $\eta^{q}_{t}\leftarrow(n^{q^{\prime}}_{t-1}\eta^{q^{\prime}2}_{t-1}+\|\hat{u}-\bar{u}^{q^{\prime}}_{t}\|_{2}^{2})/n^{q^{\prime}}_{t}$$\triangleright$ update RMSE
20:         $\theta^{k}_{t}\leftarrow$ (17) $\triangleright$ update macrocluster weights
21:         update $\bar{u}^{k^{\prime}}_{t}$ for microcluster $k^{\prime}$, where $C^{q^{\prime}}_{t}\subseteq C^{k^{\prime}}_{t}$
22:         $\hat{S}^{K}_{t}\leftarrow$ (20)
23:     else
24:         assign $\hat{u}$ to a new cluster $q^{\star}$, initialize $\eta^{q^{\star}}_{t}\leftarrow\min_{q\leq Q}2\eta^{q}_{t-1}$
25:         $\theta^{q}_{t}\leftarrow$ (17) $\triangleright$ update weights
26:         $A^{Q}_{t}\leftarrow$ merge the two microclusters with the closest centers
27:         $\hat{S}^{Q}_{t}\leftarrow$ (20)
28:         $A^{K}_{t}\leftarrow$ find $K$ centers using approximate $k$-centers on $A^{Q}_{t}$
29:         assign microclusters to macroclusters $C^{k}_{t}$, and compute $n^{k}_{t}$, $\theta^{k}_{t}$, and $\bar{u}_{t}^{k}$
30:         $\hat{S}^{K}_{t}\leftarrow$ (20)