### v2

# Fast Online Distributionally Robust Optimization via Data Compression

Abstract
We propose an online data compression approach for efficiently solving distributionally robust optimization (DRO) problems with streaming data while maintaining out-of-sample performance guarantees. Our method dynamically constructs ambiguity sets using online clustering, allowing the clustered configuration to evolve over time for an accurate representation of the underlying distribution. We establish theoretical conditions for clustering algorithms to ensure robustness, and show that the performance gap between our online solution and the nominal DRO solution can be written in terms of the distance between the true and compressed distributions. Therefore, by varying the number of clusters, our method effectively balances robustness and online computational efficiency. We show that our analysis is compatible with well-established finite-sample and asymptotic guarantees for Wasserstein DRO. Numerical experiments in mixed-integer portfolio optimization demonstrate significant computational savings, with minimal loss in solution quality.
1 Introduction
Many decision-making problems in engineering, operations research, and computer science involve solving optimization problems affected by uncertain parameters. Distributionally robust optimization (DRO) has emerged as a disciplined framework to robustify the decision-making process against distributional misspecification [43, 34]. Instead of assuming a known distribution, DRO considers an ambiguity set of possible distributions and optimizes for the worst-case expected cost under all distributions within this set. This approach is particularly appealing as it offers out-of-sample performance guarantees with a finite number of data samples. By accounting for ambiguity, DRO often outperforms non-robust methods such as sample average approximation (SAA), especially with respect to worst-case performance [18].
Since its appearance, significant effort has been devoted to constructing ambiguity sets primarily in static settings, using either structural assumptions on the distribution (e.g., moments [44, 21]) or training data directly [33, 18]. However, many real-time settings, such as healthcare resource allocation, air traffic control, and energy trading, are dynamic and new data becomes available sequentially over time. Despite its potential, DRO has been largely underdeveloped in such scenarios. Most DRO approaches construct a static ambiguity set, and model time progression through adjustable wait-and-see variables [15, 8, 6]. There are limited works focusing on how to adapt the ambiguity sets dynamically over time, where they pose strict assumptions on the problem: either by restricting the distribution to a finite set of scenarios [3], or by requiring continuous variables and objective functions with local strong concavity [35]. In particular, streaming data settings can be challenging for the popular Wasserstein-based DRO [38, 33, 18], where the ambiguity set is defined as a Wasserstein ball around the empirical distribution of a dataset. As new datapoints arrive, the problem dimension grows, increasing computational complexity [33, 54]. While a larger dataset improves confidence in the empirical measure as an approximation of the true distribution, the reduction in the ambiguity set’s radius is not enough to offset this complexity; for mixed-integer problems, even SAA is difficult to solve with large amounts of data, without any added complexity from ambiguity. This leads to computational bottlenecks that are especially limiting for real-time applications.
1.1 Our contributions
We propose an online data compression approach for efficiently solving Wasserstein DRO problems with streaming data, while also maintaining out-of-sample performance guarantees. We allow for mixed-integer formulations, and a broad class of underlying distributions. Our key contributions are:
  • Adaptive ambiguity sets via online clustering: We apply online clustering to construct adaptive ambiguity sets, formulated as Wasserstein balls of any order around a clustered empirical distribution. We establish theoretical conditions on clustering algorithms to ensure the robustness of our framework, which is compatible with any online clustering method that meets these conditions. We formalize the concept of optimal clustering, and provide three fast and memory-efficient approximation algorithms suitable for our purposes.
  • Clustering effect analysis: We prove that the performance gap between our compressed solution and the non-compressed (nominal) DRO solution can be written in terms of the distance between the true distribution and its clustered approximation, and the ambiguity set radius. Under mild assumptions, this radius can decay at a rate independent of the uncertainty dimension. We also show that our objective converges to a value dependent on the number of clusters and the quality of the clustering algorithm. This quantifies the impact of data compression, and highlights the trade-off between computational effort and optimality. We provide both finite-sample and asymptotic performance guarantees for our solution.
  • Computational gains in numerical experiments: We demonstrate the efficiency of our approach in sparse portfolio optimization. Our results show significant computational savings, even compared to SAA, with minimal loss in solution quality. We also demonstrate the possibility for the online framework to be memory-efficient, allowing the optimizer to discard datapoints once seen, with minimal impact on solution times and quality.
1.2 Related work
Distributionally robust optimization
DRO has been extensively explored in recent years, with successful applications (among others) in machine learning [45], finance [9], and medicine [49]. Typical ambiguity sets that appeared in the literature include support, moment, or distance-based sets of distributions or mixtures thereof. We focus our attention on discrepancy-based ambiguity sets, defined as a ball in the space of probability distributions around a nominal or most-likely distribution, which is constructed from data. In this setting, the distance, commonly expressed in terms of, e.g., the $\phi$-divergence [5], the total variation norm [50], the kernel mean embedding [22], contamination techniques [32], or optimal transport based-distances including the celebrated Wasserstein distance [38, 29], signifies the “trust” in the data at hand. More recently, extensions, such as the trade-off ambiguity set [48] and the globalized ambiguity set [36], have been introduced to mitigate some of the conservatism of the classical DRO models. Due to the favorable properties of Wasserstein distance in terms of expressivity and statistical properties, we will focus on Wasserstein DRO.
Online learning and DRO
Online learning is a well-established framework providing algorithms for solving repetitive problems over time. Recently, it has been applied to robust optimization problems [25, 14], as well as to DRO formulations. For example, [41] considers an online DRO model with $\phi$-divergence ambiguity sets and proposes an alternating mirror descent algorithm to solve it, while [42] proposes a duality-free online stochastic method for regularized DRO problem with KL-divergence regularization. We focus here on a different problem, where the same DRO problem needs to be solved repeatedly over time with a growing knowledge of the uncertainty distribution. In this sense, our work is more closely related to [3], where a projected gradient descent method is used to adapt the ambiguity set to the samples collected progressively over time. However, [3] only considers discrete distributions. As a result, this work naturally disregards the fundamental computational challenge of incorporating streaming data in continuous settings, which is instead our main focus. Recently, [35] also considers the online DRO problem with data assimilation, but restricts their framework to continuous variables and strongly concave functions with accessible gradients. Our approach is therefore more general, and importantly, allows for mixed-integer formulations.
Scenario reduction and DRO
It is well-known that the size and computational complexity of data-driven optimization problems generally increase with the number of samples, resulting in a fundamental trade-off between statistics and computation. To overcome the computational bottleneck, scenario reduction techniques are often applied. They aim to reduce the number of scenarios while retaining a good enough representation of the underlying uncertainty and thus an accurate solution to the data-driven optimization problem. A popular approach consists of generating new scenarios and assigning probabilities to minimize the distance to the original distribution [16]. More recently, decision-focused scenario reduction techniques have been proposed, where the loss function itself is used in the construction of metrics to aggregate scenarios [7, 56]. A variety of techniques to aggregate scenarios have been suggested in the literature, based on clustering [24], moment matching [37], objective approximation [55], and nested distances [27]. We consider scenario reduction for DRO problems closely related to recent developments in two-stage robust optimization, such as [7, 56]. Our previous work [54] studies the impact of clustering in static and finite-sample DRO problems, using it as a tool to bridge robust and distributionally robust optimization. Closely related is also [4], which embeds scenario reduction into a DRO framework and provides suboptimality bounds—albeit limited to monotonically homogeneous uncertain objectives with strictly positive uncertainty. However, both [54] and [4] assume that data are available a priori, while our focus here is on online scenario reduction for DRO problems with streaming data.
Clustering with streaming data
Due to the ongoing data revolution, data stream clustering has recently attracted attention for emerging applications that involve large amounts of streaming data. They can be broadly classified into partition-based algorithms that partition data into clusters using distance-based similarity metrics, density-based algorithms that define clusters as dense partitions separated by sparse areas that dynamically change over time to adapt to data evolving distributions (such as DenStream [12]), hierarchical-based algorithms that maintain a tree-like structure by grouping similar clusters at different levels [40]. Among the first class, the most popular ones are undoubtedly incremental $k$-mean [1] and CluStream [2], which we adapt to fit our framework. We point to the survey [13] for further details. We highlight that a major strength of our approach is its ability to incorporate state-of-the-art online clustering algorithms, with various memory and run-time complexities, with the choice left to the user’s discretion.
1.3 Layout of the paper
In Section 2, we state the problem and introduce our online algorithm, and in Section 3, analyze the clustering effect by providing finite sample and asymptotic performance guarantees. In Section 4, we give theoretical results for the specific case of bounded support. In Section 5, we describe various online clustering algorithms, and in Section 6, give guidelines for choosing the number of clusters and the ambiguity set radii experimentally. In Section 7 we demonstrate our results on a portfolio optimization example.
2 The online stochastic problem
We consider a stochastic optimization (SO) problem of the form
$H_{\star}=\min_{x\in\mathcal{X}}\mathbf{E}_{u\sim\mathbf{P}}[f(u,x)],$ (1)
where $x\in\mathcal{X}\subseteq\mathbf{R}^{n}$ is the decision variable, $\mathcal{X}$ a compact set, $u\in S\subseteq\mathbf{R}^{d}$ the vector of uncertain parameters that is governed by some probability distribution $\mathbf{P}$, and $H_{\star}$ the optimal objective value. We assume that the support $S$ of $\mathbf{P}$} lives within the domain of $f$ for the variable $u$, which we will refer to as $\mathop{\bf dom}_{u}{f}$, i.e., $S\subseteq\mathop{\bf dom}_{u}{f}$. The function $f:\mathcal{X}\times S\rightarrow(-\infty,\infty]$ is assumed to be of the form
$f(u,x)=\max_{j\leq J}f_{j}(u,x),$
with each $f_{j}$ being proper, concave, and upper-semicontinuous in $u$ for all $x$. Additionally, we require the functions $f_{j}$ to be either Lipschitz or smooth for all $x\in\mathcal{X}$, according to the following definitions. We assume the existence of the global constants independent of $x$ by the compactness of $\mathcal{X}$.
Assumption 2.1 (Lipschitzness and smoothness).
Let at least one hold.
  1. 1.
    For all $x\in\mathcal{X}$, the constituent functions $f_{j}$ of $f$ are Lipschitz continuous with constants $M_{j}$, i.e., $\left|f_{j}(v,x)-f_{j}(u,x)\right|\leq M_{j}\|u-v\|_{2},\;\forall u,v\in\mathop{\bf dom}_{u}f$.
  2. 2.
    For all $x\in\mathcal{X}$, the constituent functions $f_{j}$ of $f$ are smooth with constants $L_{j}$, i.e., $\left\|\nabla f_{j}(v,x)-\nabla f_{j}(u,x)\right\|_{2}\leq L_{j}\|u-v\|_{2},\;\forall u,v\in\mathop{\bf dom}_{u}f$.
  3. 3.
    For all $x\in\mathcal{X}$, $f$ is smooth with constant  $L$, i.e., $\left\|\nabla f(v,x)-\nabla f(u,x)\right\|_{2}\leq L\|u-v\|_{2},\;\forall u,v\in\mathop{\bf dom}_{u}f$.
Analogously to [54], we assume the following on the domain of $f$.
Assumption 2.2.
The domain $\mathop{\bf dom}_{u}{f}$ is ${\bf R}^{d}$. Otherwise, $f$ is either element-wise monotonically increasing in $u$ and only has a (potentially) lower-bounded domain, or element-wise monotonically decreasing in $u$ and only has a (potentially) upper-bounded domain. In all cases, the domain $\mathop{\bf dom}_{u}{f}$ is independent of $x$.
Streaming data
We assume the distribution $\mathbf{P}$} to be unknown, and that the value $H_{\star}$ cannot be calculated as is. However, we have access to a streaming dataset of i.i.d. realization of $u$, which we use to construct an uncertainty framework. We begin with an initial dataset with $n_{0}\geq 1$ datapoints, i.e., $\mathcal{D}_{0}=\{\hat{u}^{i}\}_{i=1}^{n_{0}}$. Then, at the end of each time step $t\geq 1$, the dataset is updated to be $\mathcal{D}_{t}=\{\hat{u}^{i}\}_{i=1}^{n_{0}+t}$, with $|\mathcal{D}_{t}|=n_{t}=n_{0}+t$. That is, we assume w.l.o.g. that one realization gets disclosed at the end of each time step. Furthermore, we define the empirical distribution $\hat{{\mathbf{P}}}_{t}=(1/{n_{t}})\sum_{i=1}^{n_{t}}\delta_{\hat{u}^{i}},$ where $\delta_{u}$ is the Dirac distribution concentrating unit mass on $u$.
2.1 DRO over time
Let $\Xi_{p}(S)$ be the space of all probability distributions $\mathbf{Q}$ defined on the support $S$, with bounded $p$-th moments, i.e., ${\mbox{\bf E}}_{\mathbf{Q}}[\|u\|^{p}]=\int_{S}\|u\|^{p}\mathbf{Q}(du)<\infty$, where $\|\cdot\|$ is an arbitrary norm on ${\bf R}^{d}$, and either $p\geq 1$ or $p=\infty$.
The DRO approach under the Wasserstein metric constructs an ambiguity set of order $r$ around $\hat{{\mathbf{P}}}_{t}$ of radius $\varepsilon_{t}\geq 0$, i.e.,
$\mathcal{P}_{t}=\mathbf{B}_{\varepsilon_{t}}^{r}(\hat{{\mathbf{P}}}_{t})=\left\{\mathbf{Q}\in\Xi_{r}({S})\mid W_{r}(\hat{{\mathbf{P}}}_{t},\mathbf{Q})\leq\varepsilon_{t}\right\},$ (2)
where the Wasserstein type-$p$ distance between two probability distributions $\mathbf{Q}_{1},\mathbf{Q}_{2}\in\Xi_{p}(S)$ is defined by
$W_{p}(\mathbf{Q}_{1},\mathbf{Q}_{2})=\inf_{\pi\in\Pi(\mathbf{Q}_{1},\mathbf{Q}_{2})}\left(\int_{S\times S}\left\|u_{1}-u_{2}\right\|^{p}\mathrm{d}\pi\left(u_{1},u_{2}\right)\right)^{1/{p}},$ (3)
and $\Pi(\mathbf{Q}_{1},\mathbf{Q}_{2})$ is the set of all probability distributions over $S\times S$ with marginals $\mathbf{Q}_{1}$ and $\mathbf{Q}_{2}$. In this work, we always refer to $r$ as the order of the Wasserstein distance used within an ambiguity set. For other Wasserstein distances used in the analysis, we refer to order $p$. In addition, whenever we use a distance metric of order $p$ on a distribution $\mathbf{Q}$, we assume that the distribution itself has finite $p$-th moments. We suppress the dependency of $\mathcal{P}_{t}$ on $r$ for conciseness, and note that unless otherwise specified, $r$ is of an arbitrary order and $\|\cdot\|$ is an arbitrary norm. Intuitively, the ambiguity set $\mathcal{P}_{t}$ contains all probability distribution in $\Xi(S)$ that can obtained by transporting probability mass from $\hat{{\mathbf{P}}}_{t}$ when the transportation budget is at most $\varepsilon_{t}$.
With this uncertainty modeling framework in place, we compute the DRO solution $x_{t}\in\mathcal{X}$ at time step $t$ by solving the optimization problem
$H_{t}=\min_{x\in\mathcal{X}}\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{u\sim\mathbf{Q}}[f(u,x)],$ (4)
where we seek to minimize the worst-case (w.r.t. any distribution in the ambiguity set) expected value of $f$.
2.2 Performance guarantees
As $t$ increases, our confidence in $\hat{{\mathbf{P}}}_{t}$ increases, allowing us to safely reduce the radius $\varepsilon_{t}$, which can be interpreted as a measure of trust in the data. With suitable choices of $\varepsilon_{t}$ [17, 38, 18], which we will refer to as the nominal radius, we obtain finite sample performance guarantees of the form
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathbf{P}}^{t}\left(H_{\star}\leq{\mbox{\bf E}}_{{\mathbf{P}}}[f(u,x_{t})]\leq H_{t}+\rho_{t}\right)\geq 1-\beta_{t},$} (5)
where for all $t\geq 1$, $\beta_{t}>0$ is a specified time-varying probability of constraint violation, $\rho_{t}$ is a residual term, and ${\mathbf{P}}^{t}$ is the product distribution of the dataset $\mathcal{D}_{t}$. In this work, we make use of two such choices of $\varepsilon_{t}$, which require the distribution $\mathbf{P}$} to satisfy distinct assumptions. We summarize them here.
Dimension-dependent
One choice is to impose a light-tailed assumption, which is satisfied by sub-Gaussian distributions and by bounded distributions.
Assumption 2.3 (Light-tailed distribution).
There exists an exponent $\alpha>1$, such that $R={\mbox{\bf E}}_{\mathbf{P}}[\exp(\|u\|^{\alpha})]=\int_{S}\exp(\|u\|^{\alpha}){\mathbf{P}}(du)<\infty.$
Then, measure concentration results for Wasserstein-$p$ distances imply the following.
Theorem 2.1 (Guarantees using Wasserstein measure concentration [17, 33]).
Let $p$ be an arbitrary order. Under Assumption 2.3 for the case $p\geq 1$, and under analogous assumptions for $p=\infty$ [47], for $n_{t}\rightarrow\infty$ we can choose confidence levels $\beta_{t}\in(0,1)$ and corresponding radii $\varepsilon_{t}=O((\log(\beta^{-1}_{t})/{n_{t}})^{\min(p/d,1/2)})$ such that $\sum_{t=1}^{\infty}\beta_{t}<\infty$, ${\mathbf{P}}^{t}(W_{p}({\mathbf{P}},\hat{{\mathbf{P}}}_{t})\geq\varepsilon_{t})\leq\beta_{t},$ and $\lim_{t\rightarrow\infty}\varepsilon_{t}=0.$ The DRO problem (4) solved with $r=p$ and these radii gives the finite-sample guarantee (5) with $\rho_{t}=0$.
We point to the cited works [17, 47, 33] for the constants involved. In Section 4, we further calculate explicit radii for the specific case of bounded support $S$. From the above characterization, it is evident that larger datasets are beneficial for reducing the distributional ambiguity around the nominal distribution. Indeed, we can use Theorem 2.1 to conclude that $\mathbf{P}^{\infty}\{\lim_{t\rightarrow\infty}W_{r}(\mathbf{P},\hat{\mathbf{P}}_{t}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=0}\}=1$, which implies $\mathbf{P}^{\infty}\{\lim_{t\rightarrow\infty}H_{t}\downarrow H_{\star}\}=1$ under the Lipschitz condition of Assumption 2.1 [38, Theorem 3.7]. However, as the radius scales with $O(n_{t}^{-r/d})$, even for moderately sized $d$ this bound suffers from the curse of dimensionality. The slow decrease in the required radius while the dataset grows makes the resolution of the DRO problem in (4) computationally challenging, as a constraint needs to be added for each datapoint constituting the empirical distribution $\hat{\mathbf{P}}_{t}$ (see formulation (9)). Clearly, for $t\rightarrow\infty$, we would end up with an infinite-dimensional problem that cannot be solved.
Dimension-free
Recent work [18] decouples the radii $\varepsilon_{t}$ from the dimension $d$ by providing finite sample guarantees for type-$r$ Wasserstein DRO, $r\in[1,2]$, with radii scaling as $\tilde{O}(n_{t}^{-1/2})$ ($\tilde{O}$ suppresses logarithmic terms). These result require $\mathbf{P}$} to satisfy a transportation-information inequality, and $f$ to satisfy certain curvature conditions, stated below for completeness.
Assumption 2.4.
If we consider $r=1$ for (2), then $\mathbf{P}$} satisfies the transportation-information inequality $T_{1}(\alpha)$ for some positive constant $\alpha$, i.e., for all $\mathbf{Q}\in\Xi_{1}(S)$, $W_{1}(\mathbf{Q},{\mathbf{P}})\leq\sqrt{\alpha H(\mathbf{Q}\|{\mathbf{P}})}$, where $H(\mathbf{Q}\|{\mathbf{P}})$ denotes the relative entropy. Furthermore, $f$ satisfies clause 1 of Assumption 2.1. If $r=2$, then $\mathbf{P}$} satisfies $T_{2}(\alpha)$ for some positive constant $\alpha$, and $f$ satisfies clause 3 of Assumption 2.1.
If the objective function class $\mathcal{F}=\{f(\cdot,x)\mid x\in\mathcal{X}\}$ additionally satisfies certain assumptions on its Rademacher complexity [18, Assumption 5], the following holds.
Theorem 2.2 (Dimension-free performance guarantees [18]).
For orders $r=1$ and $r=2$, under Assumption 2.4 and a Rademacher complexity condition [18, Assumption 5], we can choose confidence levels $\beta_{t}=O(\log(n_{t}))\in(0,1)$, radii $\varepsilon_{t}=\tilde{O}(n_{t}^{-1/2})$, and residual terms $\rho_{t}=\tilde{O}(1/n_{t})$ such that for all $x\in\mathcal{X}$,
${\mathbf{P}}^{t}\left({\mbox{\bf E}}_{{\mathbf{P}}}[f(u,x)]\leq\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x)]+\rho_{t}\right)\geq 1-\beta_{t}.$ (6)
This implies the finite-sample guarantee (5) with the above parameter values.
In exchange for dimension-free radii $\varepsilon_{t}$, this result introduces a nonzero residual $\rho_{t}$. However, even with smaller radii, the difficulty of solving the DRO problem with a large number of datapoints can be high. We are thus motivated to introduce an additional tradeoff, where a separate residual term is established in exchange for a more compact DRO formulation. This leads to our framework below.
2.3 Our online algorithm
We propose a data-compression framework for fast online DRO. Our goal is to solve an adjusted DRO problem over time, where $K$ is the chosen number of data clusters, and
$H^{K}_{t}=\min_{x\in\mathcal{X}}\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t}}\mathbf{E}_{u\sim\mathbf{Q}}[f(u,x)],\quad t=1,2,\dots$ (7)
with $\mathcal{P}^{K}_{t}$ an adaptive ambiguity set, with adaptive radius $\varepsilon^{K}_{t}\geq 0$, constructed from the data $\mathcal{D}_{t}$ observed so far. From here onward, we refer to the solution $x^{K}_{t}$ as the online solution, and the solution $x_{t}$ from (4) as the nominal DRO solution.
Definition 2.1 (Adaptive ambiguity set).
The adaptive ambiguity set at the end of time period $t$, with radius $\varepsilon^{K}_{t}\geq 0$ and reference distribution $\hat{\mathbf{P}}^{K}_{t}$, is given by
$\mathcal{P}^{K}_{t}=\mathbf{B}_{\varepsilon^{K}_{t}}^{r}(\hat{\mathbf{P}}^{K}_{t})=\left\{\mathbf{Q}\in\Xi_{r}({S})\mid W_{r}(\hat{\mathbf{P}}^{K}_{t},\mathbf{Q})\leq\varepsilon^{K}_{t}\right\}.$ (8)
Unlike the standard DRO ambiguity set, the adaptive ambiguity set is compressed in that we allow the reference distribution to be supported on only $K\leq n_{t}$ points. With access to the dataset $\mathcal{D}_{t}$, the reference distribution is formulated by clustering the datapoints into sets $C^{k}_{t}$, $k=1,\dots,K_{t},$ where $K_{t}\leq K$. The weight $\theta^{k}_{t}$ of a cluster $C^{k}_{t}$ is defined as the proportion of total points in the cluster, and the cluster mean $\bar{u}^{k}_{t}$ is the mean of these points. These values characterize the clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$i.e.,
$\theta_{t}^{k}=\frac{n^{k}_{t}}{n_{t}},\quad\bar{u}^{k}_{t}=\frac{1}{n^{t}_{k}}\sum_{\hat{u}\in C^{k}_{t}}\hat{u},\quad\hat{\mathbf{P}}^{K}_{t}=\sum_{k=1}^{K_{t}}\theta_{t}^{k}\delta_{\bar{u}_{t}^{k}},$
where $n^{k}_{t}=|C^{k}_{t}|$. In Section 3.1, we provide additional details on our clustering requirements, and in Section 5, provide efficient online clustering methods.
With an adaptive ambiguity set $\mathcal{P}^{K}_{t}$, we solve the online compressed DRO problem (7) through a direct reformulation approach [54, Section 2.4], given as
[…]
In addition to the primal variables $x\in\mathcal{X}$, we introduce dual and auxiliary variables $\lambda_{t}\in{\bf R}$, $s^{k}_{t}\in{\bf R}$, $z^{jk}_{t}\in{\bf R}^{d}$, and $y^{jk}_{t}\in{\bf R}^{d}$, for $k=1,\dots,K_{t}$ and $j=1,\dots,J$. Here, $f_{j}]^{*}(z,x)=\sup_{u\in\mathop{\bf dom}_{u}f}z^{T}u-[-f_{j}(u,x)]$ is the conjugate of $f_{j}$, $\sigma_{S}(y)=\sup_{u\in S}z^{T}u$ is the support function of $S\subseteq\mathbf{R}^{m}$, $\|\cdot\|_{*}$ is the dual norm of $\|\cdot\|$, and $\phi(q)=(q-1)^{(q-1)}/q^{q}$ for $q>1$ [33, Theorem 8]. Note that $q$ is the conjugate number of $r$, satisfying $1/r+1/q=1$, i.e., q=r/(r1)q=r/(r-1). The above reformulation is for $r<\infty$; the case for $r=\infty$ is given in [54, Appendix B]. We remark that the computational complexity of this problem is directly correlated to the value $K_{t}$, as it controls the number of constraints required.
In summary, the online problem follows a sequential process for $t\geq 0$:
  1. 1.
    We compute decision $x^{K}_{t}$ by solving (9) with ambiguity set $\mathcal{P}^{K}_{t}$;
  2. 2.
    We observe a new realization of the uncertainty $u$, denoted as $\hat{u}$;
  3. 3.
    We update the ambiguity set to $\mathcal{P}^{K}_{t+1}$ with the new data point.
Remark on the re-optimization frequency
Above, we proposed to update the ambiguity set and solve (9) after each new datapoint is collected. This could be adjusted, however, depending on the frequency of data collection and the complexity of (9). If the frequency is high, or if the data arrives in batches, step 1 (and 3, if desired) of the sequential process could instead be performed once a threshold number of datapoints has been observed. This does not affect the theoretical results in Section 3.3.2, and we do so in the example in Section 7. We can also choose to solve the robust problem only after a large enough adjustment to $\hat{{\mathbf{P}}}^{K}_{t}$ has been observed: if solving (9) is time-consuming relative to data-collection, we might wish to re-optimize only if a center $\bar{u}^{k}_{t}$ or a probability $\theta^{k}_{t}$ undergoes a change beyond a predetermined threshold $\delta>0$, i.e., when $|\bar{u}^{k}_{t}-\bar{u}^{k}_{\text{prev-solve}}|\geq\delta$ for some cluster $k$. Checking for this is at most $O(K)$ at each time step. Overall, since the computational complexity of our problem can be controlled by $K$, while the complexity of nominal DRO, and even SAA, increases over time, our method enjoys much higher flexibility in choosing a re-optimization threshold.
Our performance guarantees
To assess the performance of our online solutions, in Section 3.3 we provide finite-sample and asymptotic performance guarantees similar to (5), with adjustments due to clustering. A key strength of our approach is its ability to utilize established performance guarantees for Wasserstein DRO, with minimal additional assumptions. Informally, if we choose $\varepsilon_{t}^{K}=\varepsilon_{t}$ for a nominal radius $\varepsilon_{t}$ (e.g. from Theorem 2.1 or Theorem 2.2), we obtain
${\mathbf{P}}^{t}\left(H_{\star}\leq{\mbox{\bf E}}_{{\mathbf{P}}}[f(u,x^{K}_{t})]\leq H^{K}_{t}+\rho_{t}+\psi^{K}_{t}\leq H_{t}+\rho_{t}+\gamma^{K}_{t}\right)\geq 1-\beta_{t},$ (10)
where $\psi^{K}_{t}$ and $\gamma^{K}_{t}$ are additional residual terms of order $O(\varepsilon_{t}+d(\hat{{\mathbf{P}}}_{t},\hat{{\mathbf{P}}}_{t}^{K}))$, for some distance $d(\cdot)$. In other words, the difference between our certificates and the nominal DRO certificates converges, with rate $O(\varepsilon_{t})$, to a distance between the empirical and clustered empirical distributions. This rate depends on the radius $\varepsilon_{t}$, which we have shown to scale with $O(n_{t}^{-r/d})$(dimension-dependent), or with $\tilde{O}(n_{t}^{-1/2})$ for $r\in[1,2]$ and any $d$ (dimension-free). Our residual terms can be interpreted as the price for higher computational efficiency.
3 Clustering effect analysis
In this section, we formally show the effect of clustering on the finite samples and asymptotic performance guarantees for the online algorithm. In Section 3.1, we provide requirements on the online clustered distribution $\hat{{\mathbf{P}}}^{K}_{t}$, and in Section 3.2, provide theoretical results pertaining this distribution. In Section 3.3, we state our performance guarantees.
3.1 Clustering procedure requirements
The key of our approach is the adaptive ambiguity set $\mathcal{P}^{K}_{t}$, which is constructed using a compressed representation of the available data, $\hat{{\mathbf{P}}}^{K}_{t}$. Below, we formalize the requirements on generating $\hat{{\mathbf{P}}}^{K}_{t}$, and note that the decision-maker has the flexibility to use any method to generate $\hat{{\mathbf{P}}}^{K}_{t}$, provided it satisfies the given assumptions. These requirements ensure the validity of the asymptotic performance guarantees established later in this section, and can be satisfied by any partitioning of the support set $S$. In this section, we also outline the theoretical procedure to find the optimal clustering, and note that the goal of the online clustering procedure is to find a good approximation of the optimal clustering, while also maintaining low computational complexity.
3.1.1 Requirements
Let $K$ denote the maximum number of data points (that is, of constraints) that we can handle based on the available computational budget. We thus cluster data points together, and allow up to $K$ distinct clusters. For all methods and time periods $t$, we then maintain a set of $K_{t}\leq K$ clusters $\{C_{t}^{k}\}_{k=1}^{K_{t}}$ and a corresponding set of supports $\mathcal{S}_{t}=\{S_{t}^{k}\}_{k=1}^{K_{t}},$ where each one is defined as follows.
Definition 3.1 (Cluster support).
For each cluster $k$ and time $t$, the cluster support $S_{t}^{k}\subset S$ is defined as a region in $S$ such that any datapoint $\hat{u}\in\mathcal{D}_{t}$ that falls within this region is contained in cluster $C_{t}^{k}$. We also call these clusters the clustering induced by the cluster support.
The cluster supports must satisfy the following assumption.
Assumption 3.1.
The following holds for any time step $t$,
  • The elements in the set $\mathcal{S}_{t}=\{S_{t}^{k}\}_{k=1}^{K_{t}}$ have pairwise disjoint interiors, i.e., $\mathrm{int}(S^{k}_{t})\cap\mathrm{int}(S^{k^{\prime}}_{t})=\emptyset,~\forall S_{t}^{k},S_{t}^{k^{\prime}}\in\mathcal{S}_{t},k\neq k^{\prime}.$
  • The cluster supports cover $S$, i.e., $S=\cup_{k\leq K_{t}}S_{t}^{k}$.
  • The cluster supports have nonzero measure, i.e., ${\mathbf{P}}(S^{k}_{t})>0$ for $k=1,\dots,K_{t}$.
  • The cluster boundaries have measure 0.
Furthermore, there is some finite time $\tau<\infty$ such that $\mathcal{S}_{t}=\mathcal{S}_{\tau}$ for all $t\geq\tau$.
These assumptions ensure that all datapoints fall in a single cluster support, and could be satisfied by any partitioning of the original support set $S$. Note that until some finite time $\tau$, the datapoints are allowed to switch cluster assignments, which increases the flexibility of the clustering algorithm. By Assumption 3.1, the cluster weights $\sum_{k=1}^{K_{t}}\theta^{k}_{t}=1$, and $\hat{\mathbf{P}}^{K}_{t}$ represents an approximation of the non-clustered empirical distribution $\hat{\mathbf{P}}_{t}$, supported on $K_{t}\leq n_{t}$ atoms.
3.1.2 Optimal clustering
If the distribution $\mathbf{P}$} is known, then for any $K$, the optimal set of supports can be determined using $k$-centers clustering [20, Chapter 1].
Definition 3.2 ($k$-centers).
The $k$-centers problems, with respect to a distribution $\mathbf{Q}$ with support $Z$, and orders $p\geq 1$ (left), $p=\infty$ with compact $Z$ (right) are
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\inf_{A\subset{\bf R}^{d},|A|\leq K}{\mbox{\bf E}}_{\mathbf{Q}}\left[\min_{a\in A}\|u-a\|^{p}\right],~\quad\text{and}\quad\inf_{A\subset{\bf R}^{d},|A|\leq K}\sup_{u\in Z}\min_{a\in A}\|u-a\|,$} (11)
where $A^{\star}$ is the set that achieves this infimum.
Definition 3.3 (Minimum quantization error [20, Chapter 1]).
The minimum quantization error of a distribution $\mathbf{Q}$ with support $Z$ is the minimum Wasserstein-$p$ distance between $\mathbf{Q}$ and any discretization of $\mathbf{Q}$ to at most $K$ atoms, given as
$W^{K}_{\star,p}(\mathbf{Q})=\left({\mbox{\bf E}}_{\mathbf{Q}}\left[\min_{a\in A^{\star}}\|u-a\|^{p}\right]\right)^{1/{p}},\quad\text{and}\quad W^{K}_{\star,\infty}(\mathbf{Q})=\sup_{u\in Z}\min_{a\in A^{\star}}\|u-a\|,$
with $p\geq 1$ (left), $p=\infty$ with compact $Z$ (right), and $A^{\star}$ the set of optimal centers.
Specifically, we solve the $k$-centers problem to obtain an optimal set of centers $A^{\star}\subset{\bf R}^{d}$, then construct the Voronoi diagram of $A^{\star}$: each cluster support is defined as $S^{k}=\left\{u\in S\,\middle|\,\|u-a^{k}\|=\min_{a\in A^{\star}}\|u-a\|\right\},$ where each point $u$ is assigned to the region corresponding to its closest center $a^{k}$. This achieves the minimum quantization errori.e., $W^{K}_{\star,p}({\mathbf{P}})\leq W_{p}(\hat{{\mathbf{P}}}^{K}_{t},{\mathbf{P}})$ for any valid $\hat{{\mathbf{P}}}^{K}_{t}$.
Without knowledge of $\mathbf{P}$} but having access to a dataset $\mathcal{D}_{t}$, the optimal clustering with respect to the empirical distribution $\hat{\mathbf{P}}_{t}$ can also be found through the above process. However, the $k$-centers problem is NP-hard [26], and is therefore inefficient to solve. Furthermore, with $\hat{\mathbf{P}}_{t}$ as the reference distribution, when $t\rightarrow\infty$, the computational complexity of the $k$-centers problem increases, conflicting with our goal of decreasing computational effort. Thus, the goal of the online clustering algorithm is to find a good approximation of the $k$-centers problem. In Section 5, we give three algorithms that satisfy the given requirements.
3.2 Clustered distribution analysis
Since the performance bounds will be affected by the sequence of clustered empirical distributions $\hat{{\mathbf{P}}}^{K}_{t}$, we first derive some results pertaining these distributions. In particular, we formalize its distances, under different metrics, to the empirical distribution $\hat{{\mathbf{P}}}_{t}$, as well as prove its convergence to a limiting distribution ${\mathbf{P}}^{K}_{\star}$. We define the following values. Note that the orders and norms below are independent of the order and norm of the ambiguity set (8).
Definition 3.4.
For any clustered configuration of the dataset $\mathcal{D}_{t}$, with clusters $C^{k}_{t}$, cluster means $\bar{u}^{k}_{t}$, and clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$, we define the Wasserstein cost $W^{K}_{p}$ and clustering value $D^{K}_{t,p}$ as follows.
$W^{K}_{t,p}=W_{p}(\hat{{\mathbf{P}}}_{t},\hat{{\mathbf{P}}}^{K}_{t}),\quad D^{K}_{t,p}=\left(\frac{1}{n_{t}}\sum_{k=1}^{K_{t}}\sum_{\hat{u}\in C^{k}_{t}}\left\|\hat{u}-\bar{u}^{k}_{t}\right\|_{2}^{p}\right)^{1/{p}}.$
At each time step, we also define a function of dual variables $z^{jk}_{t}$ of (9),
$\Phi^{K}_{t}=\mathbf{1}\{J\geq 2\}\frac{1}{n_{t}}\sum_{k=1}^{K_{t}}\sum_{\hat{u}\in C^{k}_{t}}\max_{j\leq J}(-z_{t}^{{jk}})^{T}(\hat{u}-\bar{u}^{k}_{t}),$
where $\mathbf{1}\{\cdot\}$ is the 0-11 indicator function. Note that $\Phi^{K}_{t}=0$ when $J=1$, i.e., when $f$ is a concave function.
By construction, the above values all quantify the distance between the clustered and the non-clustered empirical distributions; the first two are distance metrics, while the third is a nonnegative cost function. We establish a hierarchy between these distances and provide a few bounds, making use of the following definition.
Definition 3.5.
The diameter of a set $S$ is defined
$\mathrm{diam}_{q}(S)=\max\{\|u-v\|_{q}\mid u,v\in S\}.$
Proposition 3.1.
For all $K$, $t$, and $p$, with $\|\cdot\|$ the $\ell_{2}$-norm, $W^{K}_{\star,p}(\hat{{\mathbf{P}}}_{t})\leq W^{K}_{t,p}\leq D^{K}_{t,p}.$ Under Assumption 3.1, if $K\geq n_{t}$, $D^{K}_{t,p}=W^{K}_{t,p}=\Phi^{K}_{t}=0.$ In addition, if $S$ is bounded, with radius $\rho=(1/2)\mathrm{diam}_{\infty}(S)<\infty$, we have that for all $K$ and $t$, and any $\|\cdot\|$,
$D^{K}_{t,p}\leq 2\rho,\quad\Phi^{K}_{t}\leq\mathbf{1}\{J\geq 2\}\max_{k\leq{K_{t}}}\max_{j\leq J}2\rho\|z^{jk}_{t}\|.$
Proof.
The Wasserstein distance is computed with the optimal (minimizing) coupling (joint distribution) between the two reference distributions, while $D^{K}_{t,p}$ is calculated from one particular coupling. Therefore, $W^{K}_{t,p}\leq D^{K}_{t,p}$. The lower bound on $W^{K}_{t,p}$ follows by Definition 3.3, and the equalities follow from the fact that $\hat{{\mathbf{P}}}^{K}_{t}=\hat{{\mathbf{P}}}_{t}$ when $K\geq n_{t}$. The final inequalities follow from the boundedness of the support. ∎
Furthermore, we can derive asymptotic bounds, for $t\to\infty$. This relies on Assumption 3.1, that for $t\geq\tau$, where $\tau<\infty$ is some finite time, we fix the cluster supports $S_{t}^{k}$. We make use of the following lemma.
Lemma 3.1 (Clustering consistency).
[…]
${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}W_{p}({\mathbf{P}}^{K}_{\star},\hat{{\mathbf{P}}}^{K}_{t})=0\right\}=1.$
Proof.
W.l.o.g., let $K_{\tau}=K$. By Assumption 3.1 and the Strong Law of Large Numbers (SLLN) [31], for each cluster $k\leq K$,
${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\mathbf{1}\{\hat{u}^{i}\in S^{k}_{\tau}\}=\theta_{\star}^{k}\right\}=1.$
This shows that the cluster weights converge almost surely to ${\mathbf{P}}(S^{k}_{\tau})=\theta_{\star}^{k}$, the true weights on each cluster support. Next, as ${\mathbf{P}}(S^{k}_{\tau})>0$ for all clusters $k$, we note that $n^{k}_{t}\to\infty$ when $t\to\infty$. Furthermore, since $\mathbf{P}$} has finite $p$-th moments, the conditional distributions of $\mathbf{P}$} on these supports $S^{k}_{\tau}$ must also have finite $p$-th moments. We denote these conditional distributions ${\mathbf{P}}^{k}$, and their means $\bar{u}^{k}_{\star}$. Then, by the SLLN,
${\mathbf{P}}^{\infty}\left\{\lim_{n_{t}^{k}\to\infty}\frac{1}{n_{t}^{k}}\sum_{\hat{u}\in C^{k}_{\tau}}\hat{u}=\bar{u}^{k}_{\star}\right\}=1,$
i.e., the cluster means converge almost surely to the true (finite) means of the conditional distributions ${\mathbf{P}}^{k}$. Since both the probabilities and the means converge almost surely to finite limits, it follows that
${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}\hat{{\mathbf{P}}}^{K}_{t}=\lim_{t\to\infty}\sum_{k=1}^{K}\theta^{k}_{t}\delta_{\bar{u}_{t}^{k}}=\sum_{k=1}^{K}\theta^{k}_{\star}\delta_{\bar{u}_{\star}^{k}}={\mathbf{P}}^{K}_{\star}\right\}=1,$
where ${\mathbf{P}}^{K}_{\star}$ is constructed to be a discretized distribution of $\mathbf{P}$} onto $K$ atoms. Since ${\mathbf{P}}^{K}_{\star}$ has finite $p$-th moments by construction, this implies almost sure convergence with respect to the Wasserstein-$p$ metric [53, Theorem 6.9]. ∎
We can then give an asymptotic bound on $d^{K}_{t,p}$ for $t\to\infty$.
Theorem 3.1 (Asymptotic clustering bounds).
Under Lemma 3.1, we have
$\mathbf{P}^{\infty}\left\{W^{K}_{\star,p}({\mathbf{P}})\leq\lim_{t\rightarrow\infty}W^{K}_{t,p}=W^{K}_{\star,p}=W_{p}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\right\}=1.$
Furthermore, $\lim_{K\rightarrow\infty}W_{p}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})=0$ for all $p$.
Proof.
By the continuity of the Wasserstein-$p$ metric, and by the almost-sure convergence of the empirical distributions to the respective true distributions (Theorem 2.1 and Lemma 3.1), we obtain $\mathbf{P}^{\infty}\{\lim_{t\rightarrow\infty}W^{K}_{t,p}=W_{p}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\}=1$ [52, Theorem 7.12]. For the lower bound, recall that by Definition 3.3, $W^{K}_{\star,p}(\hat{{\mathbf{P}}}_{t})\leq W^{K}_{t,p}.$ The result then holds by the asymptotic consistency of finite quantizations, given in [20, Corollary 4.24]. Specifically, we note that $\mathbf{P}^{\infty}\{\lim_{t\rightarrow\infty}W^{K}_{\star,p}(\hat{{\mathbf{P}}}_{t})=W^{K}_{\star,p}({\mathbf{P}})\}=1.$ Lastly, the asymptotic result for $K\to\infty$ holds from [20, Lemma 6.1]. ∎
The asymptotic limit of $D^{K}_{t,p}$ and the limit superior of $\Phi^{K}_{t}$ follows similarly.
Theorem 3.2 (Asymptotic clustering values).
Under Lemma 3.1, we have that
${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}D^{K}_{t,p}=D^{K}_{\star,p}=\left(\mathbf{E}_{{\mathbf{P}}}\left[\sum_{k=1}^{K}\mathbf{1}\{u\in S^{k}_{\tau}\}\|u-\bar{u}^{k}_{\star}\|^{p}_{2}\right]\right)^{1/{p}}\right\}=1.$
Furthermore, if there exist accumulation points of $z_{t}^{jk}$ for all $j\leq J,k\leq K$, then
${\mathbf{P}}^{\infty}\left\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\limsup_{t\to\infty}}~\Phi^{K}_{t}=\Phi^{K}_{\star}=\mathbf{1}\{J\geq 2\}\mathbf{E}_{{\mathbf{P}}}\left[\sum_{k=1}^{K}\mathbf{1}\{u\in S^{k}_{\tau}\}\max_{j\leq J}(-z^{jk}_{\star})^{T}(u-\bar{u}^{k}_{\star})\right]\right\}=1,$
where $\hat{z}^{jk}_{\star}$ are the accumulation points that achieve the limit superior.
Proof.
By the consistency of our clustering procedure (Lemma 3.1), the 0-measure of the cluster boundaries (Assumption 3.1), the continuity of linear functions and $\|\cdot\|^{p}$, and the existence of the accumulation points, the limit (or limit superior, for the second result) of the functions within the summations of Definition 3.4 converge almost surely to the functions within the expectations above, for $\mathbf{P}$}-almost every $u$. That is, defining $\omega_{t}(u)=(\sum_{k=1}^{K_{t}}\mathbf{1}\{u\in C^{k}_{t}\}\|u-\bar{u}^{k}_{t}\|_{2}^{p})^{1/{p}}$ for the first result, by the given assumptions, we have that $\lim_{t\to\infty}\omega_{t}(u)=\sum_{k=1}^{K}\mathbf{1}\{u\in S^{k}_{\tau}\}\|u-\bar{u}^{k}_{\star}\|^{p}_{2}=\omega_{\star}(u)$, for $\mathbf{P}$}-almost every $u$. Next, to conclude that $\lim_{t\to\infty}\mathbf{E}_{\hat{{\mathbf{P}}}_{t}}[\omega_{t}(u)]=\mathbf{E}_{{\mathbf{P}}}[\omega_{\star}(u)]$, we require the function $\omega_{t}(u)$ to have an integrable dominating function. W.l.o.g., let $u$ belong in the cluster $k$. By the triangle inequality, $|\omega_{t}(u)|=\|u-\bar{u}^{k}_{t}\|^{p}_{2}\leq(\|u\|_{2}+\|\bar{u}\|^{k}_{t})^{p}\leq\bar{C}(1+\|u\|_{2}^{p})$ for some constant $\bar{C}$. By the assumption of finite $p$-th moments, this dominating function is integrable. The result then follows from [52, Theorem 7.12(iv)], which establishes a generalized convergence of expectation. Since we have shown almost sure pointwise convergence in the first step, the cited theorem holds even without the assumption on continuity.
From Lemma 3.1, we established that the clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$ converges almost surely to the distribution ${\mathbf{P}}^{K}_{\star}$. We can also prove that its convergence rate follows the convergence rate of the non-clustered empirical distribution.
Theorem 3.3 (Clustering convergence rate).
Suppose Assumption 3.1 holds, and we have a sequence $\varepsilon_{t}$, computed using Theorem 2.1 for a order $p$, such that $\lim_{t\rightarrow\infty}\varepsilon_{t}=0$ and the confidence sequence $\beta_{t}\in(0,1)$ satisfies $\sum_{t=1}^{\infty}{\beta_{t}<\infty}$. Then, for $t\geq\tau$, the distance $W_{p}(\hat{{\mathbf{P}}}^{K}_{t},{\mathbf{P}}^{K}_{\star})$ converges to 0 with rate $O(\varepsilon_{t})$.
Proof.
By assumption, all cluster supports have nonzero measure, i.e., ${\mathbf{P}}(S^{k}_{\tau})>0$. Since $\mathbf{P}$} is light-tailed, we also know that for some $\alpha>1$, there is some finite $R<\infty$ defined by Assumption 2.3. It follows that the conditional distributions of $\mathbf{P}$} on the supports $S^{k}_{\tau}$, denoted ${\mathbf{P}}^{k}$, are also light-tailed, with the same $\alpha$ and a constant $R^{k}\leq O(R)$, i.e., $R^{k}=\int_{S_{\tau}^{k}}\exp(\|u\|^{\alpha}){\mathbf{P}}(u|u\in S^{k}_{\tau})(du)<\infty.$ By Theorem 2.1,
${\mathbf{P}}^{t}(W_{p}({\mathbf{P}}^{k},\hat{{\mathbf{P}}}^{t,k})\leq O(\varepsilon_{t}))\geq 1-\beta_{t},$
where $\hat{{\mathbf{P}}}^{t,k}$ is the conditional distribution of $\hat{{\mathbf{P}}}_{t}$ on support $S^{k}_{\tau}$. By Kantorovich-Rubinstein duality and the ordering of Wasserstein distances,
$\left|{\mbox{\bf E}}_{{\mathbf{P}}^{k}}[u_{i}]-{\mbox{\bf E}}_{\hat{{\mathbf{P}}}^{k,t}}[u_{i}]\right|\leq W_{1}({\mathbf{P}}^{k},\hat{{\mathbf{P}}}^{k,t})\leq W_{p}({\mathbf{P}}^{k},\hat{{\mathbf{P}}}^{k,t}),\quad i=1,\dots,d.$
This implies that the means $\bar{u}^{k}_{t}$ converge to the true means $\bar{u}^{k}_{\star}$ with rate $O(\varepsilon_{t})$, with respect to the infinity norm, i.e., $\|\bar{u}^{k}_{t}-\bar{u}^{k}_{\star}\|_{\infty}\leq O(\varepsilon_{t}).$ Then, by the equivalence of norms on ${\bf R}^{d}$ and the definition of the Wasserstein distance on discrete distributions,
$W_{p}(\hat{{\mathbf{P}}}^{K}_{t},{\mathbf{P}}^{K}_{\star})\leq\left(\sum_{k=1}^{K}|\theta^{k}_{t}-\theta^{k}_{\star}|\|\bar{u}^{k}_{t}-\bar{u}^{k}_{\star}\|^{p}\right)^{1/{p}}\leq O(\varepsilon_{t}),$
since the weights $\theta$ are bounded. ∎
3.3 Performance guarantees
We can now derive performance guarantees for the online problem. We begin by summarizing the assumptions.
Assumption 3.2.
Fix an order $r$ and let the following hold.
  1. 1.
    For the chosen order $r$ and all $t$, $\varepsilon_{t},\beta_{t}$, and $\rho_{t}$ are constructed according to either Theorem 2.1 or Theorem 2.2, under their respective assumptions.
  2. 2.
    The distribution $\hat{{\mathbf{P}}}^{K}_{t}$ is constructed using $\mathcal{S}^{k}$ that satisfy Assumption 3.1.
  3. 3.
    The adaptive ambiguity set $\mathcal{P}^{K}_{t}=\mathbf{B}_{\varepsilon^{K}_{t}}^{r}(\hat{\mathbf{P}}^{K}_{t})$, with $\varepsilon^{K}_{t}=\varepsilon_{t}$ and the Wasserstein distance defined with an arbitrary norm $\|\cdot\|$.
3.3.1 Concave case
We first present a simplified set of results for a concave objective function $f$, with unbounded support $S={\bf R}^{d}$. The proofs can be derived from the general case, which will be presented subsequently.
Theorem 3.4 (Concave finite sample guarantee).
Suppose Assumptions 2.12.2, and 3.2 hold, $f$ is concave in $u$, and $S={\bf R}^{d}$. Then, the optimal solution $x^{K}_{t}$ and the optimal value $H^{K}_{t}$ of the compressed DRO problem (7) at time $t$ satisfies
$H_{t}\leq H_{t}^{K}\leq H_{t}+\bar{\psi}_{t}^{K},\quad\bar{\psi}_{t}^{K}=\min\left\{{\frac{L}{2}}(D^{K}_{t,2})^{2},M(2\varepsilon_{t}+W^{K}_{t,1})\right\}$ (12)
where $H_{t}$ is the nominal DRO value. Furthermore, we obtain the finite-sample probabilistic guarantee
${\mathbf{P}}^{{t}}\left(H_{\star}\leq\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]\leq H^{K}_{t}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho_{t}}\right)\geq 1-\beta_{t},$ (13)
where the certificate $H^{K}_{t}+\rho_{t}$ satisfies
${\mathbf{P}}^{{t}}\left(H_{\star}\leq H_{t}+\rho_{t}\leq H^{K}_{t}+\rho_{t}\leq H_{t}+\rho_{t}+\bar{\psi}_{t}^{K}\right)\geq 1-\beta_{t}.$ (14)
If any among the Lipschitz and smoothness conditions from Assumption 2.1 does not hold, the corresponding term in $\bar{\psi}_{t}^{K}$ is set to $\infty$.
Theorem 3.4 shows that, when the adaptive ambiguity set $\mathcal{P}^{K}_{t}$ is constructed with the same radius as nominal DRO, we can relate their optimal values in terms of this radius and the distances given in Definition 3.4. This yields a finite-sample guarantee, with a certificate within a bounded distance from the nominal DRO certificate.
In particular, equation (13) states that with probability $1-\beta_{t}$, the out-of-sample performance $\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]$ upper bounds the true optimal value $H_{\star}$, and this value is in turn upper bounded by the in-sample objective $H^{K}_{t}$ and the nominal DRO residual $\rho_{t}$, which depends on the radius $\varepsilon_{t}$. Using the bounds established in (12), equation (14) relates this certificate to the one obtained by nominal DRO, providing upper and lower bounds. Specifically, the certificate is an upper bound on the nominal certificate $H_{t}+\rho_{t}$, but the increase in suboptimality (as we are solving a minimization problem, a larger certificate is less optimal) is at most $\bar{\psi}_{t}^{K}$. Similar to the nominal residual, which is the price paid in exchange for smaller radii, $\bar{\psi}_{t}^{K}$ is the price we pay for increased computational efficiency. As the radius $\varepsilon_{t}$ decreases over time at the chosen rate, which could be as fast as $\tilde{O}(n_{t}^{-1/2})$, this price converges to, or already is, a function of only the clustering distance. Note that the distances for the clustering discrepancy are defined with orders $p=1$ or $p=2$, unrelated to the order $r$ of the ambiguity set. The order $r$ instead affects the bounds through the radius $\varepsilon_{t}$ and the residual $\rho_{t}$.
Depending on the smoothness and Lipschitz conditions on the function $f$, we can choose the best out of the two clauses given for $\bar{\psi}_{t}^{K}$. Note that when $f$ is affine, $\bar{\psi}_{t}^{K}$ reduces to 0, and we recover nominal DRO guarantees regardless of $K$. Otherwise, when the Lipchitz condition is satisfied, we can compute the bound involving $W^{K}_{t,1}$. This is the Wasserstein-1 distance between two empirical distributions, and can be calculated using a linear program. When the smoothness condition is satisfied, $D_{t,2}^{K}$ can be easily computed using the empirical distributions.
Asymptotically, if $\varepsilon_{t}$ is chosen through Theorem 2.1, the bound converges as in the following theorem. Note that while $\varepsilon_{t}$ chosen through Theorem 2.2 allows smaller finite-sample differences between the certificates, i.e., $\tilde{O}(n_{t}^{-1/2})$ scales better than $O(n_{t}^{-p/d})$, in the asymptotic case we can assume $O(n_{t}^{-p/d})$ to converge to 0 for any $d$. In fact, our method is more compatible with asymptotic analysis than nominal DRO, as our computational complexity is fixed regardless of the sample size.
Theorem 3.5 (Concave asymptotic guarantee).
Suppose Assumptions 2.12.23.1 and 3.2 hold, $f$ is concave, $S={\bf R}^{d}$, and $\varepsilon_{t},\beta_{t}$ are constructed according to Theorem 2.1 for all $t$. Asymptotically, the sequence of optimal solutions $x^{K}_{t}$ and optimal values $H^{K}_{t}$ of the compressed DRO problem (7) satisfies
$\mathbf{P}^{\infty}\left\{H_{\star}\leq\limsup_{t\rightarrow\infty}\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]\ \leq\limsup_{t\rightarrow\infty}H^{K}_{t}\leq H_{\star}+\bar{\psi}_{\star}^{K}\right\}=1,$
$\bar{\psi}_{\star}^{K}$ $=\min\left\{{\frac{L}{2}}(D^{K}_{\star,2})^{2},M{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}W^{K}_{\star,1}}\right\}.$
This shows that in the asymptotic case, the performance gap between the online solution and the nominal DRO solution is controlled by our defined distances between the true distribution and a chosen discretization to $K$ atoms. It follows that a well-chosen clustering algorithm and a balanced $K$ are crucial for good performance.
3.3.2 General case
We now extend these results to the maximum-of-concave case with arbitrary support $S$, which requires the following lemma. The proofs of this lemma and the upcoming results are in Appendix A.1.
Lemma 3.2 (Effect of the support).
Let $\Delta_{t}=\tilde{H}_{t}-H_{t}$, where $H_{t}$ is the value of the DRO problem with arbitrary support $S$, while $\tilde{H}_{t}$ fixes $S={\bf R}^{d}$, i.e., $\tilde{H}_{t}=\max_{\mathbf{Q}\in\hat{\mathcal{P}}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x_{t})]$, $\hat{\mathcal{P}}_{t}=\{\mathbf{Q}\in\Xi_{r}({{\bf R}^{d}})\mid W_{r}(\hat{{\mathbf{P}}}_{t},\mathbf{Q})\leq\varepsilon_{t}\}.$ If the sequence of $\varepsilon_{t}$ is constructed according to Theorem 2.1, we have ${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}\Delta_{t}=0\right\}=1.$ For any $\varepsilon_{t}\leq 1$, if $f$ satisfies clause 1 or 3 of Assumption 2.1, with $r\geq 2$ in the latter case, then $\Delta_{t}=O(\varepsilon_{t})$.
This lemma formalizes the discrepancy between nominal DRO objectives with and without support constraints. This discrepancy appears in the following bounds.
Theorem 3.6 (General finite sample guarantee).
Suppose Assumptions 2.12.2, and 3.2 hold. Then, the optimal solution $x^{K}_{t}$ and the optimal value $H^{K}_{t}$ of the compressed DRO problem (7) at time $t$ satisfies
$\underline{\psi}_{t}^{K}\leq H_{t}^{K}-H_{t}\leq\bar{\psi}_{t}^{K},$ (15)
where $H_{t}$ is the nominal DRO value, and the bounds are defined as
$\underline{\psi}_{t}^{K}$ $=\min\left\{\Phi^{K}_{t},\max_{j\leq J}M_{j}(2\varepsilon_{t}+W^{K}_{t,1}),{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\nabla f_{x^{K}_{t}}\|_{L_{2}(\hat{{\mathbf{P}}}^{K}_{t})}(2\varepsilon_{t}+W^{K}_{t,2})+\frac{L}{2}(2\varepsilon_{t}+W^{K}_{t,2})^{2}}\right\},$
$\bar{\psi}_{t}^{K}$ $=\min\left\{\Delta_{t}+{\max_{j\leq J}\frac{L_{j}}{2}}(D^{K}_{t,2})^{2},\max_{j\leq J}M_{j}(2\varepsilon_{t}+W^{K}_{t,1})\right\},$
with $\|\nabla f_{x}\|_{L_{2}(\mathbf{Q})}=(\int\|\nabla f(u,x)\|^{2}d\mathbf{Q})^{1/2},$ $\Phi^{K}_{t}$ given in Definition 3.4, and $\Delta_{t}$ given in Lemma 3.2. Furthermore, we obtain the finite-sample probabilistic guarantee
${\mathbf{P}}^{{t}}\left(H_{\star}\leq\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]\leq H^{K}_{t}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho_{t}}+\underline{\psi}_{t}^{K}\right)\geq 1-\beta_{t},$ (16)
where the certificate $H^{K}_{t}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\rho_{t}}+\underline{\psi}_{t}^{K}$ satisfies
${\mathbf{P}}^{{t}}\left(H_{\star}\leq H_{t}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\rho_{t}}\leq H^{K}_{t}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\rho_{t}}+\underline{\psi}_{t}^{K}\leq H_{t}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\rho_{t}}+\underline{\psi}_{t}^{K}+\bar{\psi}_{t}^{K}\right)\geq 1-\beta_{t}.$ (17)
If any among the Lipschitz and smoothness conditions from Assumption 2.1 does not hold, the corresponding term(s) in the bounds are set to $\infty$.
Theorem 3.6 shows that, for maximum-of-concave $f$ with arbitrary support, the online objective $H^{K}_{t}$ is no longer guaranteed to be an upper bound on $H_{t}$. This leads to performance guarantees and performance gaps with additional terms compared to that of Theorem 3.4. However, the additional terms are also functions of the ambiguity set radius $\varepsilon_{t}$ and the clustering distances, and can be readily approximated or computed. For instance, $\Delta_{t}\approx O(\varepsilon_{t})$, and the value $\Phi^{K}_{t}$ can be computed using the dual variables of the solved optimization problem. Asymptotically, we can also obtain limits for the generalized bounds.
Theorem 3.7 (General asymptotic guarantee).
Suppose Assumptions 2.12.23.1 and 3.2 hold, and $\varepsilon_{t},\beta_{t}$ are constructed according to Theorem 2.1 for all $t$. Asymptotically, the sequence of optimal solutions $x^{K}_{t}$ and optimal values $H^{K}_{t}$ of the compressed DRO problem (7) satisfies
$\mathbf{P}^{\infty}\left\{H_{\star}\leq\limsup_{t\rightarrow\infty}\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]\ \leq\limsup_{t\rightarrow\infty}H^{K}_{t}+\underline{\psi}_{\star}^{K}\leq H_{\star}+\underline{\psi}_{\star}^{K}+\bar{\psi}_{\star}^{K}\right\}=1,$
$\underline{\psi}_{\star}^{K}$ $=\min\left\{\limsup_{t\rightarrow\infty}{\Phi^{K}_{t}},\max_{j\leq J}M_{j}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}W^{K}_{\star,1}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\nabla f_{\hat{x}^{K}_{\star}}\|_{L_{2}({\mathbf{P}}^{K}_{\star})}W^{K}_{\star,2}+\frac{L}{2}(W^{K}_{\star,2})^{2}}\right\},$
$\bar{\psi}_{\star}^{K}$ $=\min\left\{{\max_{j\leq J}\frac{L_{j}}{2}}(D^{K}_{\star,2})^{2},\max_{j\leq J}M_{j}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}W^{K}_{\star,1}}\right\},$
where $\hat{x}^{K}_{\star}$ is the accumulation point of $x^{K}_{t}$ that yields the largest value of the bound.
Proof.
This follows from the summability of $\beta_{t}$ and the Borel–Cantelli Lemma [28, Theorem 2.18], as well as Lemma 3.2. The convergence of the bounds follow from the results in Section 3.2. The existence of $\hat{x}^{K}_{\star}$ follows from the compactness of $\mathcal{X}$.
In order to obtain a closed-form limit superior for $\Phi^{K}_{t}$}, we require the convergence of the online algorithm. Below, we prove that the online optimal value converges to a clustering-dependent value, at the same rate of convergence as nominal DRO.
Theorem 3.8.
Suppose Assumption 3.1 and the Lipschitz clause of Assumption 2.1 holds, and let $\varepsilon^{K}_{t}\sim O(\varepsilon_{t})$, where $\varepsilon_{t}$ is computed according to Theorem 2.1. With this radius, the optimal value $H^{K}_{t}$ of the compressed DRO problem (7) at time $t\geq\tau$ satisfies
${\mathbf{P}}^{{t}}\left(H^{K}_{\star}\leq{\mbox{\bf E}}_{{\mathbf{P}}^{K}_{\star}}[f(u,x^{K}_{t})]\leq H^{K}_{t}\right)\geq 1-\beta_{t},$
[…]
Proof.
This follows from Theorem 3.3 and an adaptation of [38, Theorem 3.6], by treating ${\mathbf{P}}^{K}_{\star}$ as the true distribution. ∎
With this result, for special cases, we can obtain a limit superior for $\Phi^{K}_{t}$.
Theorem 3.9.
Suppose the assumptions of Theorem 3.8 hold. When the function $f(u,x)$ is maximum-of-affine in both $u$ and $x$, and the support $S={\bf R}^{d}$, we have $\limsup_{t\to\infty}\Phi^{K}_{t}=\Phi^{K}_{\star},$ where the latter is defined in Theorem 3.2.
Proof.
When $f$ is maximum-of-affine and $S={\bf R}^{d}$, well known conjugation results show that the variables $z^{jk}=P^{jk}x$, for constant matrices $P^{jk}$ [38, Corollary 5.1]. It follows that the accumulation points $\hat{z}^{jk}_{\star}$ exist as continuous functions of an accumulation point $\hat{x}^{K}_{\star}$, which exist by Theorem 3.8.
The above results hold in general, for any clustering algorithm satisfying Assumption 3.1. For completeness, below we instead assume optimal clustering, and state the best bounds we can obtain.
Theorem 3.10 (Optimal clustering bounds).
Suppose Assumptions 2.12.23.1 and 3.2 hold. If the clustering algorithm is optimal with respect to $p=2$ and the $\ell_{2}$-norm, i.e., for all $t$, we achieve the minimum quantization error of $\hat{{\mathbf{P}}}_{t}$ on $K$ clusters, such that $D^{K}_{t,2}=W^{K}_{t,2}=W^{K}_{\star,2}(\hat{{\mathbf{P}}}_{t}),$ then almost surely with respect to $\mathbf{P}$$^{\infty}$,
$\bar{\psi}_{\star}^{K}=\min\left\{{\max_{j\leq J}\frac{L_{j}}{2}}(W^{K}_{\star,2}({\mathbf{P}}))^{2},\max_{j\leq J}M_{j}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}W^{K}_{\star,1}}\right\}.$
When $f(u,x)$ is maximum-of-affine in both $u$ and $x$, and the support $S={\bf R}^{d}$, if the clustering-induced coupling is optimal with respect to $p=1$, i.e., $D^{K}_{t,1}=W^{K}_{t,1},$ then almost surely with respect to $\mathbf{P}$$^{\infty}$,
${\underline{\psi}_{\star}^{K}\leq\Phi^{K}_{\star}}\leq\max_{j\leq J}M_{j}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}W^{K}_{\star,1}}.$
Proof.
The first result follows from Proposition 3.1 and Theorem 3.1, where the pertinent inequalities are made tight by the assumption of optimal clustering. For the second result, note that $\Phi^{K}_{t}$ is computed with respect to the clustering-induced coupling, i.e., each $\hat{u}$ is associated with the mean $\bar{u}^{k}_{t}$ of the cluster it belongs to. If this is equivalent to the optimal coupling, then by the Cauchy-Schwarz inequality,
$(-z^{jk}_{t})^{T}(\hat{u}-\bar{u}^{k}_{t})\leq\|z^{jk}_{t}\|\|\hat{u}-\bar{u}^{k}_{t}\|\leq M_{j}\|\hat{u}-\bar{u}^{k}_{t}\|=M_{j}\min\left\{\|\hat{u}-\bar{u}\|\mid\bar{u}\in\{\bar{u}^{k}_{t}\}_{k=1}^{K}\right\},$
where $\|z^{jk}_{t}\|\leq M_{j}$ $\forall j\leq J$, $k\leq K$, and $t\geq 1$ by the conjugates of maximum-of-affine functions and the compactness of $\mathcal{X}$. This holds for all $t$, $j$, and $\hat{u}$, which implies
$\Phi^{K}_{t}\leq\max_{j\leq J}M_{j}\frac{1}{n_{t}}\sum_{k=1}^{K}\sum_{\hat{u}\in C^{k}_{t}}\|\hat{u}-\bar{u}^{k}_{t}\|=\max_{j\leq J}M_{j}W^{K}_{t,1}.$
The result follows by taking the limit superior as $t\to\infty$.
In all cases, our bounds scale with the clustering quality; when the clustering algorithm is optimized, the bound is tighter. Theorem 3.8 shows a direct correlation between the optimality of the clustering procedure and the optimality of the online solution; the algorithm will converge to a value dependent on the supports of the final clustering, so it follows that a well-chosen clustering procedure will lead to a better performing solution. For all these bounds, we can verify using either Proposition 3.1 or Theorem 3.1 that as the number of clusters $K\to\infty$, we recover the DRO performance guarantees. This illustrates the importance of finding a $K<\infty$ which balances the tradeoff between performance and computational effort.
In fact, this leads to an interesting result: if we do not have a strict budget $K$, but would like to decrease computational effort, we can obtain the following more controllable bounds.
Theorem 3.11 (Fixed radii clustering).
[…]
$H_{\star}-2\max_{j\leq J}M_{j}\eta\leq\limsup_{t\rightarrow\infty}H^{K_{t}}_{t}\leq H_{\star}+2\min\left\{{\max_{j\leq J}L_{j}}\eta^{2},\max_{j\leq J}M_{j}\eta\right\}.$
Proof.
Under this clustering regime, the distance $W_{p}(\hat{{\mathbf{P}}}_{t},\hat{{\mathbf{P}}}^{K_{t}}_{t})\leq 2\eta$, where $K_{t}$ is allowed to increase with $t$. The result then follows from Theorem 3.6. ∎
4 Special case: bounded support
In the dimension-dependent portion of Section 2.2, we gave performance guarantees for light-tailed supports $S$, and set the ambiguity set radius $\varepsilon^{K}_{t}$ with the nominal DRO radius $\varepsilon_{t}$, calculated using Theorem 2.1. In this section, for the case of bounded support, we give an explicit formulation of the radius in terms of the diameter of the support $S$. We also derive performance guarantees using a different theoretical approach from that of Section 3.3: using measure concentration results, we instead increase the ambiguity set radius by a finite value. In this case, the online objective value $H^{K}_{t}$ is guaranteed to be an upper bound on the DRO objective value $H_{t}$, regardless of the curvature of $f$.
4.1 Wasserstein radius with bounded support
We begin with calculating the ambiguity set radius for the case of bounded support. Suppose the following assumption holds true.
Assumption 4.1 (Compact support).
The true distribution $\mathbf{P}$ has compact support $S\subset\mathbf{R}^{d}$.
Then, the result follows from propositions 24 and 25 of the online version of [11], which makes use of the concentration bound below.
Proposition 4.1 (Concentration bound).
Let Assumption 4.1 hold, and let $p<d/2$. Then, for the true distribution $\mathbf{P}$} and an empirical distribution $\hat{{\mathbf{P}}}^{N}$ supported on $N$ points,
$\mathbf{P}^{N}\left(W_{p}\left(\hat{\mathbf{P}}^{N},\mathbf{P}\right)\geq\mathbf{E}\left[W_{p}\left(\hat{\mathbf{P}}^{N},\mathbf{P}\right)\right]+t\right)\leq e^{-Nt^{2p}/\left(2\tilde{\rho}^{2p}\right)}\quad\forall t\geq 0,$
where $\tilde{\rho}=\mathrm{diam}_{2}(S)$.
This follows from [10, Proposition A.2] and exploits the fact that bounded distributions are sub-Gaussians. Then, by bounding the expected Wasserstein distance using a function of $d,p,\tilde{\rho}$, and $N$, we obtain the following explicit characterization of the nominal ambiguity radius. For details, see the proofs in the online version of the referenced paper.
Theorem 4.1 (Explicit nominal radius [11][Propositions 23, 24]).
Assume that the probability measure $\mathbf{P}$ is supported on $S\subset\mathbf{R}^{d}$ with $\rho=\frac{1}{2}\operatorname{diam}_{\infty}(S)<\infty$ and that $p<d/2$. Then, for a given confidence $\beta\in(0,1)$, the nominal ambiguity radius is given by
$\varepsilon_{N}(\beta,\rho)=2\rho\left(CN^{-\frac{1}{d}}+\sqrt{d}\left(2\ln\beta^{-1}\right)^{\frac{1}{2p}}N^{-\frac{1}{2p}}\right),$
where
$C=\sqrt{d}2^{(d-2)/(2p)}\left(\frac{1}{1-2^{p-d/2}}+\frac{1}{1-2^{-p}}\right)^{1/p}.$
Furthermore, the radius can be written with the compact form
$\varepsilon_{N}(\beta,\rho)=2\rho\left(\frac{\ln\left(C^{\star}\beta^{-1}\right)}{c^{\star}}\right)^{\frac{1}{d}}N^{-\frac{1}{d}},$
with $C^{\star}=C^{d}/(2\sqrt{d}^{d})$ and $c^{\star}=1/(2^{d}\sqrt{d}^{d})$.
For any time step $t$ and $\beta_{t}$, the radius $\varepsilon_{t}$ calculated from Theorem 4.1 can be used to obtain the performance guarantees in Theorem 2.1, since distributions with bounded support are trivially light-tailed.
On the other hand, we can also use this radius to obtain performance guarantees specifically for bounded distributions. In the following, we show that by suitably enlarging the radius $\varepsilon_{t}$, we can also obtain performance guarantees for the online compressed DRO problem.
4.2 Performance guarantees with bounded support
In this approach, we assume the clustering algorithm to satisfy the requirements in Section 3.1, and additionally restrict the diameter of each cluster to a value $2\eta_{K}$. In Section 5.1, we give such an algorithm, which is based on covering the bounded support with $K$ balls of finite radius $\eta_{K}$. By setting the ambiguity set radius $\varepsilon^{K}_{t}$ as the nominal radius $\varepsilon_{t}$ plus a function of $\eta_{K}$, we ensure that the true distribution lies within the adaptive ambiguity set with high probability. This then allows us to derive guarantees on the out-of-sample performance of the online solution $x^{K}_{t}$. To show this, we first quantify the discrepancy between the distributions of interest. Note that the proofs of this section are delayed to Appendix A.2.
Lemma 4.1 (Distance between distributions).
Let $\hat{{\mathbf{P}}}^{K}_{t}$ be the clustered empirical distribution at time $t$, with cluster diameters $\mathrm{diam}_{q}(C^{k}_{t})\leq 2\eta_{K}$, defined according to Definition 3.5 and an arbitrary norm $q$. Then, for all $t$ and $p$, and with the same norm, we have $W^{K}_{t,p}\leq 2\eta_{K}.$ With probability $1-\beta_{t}$, we have $W_{p}(\mathbf{P},\hat{\mathbf{P}}^{K}_{t})\leq\varepsilon_{t}+2\eta_{K},$ where $\beta_{t}$ and $\varepsilon_{t}$ are computed as in Theorem 4.1.
Lemma 4.1 implies that by selecting the radius $\varepsilon^{K}_{t}=\varepsilon_{t}+2\eta_{K}$, where $\varepsilon_{t}$ is the nominal radius for Wasserstein DRO, we can obtain the same finite sample performance guarantee for the online algorithm as Wasserstein DRO. We formalize this result in the following theorem.
Theorem 4.2 (Finite sample performance guarantee).
Under the assumptions of Theorem 4.1, let the adaptive ambiguity set $\mathcal{P}^{K}_{t}$ be defined with $\varepsilon^{K}_{t}=\varepsilon_{t}+2\eta_{K}$. Then, the solution $x^{K}_{t}$ and optimal value $H^{K}_{t}$ of the compressed DRO problem (7) implies the finite sample performance guarantee
${\mathbf{P}}^{t}\left(H_{\star}\leq{\mbox{\bf E}}_{{\mathbf{P}}}[f(u,x^{K}_{t})]\leq H^{K}_{t}\right)\geq 1-\beta_{t}.$
Proof.
This follows from Lemma 4.1 and the same logic as nominal DRO results [38, Theorem 3.5]. ∎
In addition, we can derive the following asymptotic result by applying a similar approach as [38, Lemma 3.7].
Lemma 4.2 (Convergence of distributions).
Assume for all $t$, $\varepsilon_{t}$ is computed as in Theorem 4.1, with a sequence $\beta_{t}$ such that $\sum_{t=0}^{\infty}\beta_{t}<\infty$ and $\lim_{t\to\infty}\varepsilon_{t}=0$. Then, any sequence $\mathbf{Q}_{t}\in\mathcal{P}^{K}_{t}$, where $\mathcal{P}^{K}_{t}$ has radius $\varepsilon^{K}_{t}=\varepsilon_{t}+2\eta_{K}$, satisfies
$\mathbf{P}^{\infty}\left\{\lim_{t\rightarrow\infty}W_{p}(\mathbf{P},\mathbf{Q}_{t})\leq 4\eta_{K}\right\}=1.$
This lemma dictates that asymptotically, the true distribution lies in a ball around the clustered distribution with radius $4\eta_{K}$; the increased radius is the price to pay due to the limited computational budget available to solve (4), and is directly related to the quality of our approximation $\hat{{\mathbf{P}}}^{K}_{t}$. Clearly, if $K$ is high, then $\eta_{K}$ will be low and our asymptotic estimate will be more accurate; viceversa for low $K$. Following [38, Theorem 3.6], we now derive the following result.
Theorem 4.3 (Asymptotic performance guarantee).
Assume for all $t$, $\varepsilon_{t}$ is computed as in Theorem 4.1, with a sequence $\beta_{t}$ such that $\sum_{t=0}^{\infty}\beta_{t}<\infty$ and $\lim_{t\to\infty}\varepsilon_{t}=0$. Also assume the adaptive ambiguity set $\mathcal{P}^{K}_{t}$ is defined with $\varepsilon^{K}_{t}=\varepsilon_{t}+2\eta_{K}$, and that for all $x\in\mathcal{X}$, the constituent functions $f_{j}$ of $f$ satisfy the $M$-Lipschitz condition given in Assumption 2.1 with constants $M_{j}$. Then, the sequence of solutions and optimal values $x^{K}_{t}$ and $H^{K}_{t}$ of the compressed DRO problem (7) satisfies
$\mathbf{P}^{\infty}\left\{H_{\star}\leq\limsup_{t\rightarrow\infty}{\mbox{\bf E}}_{\mathbf{P}}[f(u,x^{K}_{t})]\leq\limsup_{t\rightarrow\infty}H^{K}_{t}\leq H_{\star}+4\max_{j\leq J}M_{j}\eta_{K}\right\}=1.$
Theorem 4.3 shows that, asymptotically, our online in-sample value is an upper bound on the out-of-sample and true values, but is also no more than $4\max_{j\leq J}M_{j}\eta_{K}$ suboptimal compared to the true solution. In other words, there exists some value $\tau$ such that for all $t\geq\tau$, the online solution $H^{K}_{t}\leq H_{\star}+4\max_{j\leq J}M_{j}\eta_{K}$. We note that the upper bound on $H^{K}_{t}$ may be less tight than the ones derived in Section 3.3, but $H^{K}_{t}$ is now an upper bound on $H_{\star}$, regardless of the curvature of $f$, without the need for the term $\Phi^{K}_{t}$. Therefore, for maximum-of-concave $f$, this bound may perform better than the ones in Section 3.3, depending on the specific robust optimization problem.
Overall, to tune the tradeoff between performance and computational effort, we would again be inclined to choose a number $K$ that provides a reasonable $\eta_{K}$ while not being too computationally expensive.
5 Online algorithms for data-compression
In Section 3.1, we gave requirements for the clustering algorithm. In this section, we provide three online clustering algorithms satisfying Assumption 3.1. These algorithms maintain a reasonably accurate approximation of the optimal discretization while minimizing computational effort. Note that they can be easily adjusted to perform updates with a batch of data, instead of a single datapoint, in case the data-arrival frequency is high. Below, we summarize the time and memory complexities of the algorithms.
Table 1: Time and memory complexities of our clustering algorithms. Recall that $n_{t}$ is the total number of datapoints at time $t$, $K$ is the number of clusters, and $d$ is the dimension of the data. We also let II be the number of iterations required for the clustering algorithm’s convergence. The values $G$ and $Q$ are defined in the respective algorithms.
Algorithm Time (initialization) Time (online) Memory
SupCover $O(GK)$ $O(Kd)$ $O(Kd)$
Reclustering $O(In_{0}Kd)$ $O(In_{t}Kd)$ $O((K+n_{t})d)$
OnlineClustering $O(In_{0}Qd)$ $O(IQKd)$ $O((K+Q)d)$
The most time and memory efficient algorithm, as we will describe below, is the SupCover algorithm, but it is the least flexible in terms of cluster assignments, and is only for bounded supports. The Reclustering approach gives the best approximation, but has the highest time and memory complexities. The OnlineClustering approach is suboptimal compared to reclustering, but has lower runtime and memory complexities; it doesn’t need to store the observed dataset.
5.1 A bounded coverage of the support (SupCover)
We begin with an intuitive approach for a bounded support set $S$: partitioning $S$. When $S$ is bounded, we can always cover $S$ with $K$ balls $B_{\eta_{K}}(a^{k})$ with centers $A=\{a^{k}\}_{k=1}^{K}\subseteq S$ and fixed finite radius $\eta_{K}>0$. Specifically, the centers and the radius are selected as follows.
Assumption 5.1.
The parameters $\{a^{k}\}_{k=1}^{K}$ and $\eta_{K}$ are chosen such that, $\forall u\in S$
  1. 1.
    $\min_{k\leq K}\|u-a^{k}\|\leq\eta_{K}$;
  2. 2.
    $u\subseteq\cup_{k\in K}B_{\eta_{K}}(a^{k})$;
  3. 3.
    $\eta_{K}=\min\{\eta\mid\min_{k\leq K}\|u-a^{k}\|\leq\eta,~u\in S\}$.
In fact, such a partitioning is equivalent to an $\eta_{K}$-net of the support $S$, or a $k$-centers problem with order 1. Solving this problem is NP-hard, but we can use an efficient greedy-algorithm approximation [19]. In this greedy algorithm, we begin with a random point, and iteratively find the furthest point from the current point, and set these $K$ total points as the cluster centers. With a given point $a^{k}$, the process of finding the next furthest point depends on the support $S$, and can be solved using the following optimization problem
$\mathop{\rm argmax}_{u\in S}\|u-a^{k}\|,$
whose complexity depends on the shape of $S$. We let the time complexity of this operation be $G$, and note that it needs to be solved $K$ times. The radius $\eta_{K}$ can be set to the maximum of the distances found.
Once the centers are found, we can construct a simple procedure for updating $\hat{{\mathbf{P}}}^{K}_{t}$. The entire process is summarized in Algorithm 1.
Algorithm 1 SupCover
1:given $\mathcal{D}_{0},K,S$
2:choose centers $\{a^{k}\}_{k=1}^{K}$ and radius $\eta_{K}$ satisfying Assumption 5.1
3:assign datapoints to clusters $\{C^{k}_{0}\}_{k=1}^{K}$, and compute $\{n^{k}_{0}\}_{k=1}^{K}$, $\{\theta^{k}_{0}\}_{k=1}^{K}$, $\{\bar{u}_{0}^{k}\}_{k=1}^{K}$
4:for $t=0,1,\dots$ do
5:  $\hat{\mathbf{P}}^{K}_{t}\leftarrow\sum_{k=1}^{K}\theta_{t}^{k}\delta_{\bar{u}_{t}^{k}}$
6:  $C_{t}^{k^{\prime}}\leftarrow C_{t}^{k^{\prime}}\cup\{\hat{u}\} \triangleright$ assign $\hat{u}$ to cluster $k^{\prime}$ following  (18)
7:  $\theta^{k}_{t}\leftarrow$ (19) $\triangleright$ update weights
8:  $\bar{u}^{k^{\prime}}_{t}\leftarrow(n^{k^{\prime}}_{t-1}\bar{u}^{k^{\prime}}_{t-1}+\hat{u})/n^{k^{\prime}}_{t}$ $\triangleright$ update mean
Specifically, the cluster support set $\mathcal{S}_{t}$ can be seen as the Voronoi diagram of $A$, which covers the entire support $S$, and remains constant for all $t$. In addition, each Voronoi region $S^{k}$ has a diameter bounded by $\eta_{K}$. The clustering induced by this support is then
$\hat{u}\in C_{t}^{k^{\prime}}\quad\text{if}\;\hat{u}\in S^{k^{\prime}},\quad\text{where}\quad k^{\prime}=\mathop{\rm argmin}_{k\leq K}\|\hat{u}-a^{k}\|,$ (18)
and $\|\hat{u}-\zeta_{k^{\prime}}\|\leq\eta_{K}$. The set of weights $\{\theta^{k}_{t}\}_{k=1}^{K}$ follow the first-order dynamics
$\theta^{k}_{t}=\begin{cases}\frac{n^{k}_{t-1}+1}{n_{t}}&\text{if}\>\hat{u}\in C^{k}_{t}\\ \frac{n^{k}_{t-1}}{n_{t}}&\text{otherwise}\end{cases}\quad k=1,\dots,K,$ (19)
where $\hat{u}$ is the datapoint at time $t$. Additionally at each time $t$, the mean of the cluster that gained a datapoint is updated. Notably, we do not need to store $\mathcal{D}_{t}$. By keeping track of the centers, means and weights of each cluster, together with a count of the total number of datapoints seen, we can perform the above updates with only the newest datapoint. After performing the update, this datapoint can also be discarded. The algorithm is thus memory efficient; it only requires $O(Kd)$ space.
However, while the updating procedure of this algorithm is simple, its performance is highly dependent on the initial selection of centers $\{a^{k}\}_{k=1}^{K}$. If the covering found is a suboptimal representation of the data, i.e., a single cluster contains multiple distinct modes of the data, then the performance is expected to be suboptimal as well.
5.2 Approximate $k$-centers clustering with warm starts (Reclustering)
To track the given data more closely, we can recompute the approximate $k$-centers clustering at each time step, until the stopping threshold $\tau$. At each time $t$, the cluster support for each $k$ is the Voronoi region around the center $a_{t}^{k}\in A_{t}$. After the threshold $\tau$, we cluster points using the fixed centers $A_{\tau}$, similarly as above. At times $t\leq\tau$, the subsequent clustering can be warm-started with the previous set of centers. We assume without loss of generality that the number of clusters remain $K$, and summarize the algorithm in Algorithm 2.
Algorithm 2 Reclustering
1:given $\mathcal{D}_{0},K,\tau$
[…]
3:assign datapoints to clusters $\{C^{k}_{0}\}_{k=1}^{K}$, and compute $\{n^{k}_{0}\}_{k=1}^{K}$, $\{\theta^{k}_{0}\}_{k=1}^{K}$, $\{\bar{u}_{0}^{k}\}_{k=1}^{K}$
4:for $t=0,1,\dots$ do
5:  $\hat{\mathbf{P}}^{K}_{t}\leftarrow\sum_{k=1}^{K}\theta_{t}^{k}\delta_{\bar{u}_{t}^{k}}$
6:  observe datapoint $\hat{u}$
7:  if $t\geq\tau$ then
8:   $C_{t}^{k^{\prime}}\leftarrow C_{t}^{k^{\prime}}\cup\{\hat{u}\} \triangleright$ assign $\hat{u}$ to cluster $k^{\prime}$ following  (18)
9:   $\bar{u}^{k^{\prime}}_{t}\leftarrow(n^{k^{\prime}}_{t-1}\bar{u}^{k^{\prime}}_{t-1}+\hat{u})/n^{k^{\prime}}_{t}$ $\triangleright$ update mean
10:   $\{\theta^{k}_{t}\}_{k=1}^{K}\leftarrow$ (19) $\triangleright$ update weights
11:  else  repeat lines 2 and 3, warm starting at $A_{t-1}$
In the complexity analysis in Table 1, we assume the $k$-centers approximation to be $k$-means [23], with an iterative algorithm converging in $I$ steps. This method is not memory conservative, as we need to store all datapoints to recompute the cluster information. However, with warm-starts, it can be quite computationally efficient; in practice, warm-starting greatly reduces the required number of iterations.
5.3 Online clustering (OnlineClustering)
For an algorithm that is also memory conservative, we propose an online updating procedure that takes inspiration from a well-known online clustering approach, CluStream [2]. We summarize the algorithm here, and give a more detailed version in Appendix A.3. Similarly to CluStream, we maintain two sets of clusters: a set of $Q\geq K$ microclusters, and a set of $K$ macroclusters. Datapoints are assigned to microclusters on arrival, or initialized as a new cluster based on its distance to current clusters. When new microclusters are initialized, the closest microclusters are merged to maintain the budget $Q$. The $K$ macroclusters are obtained by clustering the microclusters, and can be warm-started.
Differently from CluStream, we allow for separate cluster centers and means; the centers do not shift with the addition of new datapoints, while the means are updated. Both values are updated when clusters are merged. For times $t\leq\tau$, all clustering assignments can be made according to either the microcluster centers or means. We also require the subsequent macrocluster update to be performed online, either with the addition of each datapoint, or before each time the DRO problem (9) is solved. For $t>\tau$, cluster assignments must be made according to the macrocluster centers, which are fixed at time $\tau$. For all clusters, the support is the Voronoi region around its center/mean, with adjustments based on the already-clustered datapoints, which may fall outside the region. For continuous distributions $\mathbf{P}$}, since the datapoints have measure 0, this does not need to be explicitly maintained. Therefore, this algorithm can discard all datapoints once seen, and keep only the aggregated information of size $O((K+Q)d)$. For discrete $\mathbf{P}$}, however, the adjustment to the support is no longer negligible. Nevertheless, discrete distributions generally have a finite set of atoms, so memory (for storing the support information and datapoints) is not a bottleneck.
For the complexity analysis, we again assume the $k$-centers approximation on the $Q$ microclusters to be $k$-means [23]. Compared to the reclustering approach, we are solving a smaller problem, which reduces the theoretical time complexity; the set of $Q$ microclusters allows this method to interpolate between SupCover and Reclustering. Practically, however, the main advantage of this approach is the lower memory requirement.
6 Choosing hyper-parameter values
Above, we detailed various theoretical bounds for the online algorithm, with a given $K$ and with $\varepsilon^{K}_{t}$ chosen according to measure concentration results. In practice, however, we tune these hyper-parameters based on the performance on a validation set of size $N^{\rm val}$.
Choosing $K$
If the initial dataset $\mathcal{D}_{0}$ is large enough, the value of $K$ can be chosen a-priori, using the elbow method on the clustering value $(D^{K}_{0,2})^{2}$ [46, 54]. If the initial dataset is small, we can initialize with a small value of $K$ and adapt it based on the performance on validation data.
Choosing $\varepsilon^{K}_{t}$
Similarly to existing literature on Wasserstein DRO, we tune radius $\varepsilon_{t}$ to minimize validation performance: for every time step $t$, upon constructing the clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$, we solve the optimization problem (9) with a finite set of candidate radii $\varepsilon^{K}_{t}$. For each solution, we use the validation dataset to compute the validation performance. Averaging across multiple repetitions of the experiment, we select the sequence of $\varepsilon^{K}_{t}$ that gives the best validation performance.
7 Numerical examples
We now illustrate the computational performance and robustness of the proposed method on a numerical example. All the code to reproduce our experiments is available, in Python, at
https://github.com/stellatogrp/online_mro.
We run the experiments on the Princeton Institute for Computational Science and Engineering (PICSciE) facility with 35 parallel 2.4 GHz Skylake cores. We solve all optimization problems with the MOSEK [39] optimizer with default settings.
Baselines and metrics
We compare four different approaches, described below.
  • Online clustering (our memory-efficient method). We use our online clustering algorithm, as described in Algorithm 3, to update cluster assignments. This method discards the datapoints once seen, and is thus memory-efficient. Upon constructing the empirical clustered distribution, we solve (9) for the online solution.
  • $k$-means reclustering (our method). Same as above, except using $k$-means clustering at each time step, warm-starting at the previous centers. The clustering procedure is described in Algorithm 2.
  • Wasserstein DRO. At time $t$, we solve (9) with the ambiguity set $\mathcal{P}_{t}$.
  • Sample average approximation SAA. At time $t$, we solve the stochastic optimization problem with respect to the empirical measure $\hat{{\mathbf{P}}}_{t}$ of the training data. We denote this in-sample value as $H^{\rm SAA}_{t}$.
We use a training dataset of up to $T=2000$ time steps (datapoints) for all approaches. For evaluation, we employ both validation and testing datasets, comprising of $N^{\rm val}=N^{\rm test}=200$ datapoints. To limit the total solving time, particularly for the nominal DRO problem, we solve the problems every 100 time steps. We compare the following metrics, with values averaged over 30 repetitions of each experiment. We additionally show the 25-75th percentiles using shaded regions.
  • In-sample certificate. We compute the in-sample certificates obtained with respect to the training data. For nominal Wasserstein DRO, we use the in-sample objective $H_{t}$. For our online methods, we use $H^{K}_{t}+\underline{\psi}_{t}^{K}$ as defined in Theorem 3.6. For all DRO approaches, we do not include the theoretical residual $\rho_{t}$, as the radii $\varepsilon_{t}$ are empirically determined as outlined in Section 6.
  • Out-of-sample objectives. For all methods, we also compute the out-of-sample expected value, which is the stochastic optimization objective computed using the empirical measure of the testing dataset, at the given solutions.
  • Empirical confidence. The empirical confidence, denoted $1-\hat{\beta}_{t}$, is computed as the probability that the certificate upper bounds the out-of-sample value.
  • Computation times. We compare the per-iteration computation times, which include both clustering and solving times for our data-compressed methods.
7.1 Sparse portfolio optimization
We consider a market that forbids short-selling and has $d$ assets [54, 33]. The uncertain parameters are the daily returns, given by $v\in{\bf R}^{d}$. The decisions $x\in{\bf R}^{n}$} represent the percentage weights (of the total capital) invested in the assets. We restrict our selection to at most $\gamma$ assets, given by the 0-th norm cardinality constraint. The distribution $\mathbf{P}$} is unknown, but we have access to a streaming dataset $\mathcal{D}_{t}$, updated each day with a new returns vector. Our objective is to minimize the $\mathop{\bf CVaR}$ with respect to $x$,
$\begin{array}[]{ll}\text{minimize }&\mathop{\bf CVaR}(-v^{T}x,\omega)\\ \text{subject to}&\mathbf{1}^{T}x=1,\quad x\geq 0,\quad\|x\|_{0}\leq\gamma,\end{array}$
which represents the average of the $\omega$ largest portfolio losses that occur. In other words, the $\mathop{\bf CVaR}$ term seeks to ensure that the expected magnitude of portfolio losses, when they occur, is low. The objective can be written as the expectation of the maximum-of-affine functions [51], i.e., ${\mbox{\bf E}}_{{\mathbf{P}}}\left[\tau+(1/\omega)\max\{-v^{T}x-\tau,0\}\right]$. We solve the online DRO problem using the reformulation (9), with $r=1$ and $\|\cdot\|$ the $\ell_{2}$-norm. As $f$ is maximum-of-affine, $\bar{\psi}_{t}^{K}=0$, and $\Phi^{K}_{t}$ converges to $\Phi^{K}_{\star}$, as proven in Theorem 3.9.
Problem setup
We take stock data from the past 5 years of S&P500 (1/1/2020 to 1/1/2025) daily returns, and generate synthetic data from their fitted general Pareto distributions, with correlations preserved using a Gaussian copula. These distributions are chosen to model the data more accurately than a Gaussian fit. We let $\omega=20\%$, $d=50$ stocks, and restrict our portfolio to at most $\gamma=8$ stocks. We initialize with a dataset of size $n_{0}=5$, and show results for $K=15$ and $K=25$. This range is obtained using the elbow method on the initial dataset, and allowing for adjustments.
Figure 1: Sparse portfolio, $K=25$. In-sample certificates, empirical confidence, and per-iteration computation times for the different methods, at $\varepsilon^{K}_{t}=0.0025{(t+n_{0})^{-1/40}}$.
Figure 2: Sparse portfolio. Left: $K=15$, right: $K=25$. Out-of-sample expected values for $\varepsilon_{t}=0.0025{(t+n_{0})^{-1/40}}$.
Results
For all DRO settings, the radius we obtain through cross-validation is $\varepsilon_{t}=0.0025{(t+n_{0})^{-1/40}}$. In Figure 1, we compare the certificates, empirical confidence, and per-iteration computation times at the chosen $\varepsilon_{t}$ sequence, and in Figure 2, compare the out-of-sample performance. We observe that our online data-compressed approaches introduce slight sub-optimality in out-of-sample performance, but provide high-confidence certificates, and offer significant speed-ups compared to nominal DRO. As expected, $K=25$ clusters out-performs $K=15$ clusters, but $K=15$ already gives a good approximation of the nominal DRO performance. The SAA approach improves in performance as the number data points increases, but does not offer a certificate of optimality: the empirical confidence is near 0 for all $t$. Furthermore, it also grows in complexity with sample size and is less efficient than our online approaches. In the long-term, the computation times for our data-compressed approaches can be multiple orders of magnitude faster than that of both SAA and nominal DRO. We find that the $k$-means warm-starting algorithm (reclustering approach) is particularly efficient, achieving lower computation times and better solution quality than the online clustering method, which entails an extra trade-off between optimality and memory efficiency. Nonetheless, the differences are minimal.
In Figure 3, we show the various certificates given in equation (17) (disregarding $\rho_{t}$ in favor of cross-validation), and note that the relationships follow the hierarchy presented. The data-compressed in-sample objective value for this maximum-of-affine problem is lower than the in-sample objective of nominal DRO, but adding the clustering discrepancy makes it a valid certificate. We also show the clustering values for a longer horizon, up to $T=10000$, with $\tau=8500$. We observe that the $p=1$ and $p=2$ clustering distances, $W^{K}_{t,1}$ and $(D^{K}_{t,2})^{2}$, as well as $\Phi^{K}_{t}$, are converging. Moreover, $W^{K}_{t,1}\geq\Phi^{K}_{t}$ and, hence, $\max_{j\leq J}M_{j}W^{K}_{\star,1}\geq\Phi^{K}_{t}$ where the maximum Lipschitz constant is $M_{1}=1/\omega=5$. This illustrates the relative optimality of $\Phi^{K}_{t}$, as opposed to the Wasserstein-1 distance, as proven in Theorem 3.10. In fact, although we have not assumed (or achieved) optimal clustering-induced coupling, the result still held.
Figure 3: Sparse portfolio, $K=25$. The radius $\varepsilon_{t}=0.0025{(t+n_{0})^{-1/40}}$ for all methods. Left: certificates from (17). Right: clustering distances given in Definition 3.4, with $T=10000$. The lighter (red) shade is for the reclustering method.
8 Conclusions
We have introduced an online data compression framework for solving Wasserstein DRO problems with streaming data. Our method constructs adaptive ambiguity sets using online clustering, allowing the uncertainty model to evolve as new data arrives, while maintaining out-of-sample performance guarantees. We analyzed the impact of data compression on solution quality, providing finite-sample and asymptotic performance bounds that relate to nominal bounds for Wasserstein DRO. The framework is compatible with a broad class of clustering algorithms and supports efficient, memory-aware implementations. Numerical experiments in sparse portfolio optimization demonstrate significant computational savings with minimal loss in solution quality.
Acknowledgments
I. Wang and B. Stellato are supported by the NSF CAREER Award ECCS-2239771. I. Wang is also supported by the Princeton Wallace Memorial Fellowship. B. Stellato is also supported by the ONR YIP Award N000142512147. M. Fochesato is supported by G-Research. The authors are pleased to acknowledge that the work reported on in this paper was substantially performed using the Princeton Research Computing resources at Princeton University.
Appendix A Appendices
A.1 Proof of the performance guarantees in Section 3.3
Proof of Lemma 3.2.
For all $t$, let $\tilde{\mathbf{Q}}_{t}\in\tilde{\mathbf{B}}_{\varepsilon_{t}}^{r}(\hat{{\mathbf{P}}}_{t})$ and $\mathbf{Q}_{t}\in\mathcal{P}_{t}$. By [38, Lemma 3.7], we observe ${\mathbf{P}}^{\infty}\{\lim_{t\to\infty}\tilde{\mathbf{Q}}_{t}\rightarrow{\mathbf{P}}\}=1$ and ${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}\mathbf{Q}_{t}\rightarrow{\mathbf{P}}\right\}=1$. The first result follows. If $f$ is Lipschitz, by the Kantorovich-Rubinstein duality [30], and by noting $W_{1}(\tilde{\mathbf{Q}}_{t},\mathbf{Q}_{t})\leq W_{r}(\tilde{\mathbf{Q}}_{t},\mathbf{Q}_{t})\leq 2\varepsilon_{t}$ (by the ordering of Wasserstein distances and the triangle inequality), we obtain $\Delta_{t}\leq 2\max_{j\leq J}M_{j}\varepsilon_{t}$. If $f$ is $L$-smooth, then by the descent lemma and (3), $\mathbf{E}_{\mathbf{\tilde{Q}_{t}}}[f(u,x_{t})]-\mathbf{E}_{\mathbf{{Q}_{t}}}[f(u,x_{t})]\leq\inf_{\pi\in\Pi(\tilde{\mathbf{Q}}_{t},\mathbf{Q}_{t})}\int_{S\times S}\nabla f(v,x_{t})^{T}(u-v)d\pi(u,v)+(L/2)W_{2}(\tilde{\mathbf{Q}}_{t},\mathbf{Q}_{t})^{2}.$ By the Cauchy-Schwartz inequality and Hölder’s inequality with order 2, the integral is upper bounded by $(\textstyle\int\|\nabla f(v,x_{t})\|^{2}d{\mathbf{Q}}_{t})^{1/2}(\int\|u-v\|^{2}d\pi(u,v))^{1/2}=\|\nabla f_{x_{t}}\|_{L_{2}({\mathbf{Q}}_{t})}W_{2}(\tilde{\mathbf{Q}}_{t},\mathbf{Q}_{t})$. When $r\geq 2$, this simplifies to $\Delta_{t}\leq\|\nabla f_{x_{t}}\|_{L_{2}({\mathbf{Q}}_{t})}2\varepsilon_{t}+2L\varepsilon_{t}^{2}$. The result follows.
Proof of Theorem 3.6.
We first prove (15). We begin by observing
$H_{t}=\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x_{t})]\leq\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]\leq\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]+\Phi^{K}_{t},$
with $\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]=H^{K}_{t}$ in the final clause. The first inequality follows from the optimality of $x_{t}$, the second follows from [54, Theorem 5], and $\Phi^{K}_{t}$ is defined in Definition 3.4. From [54, Theorem 4], we note that $\Phi^{K}_{t}=0$ for concave functions ffi.e., when $J=1$. When $f$ is Lipschitz, using Kantorovich-Rubinstein duality [30], we can obtain a second bound on $H_{t}$,
$\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]$ $\leq\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]+\max_{j\leq J}M_{j}(2\varepsilon_{t}+W_{1}(\hat{{\mathbf{P}}}_{t},\hat{{\mathbf{P}}}^{K}_{t})).$
When $f$ is $L$-smooth, using the same procedure as in the proof of Lemma 3.2 above, we obtain a third bound in terms of the $W_{2}$ distance. Combining all gives the bound on $H_{t}-H^{K}_{t}$. For the other side, from the optimality of $H^{K}_{t}$ and [54, Theorem 5],
$H^{K}_{t}\leq\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x_{t})]$ $\leq H_{t}+(\tilde{H}_{t}-H_{t})+{\max_{j\leq J}\frac{L_{j}}{2}}(D^{K}_{t,2})^{2},$
where $\tilde{H}_{t}-H_{t}=\Delta_{t}$ is given in Lemma 3.2. Kantorovich-Rubinstein duality yields
$H^{K}_{t}\leq\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x_{t})]\leq H_{t}+\max_{j\leq J}M_{j}(2\varepsilon_{t}+W^{K}_{t,1}).$
Combining everything, we obtain (15). Using these relationships, and the finite-sample performance guarantee (5) implied by the chosen radii sequence in Assumption 3.2(1), we obtain (17). To prove (16), we note that if the radii are chosen using Theorem 2.1, then with probability $1-\beta_{t}$, ${\mathbf{P}}\in\mathcal{P}_{t},$ and we have, with $\rho_{t}=0$,
${\mathbf{P}}^{{t}}\left(H_{\star}\leq\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]\leq\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]+\rho_{t}\right)\geq 1-\beta_{t}.$ (20)
Applying the bounds we derived for $\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]$ gives the desired result. If the radii are chosen using Theorem 2.2, then as $x^{K}_{t}\in\mathcal{X}$, (20) holds by (6).
A.2 Proofs of Section 4 results
Proof of Lemma 4.1.
By definition of the Wasserstein metric,
$(W^{K}_{t,p})^{p}=W_{p}(\hat{{\mathbf{P}}}_{t},\hat{\mathbf{P}}^{K}_{t})^{p}$ $=\inf_{\pi\in\Pi}\left\{\int_{{S}^{2}}\|u-u^{\prime}\|^{p}\pi(du,du^{\prime})\right\}$
$\leq\sum_{k=1}^{K}\frac{n^{t}_{k}}{n_{t}}\int_{S^{2}}\|u-\bar{u}^{k}_{t}\|^{p}\hat{{\mathbf{P}}}_{t}(u|u^{\prime}=\bar{u}^{k}_{t})(du)$
$\leq\sum_{k=1}^{K}\frac{n_{k}^{t}}{n_{t}}\frac{1}{n_{k}^{t}}\sum_{\hat{u}\in C^{k}_{t}}\|\hat{u}-\bar{u}^{k}_{t}\|^{p}$
$=\frac{1}{n_{t}}\sum_{k=1}^{K}\sum_{\hat{u}\in C^{k}_{t}}\|\hat{u}-\bar{u}^{k}_{t}\|^{p}$
$\leq(2\eta_{K})^{p},$
where the final inequality follows from the cluster diameter of $2\eta_{K}$, using Definition 3.5. Next, by exploiting the triangular inequality, with probability $1-\beta$, we get
$W_{p}(\mathbf{P},\hat{\mathbf{P}}^{K}_{t})\leq W_{p}(\mathbf{P},\hat{{\mathbf{P}}}_{t})+W_{p}(\hat{\mathbf{P}}_{t},\hat{\mathbf{P}}^{K}_{t})\leq\varepsilon_{t}+2\eta_{K},$ (21)
where $\varepsilon_{t}$ is computed as in Theorem 4.1. ∎
Proof of Lemma 4.2.
As $\mathbf{Q}_{t}\in\mathbf{B}_{\varepsilon^{K}_{t}}(\hat{\mathbf{P}}_{t}^{K})$, from the triangular inequality,
$W_{p}(\mathbf{P},\mathbf{Q}_{t})\leq W_{p}(\mathbf{P},\hat{\mathbf{P}}^{K}_{t})+W_{p}(\hat{\mathbf{P}}^{K}_{t},\mathbf{Q}_{t})\leq W_{p}(\mathbf{P},\hat{\mathbf{P}}^{K}_{t})+\varepsilon^{K}_{t}$
From (21) we have $W_{p}(\mathbf{P},\hat{\mathbf{P}}^{K}_{t})\leq\varepsilon_{t}+2\eta_{K}=\varepsilon^{K}_{t}$. Thus, we have $\mathbf{P}(W_{p}(\mathbf{P},\mathbf{Q}_{t})\leq 2\varepsilon^{K}_{t})\geq 1-\beta_{t}$. Since by definition $\sum_{t=0}^{\infty}\beta_{t}<\infty$, then by the Borel-Cantelli Lemma we get
$\mathbf{P}^{\infty}\{W_{p}(\mathbf{P},\mathbf{Q}_{t})\leq 2\varepsilon^{K}_{t}\text{~~for all large enough $t$}\}=1.$
Since by definition $\lim_{t\rightarrow\infty}2\varepsilon^{K}_{t}=4\eta_{K}$, we conclude $\lim_{t\rightarrow\infty}W_{p}(\mathbf{P},\mathbf{Q}_{t})\leq 4\eta_{K}$ almost surely. ∎
Proof of Theorem 4.3.
By the finite sample performance guarantee in Theorem 4.2 and the summability of $\beta_{t}$, the Borel–Cantelli Lemma implies that
$\mathbf{P}^{\infty}\left\{H_{\star}\leq\limsup_{t\rightarrow\infty}{\mbox{\bf E}}_{\mathbf{P}}[f(u,x^{K}_{t})]\leq\limsup_{t\rightarrow\infty}H^{K}_{t}\right\}=1.$
Now, choose any $\gamma>0$, and fix a $\gamma$-optimal solution $x_{\gamma}\in\mathcal{X}$ of (1) such that ${\mbox{\bf E}}_{{\mathbf{P}}}[f(x_{\gamma},u)]\leq H_{\star}+\gamma$. For each $t$, we choose a $\gamma$-optimal distribution $\mathbf{Q}_{t}\in\mathcal{P}^{t-1}$ such that $\sup_{\mathbf{Q}\in\mathcal{P}^{t-1}}{\mbox{\bf E}}_{\mathbf{Q}}[f(x_{\gamma},u)]\leq{\mbox{\bf E}}_{\mathbf{Q}_{t}}[f(x_{\gamma},u)]+\gamma.$ Then, we observe that
$\limsup_{t\rightarrow\infty}H^{K}_{t}$ $\leq\limsup_{t\rightarrow\infty}\sup_{\mathbf{Q}\in\mathcal{P}^{t-1}}{\mbox{\bf E}}_{\mathbf{Q}}[f(x_{\gamma},u)]$
$\leq\limsup_{t\rightarrow\infty}{\mbox{\bf E}}_{\mathbf{Q}_{t}}[f(x_{\gamma},u)]+\gamma$
$\leq\limsup_{t\rightarrow\infty}{\mbox{\bf E}}_{\mathbf{{\mathbf{P}}}}[f(x_{\gamma},u)]+\max_{j\leq J}M_{j}W_{1}(\mathbf{Q}_{t},{\mathbf{P}})+\gamma$
$\leq H_{\star}+4\max_{j\leq J}M_{j}\eta_{K}+2\gamma,$
where the first inequality follows from the optimality of $x^{K}_{t}$ and $H^{K}_{t}$, and the third inequality follows from Kantorovich-Rubinstein duality [30] and the Lipschitz condition of $f$. The final inequality follows from Lemma 4.2. Since we chose $\gamma>0$ arbitrarily, we can conclude that $\limsup_{t\rightarrow\infty}H^{K}_{t}\leq H_{\star}+4\max_{j\leq J}M_{j}\eta_{K}$. ∎
A.3 Online clustering algorithm (Section 5.3)
We give details for the online clustering algorithm, introduced in Section 5.3. We keep track of two sets of clusters: a set of $Q\geq K$ microclusters, and a set of $K$ macroclusters. For simplicity, below we assume the number of microclusters to be $Q$ and the number of macroclusters to be $K$; if the number of datapoints is smaller, i.e., $Q_{t}\leq Q$ and $K_{t}\leq K$, the arrival of new datapoints result in new clusters.
Initialization
We initialize the problem by solving approximately the $k$-centers problem (11) with respect to $\hat{{\mathbf{P}}}^{0}$, allowing up to $Q$ clusters. This set of centers is denoted $A^{Q}_{0}$. We assign all points to the closest center, and define the support of the $q$-th cluster $S^{q}_{0}$ as the Voronoi region around its center $a_{0}^{q}\in A^{Q}_{0}$. Note that the Voronoi regions need not be explicitly stored; they are implicitly defined by their centers. For each cluster, we note its center, mean, root-mean-squared-error (RMSE), and weight. These are denoted the microclusters. We refer to the RMSE of each cluster as $\eta^{q}_{t}$, and for clusters with only a single datapoint, heuristically initialize it as twice the minimum RMSE of all clusters.
To create $K$ macroclusters, we solve approximately the $k$-centers problem with respect to $A^{Q}_{0}$, and obtain a set of centers $A^{K}_{0}$. Each microcluster is assigned to the macrocluster with the closest center; each macrocluster then contains all the points of its constituent microclusters, and has their combined weight. The mean of the macrocluster is therefore the weighted average of the means of the constituent microclusters. The support of the $k$-th macrocluster is the Voronoi region around the $k$-th center $a_{0}^{k}\in A^{K}_{0}$, plus the datapoints assigned to it, and minus the datapoints assigned to other clusters, i.e.,
$S^{k}_{t}=(V(a_{t}^{k})\cup\{C^{k}_{t}\})/\{C^{k^{\prime}}_{t}\}_{k^{\prime}\neq k}.$ (22)
In this manner, the set of supports $\mathcal{S}_{t}$ satisfy Assumption 3.1. The necessity of this adjustment of the support follows from the two-layer nature of the clustering algorithm. It is possible for a datapoint to be clustered into the $q$-th microcluster, which is assigned to the $k$-th macrocluster, but distance-wise is actually closer to the center of $k^{\prime}$-th macrocluster for some $k^{\prime}\neq k$.
Online updating procedure
For $t\geq 1$, when we observe a new datapoint $\hat{u}$, we calculate its distance to the microcluster whose support it falls on, i.e., if it falls on support $S^{q}_{t-1},$
$d_{t}=\|\hat{u}-a^{q}_{t-1}\|.$
If this value is below $2\eta^{q}_{t-1}$, where $\eta^{q}_{t-1}$ is the RMSE, we assign it to this cluster. The multiplier of 2 is selected according to CluStream sensitivity analysis [2, Section 6.4]. We then update all microclusters weights, as well as the mean and RMSE of the assigned microcluster. The macroclusters are also adjusted accordingly, based on their constituent microclusters.
On the other hand, if $d_{t}>2\eta^{q}_{t-1}$, we create a new microcluster with this datapoint as its center, and assign it the weight $1/n_{t}$. The weights of the other microclusters are decreased accordingly, following (19). The RMSE of this cluster is initialized heuristically to be twice the minimum RMSE of all other existing clusters. If the total number of microclusters exceeds $Q$ with this addition, we merge the two closest microclusters based on the distances between their centers. The center of the merged microcluster will be the weighted average of the centers of the two constituent microclusters, with the new mean, RMSE, and weight calculated accordingly. The supports of the microclusters will be reassigned as in (22). Then, we again perform approximate $k$-centers on the $Q$ microclusters, warm-staring at the previous centers, and generate $K$ macroclusters. The parameters of the macroclusters are assigned the same manner as in initialization.
At some time step $\tau<\infty$, we terminate the support updating procedure, and only cluster points based on the latest support $\hat{S}^{K}_{\tau}$.
The online DRO algorithm with online clustering is summarized in Algorithm 3. If a parameter is not explicitly updated, we assume it inherits the value from the previous time step.
Remark
In the algorithm and description, the cluster assignments are assumed to use the cluster centers. The mean can also be used, however, up to time $t\leq\tau$.
Overall, by keeping $Q\geq K$ microclusters, we allow for a finer clustering algorithm than keeping only $K$ clusters at all times. In this manner, the microclusters are allowed to switch macrocluster assignments when the $k$-centers update is performed, thereby minimizing the information loss induced by the $K$-cluster budget.
Algorithm 3 OnlineClustering
1:given $\mathcal{D}_{0},K,Q,S,\tau,\{\varepsilon^{K}_{t}\}_{t=0,1,\dots}$
[…]
3:assign all datapoints to microclusters $C^{q}_{0}$, and compute $n^{q}_{0}$, $\theta^{q}_{0}$, $\bar{u}_{0}^{q}$, and $\eta^{q}_{0}$
[…]
6:assign corresponding microclusters to macroclusters $C^{k}_{0}$, and compute $n^{k}_{0}$, $\theta^{k}_{0}$, and $\bar{u}_{0}^{k}$
[…]
8:for $t=1,2,\dots$ do
9:  $\hat{\mathbf{P}}^{K}_{t}\leftarrow\sum_{k=1}^{K}\theta_{t}^{k}\delta_{\bar{u}_{t}^{k}}$
10:  observe datapoint $\hat{u}$, and set $d_{t}\leftarrow\|\hat{u}-a^{q^{\prime}}_{t-1}\|$, where $q^{\prime}$ is chosen such that $\hat{u}\in S^{q^{\prime}}_{t-1}$
11:  if $t\geq\tau$ then
12:   $C_{t}^{k^{\prime}}\leftarrow C_{t}^{k^{\prime}}\cup\{\hat{u}\} \triangleright$ assign $\hat{u}$ following  (18)
13:   $\theta^{k}_{t}\leftarrow$ (19) $\triangleright$ update weights
14:   $\bar{u}^{k^{\prime}}_{t}\leftarrow(n^{k^{\prime}}_{t-1}\bar{u}^{k^{\prime}}_{t-1}+\hat{u})/n^{k^{\prime}}_{t}$ $\triangleright$ update mean
15:  else if $d_{t}\leq 2\eta^{q^{\prime}}_{t}$ then
16:   $C_{t}^{q^{\prime}}\leftarrow C_{t}^{q^{\prime}}\cup\{\hat{u}\}$
17:   $\theta^{q}_{t}\leftarrow$ (19) $\triangleright$ update microcluster weights
18:   $\bar{u}^{q^{\prime}}_{t}\leftarrow(n^{q^{\prime}}_{t-1}\bar{u}^{q^{\prime}}_{t-1}+\hat{u})/n^{q^{\prime}}_{t}\triangleright$ update mean
19:   $\eta^{q}_{t}\leftarrow(n^{q^{\prime}}_{t-1}\eta^{q^{\prime}2}_{t-1}+\|\hat{u}-\bar{u}^{q^{\prime}}_{t}\|_{2}^{2})/n^{q^{\prime}}_{t}\triangleright$ update RMSE
20:   $\theta^{k}_{t}\leftarrow$ (19) $\triangleright$ update macrocluster weights
21:   update $\bar{u}^{k^{\prime}}_{t}$ for microcluster $k^{\prime}$, where $C^{q^{\prime}}_{t}\subseteq C^{k^{\prime}}_{t}$
22:   $\hat{S}^{K}_{t}\leftarrow$ (22)
23:  else
24:   assign $\hat{u}$ to a new cluster $q^{\star}$, initialize $\eta^{q^{\star}}_{t}\leftarrow\min_{q\leq Q}2\eta^{q}_{t-1}$
25:   $\theta^{q}_{t}\leftarrow$ (19) $\triangleright$ update weights
26:   $A^{Q}_{t}\leftarrow$ merge the two microclusters with the closest centers
27:   $\hat{S}^{Q}_{t}\leftarrow$ (22)
28:   $A^{K}_{t}\leftarrow$ find $K$ centers using approximate $k$-centers on $A^{Q}_{t}$
29:   assign microclusters to macroclusters $C^{k}_{t}$, and compute $n^{k}_{t}$, $\theta^{k}_{t}$, and $\bar{u}_{t}^{k}$
30:   $\hat{S}^{K}_{t}\leftarrow$ (22)